A Bibliography of Publications in American Mathematical Monthly: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 January 2024 Version 1.23

Title word cross-reference

(log_{x} N)′ [333]. (x + 1)^{y} − x^{z} = 1 [846]. 0 [495, 1]. 1 [495, 1]. 1/p [511]. 10 [140]. 168 [538]. 2 [335]. $29.50 [793]. 2^{n}.n! [1005]. 2 × 2 [26]. 3 [829, 583]. 4^{n} − 2 [749]. A [435]. A^{∗}A = B^{∗}B [607]. AB [620]. BA [620]. \binom{2n}{n} [260]. mod5 [983]. C^{1} [833]. cos(2π/n) [322]. d [845]. dy/dx [449]. e^{x} [860]. f(x, y) [469]. f − 1 = 1/f [853]. Γ [582, 949]. H [342]. ∫_{−∞}^{+∞} e^{ax(x−2b)}dx [855]. ∫_{0}^{∞} cos x^{2} dx, [20]. ∫_{0}^{∞} e^{−x^{2}} dx [20]. ∫_{0}^{∞} sin x^{2} dx [20]. k [934, 588]. L^{1}(μ) = L_{loc}^{1}(μ) [184]. L^{4} [21]. L^{p} [256]. L^{p}(μ) ⊆ L^{q}(ν) [145]. l_{2} [579, 57]. l_{p} [414]. ln k [832]. M_{11} [524]. M_{23} [524]. M_{n}(Z) [269]. n! [96]. N [28, 872, 964, 592, 862, 247, 429, 530, 908].

NF [738]. ω(n) [82]. p [619, 149, 695, 412]. P^{2} [357]. p ≡ 1 (mod 4) [34]. ϕ [674]. ϕ(30n + 1) [948]. Φ_{pq}(X) [618]. π [741, 685, 694, 951, 489]. π^{2} [105]. ±1 [21]. Q_{c}(x) = x^{2} + c [399]. R [35, 226]. R^{3} [357]. R^{n} [62, 588]. S_{6} [315]. σ [19]. \sqrt{−1} [997]. \sqrt{2} [473]. \sqrt{2} × \sqrt{3} = \sqrt{6} [257]. ∑_{n=0}^{∞} 1/n! [619]. tan^{(k)} z [160]. ℘ [512]. x [860]. x/(sin x) [756]. x = f(x′) [833]. x^{2} + y^{m} = z^{2n} [7]. x^{2} + y^{m} = z^{2n} [65]. x^{5} + ax + b [465]. x^{n} = 1 [235]. Z^{3} [976]. Z_{2} × Z_{n} [870]. Z_{m_{1}} × ⋯ × Z_{m_{r}}intoZ_{k_{1}} × ⋯ × Z_{k_{s}} [789]. Z_{n} [984]. ζ(3) [41].

-adic [619, 149]. -body [28]. -dimensional [845]. -ellipses [908]. -function [512, 582]. -Game [872, 19]. -generator [335]. -gons [530]. -norm [414]. -polygons [342].

-queens [429]. -smooth [829]. -solutions [833]. -sphere [583]. -Tuples [964]. -volume [588].

101 [462, 408, 442]. 103xx [367]. 168pp [793]. 17th [644]. 1832 [793]. 1860s [351].

500th [448]. 51030a [442]. 53 [495].

8 [495]. 83g [70]. 85m [655]. 87e [68]. 89 [70]. 89c [67]. 89f [65]. 89g [66]. 89m [69].

90e [39]. 93 [68]. 93a [740]. 93c [382]. 93f [441]. 94 [67]. 95 [66, 65, 69]. 95c [462]. 95i [623]. 95k [442]. 96 [39]. 99 [441, 740, 382].

= [404].

Abe [655, 899, 719, 743, 516, 536]. Abel [518]. Abelian [336, 133, 645, 488]. abstract [605, 516, 536]. accumulation [613]. acting [372]. Actions [775]. actuarial [617]. add [540]. Addendum [65, 495, 446]. Adding [493, 33]. Addition [984, 1115]. Additive [1025]. adic [619, 149]. Advanced [8, 86, 107, 369, 368]. aerodynamic [903]. African [1086]. After [837]. aftermath [902]. again [334, 30]. Age [565]. aid [612]. aided [717]. Aitken [97]. Alan [878, 517]. Alexander [354]. Algebra [836, 496, 899, 712, 27, 173, 156, 193, 288, 78, 649, 353, 516, 536, 970, 970]. Algebraic [133, 655, 397]. Algebras [776, 98, 216]. algorithm [686, 258, 26, 489, 30]. Algorithms [737, 786, 479]. Alice [771]. Almost [800, 665]. Also [565, 392]. alternating [197]. alternative [36, 186]. AM [1114]. amazing [662, 584]. Ambiguous [64]. Amer [66, 67, 65, 441, 740, 462, 495, 382, 68, 69, 408, 442, 39, 70]. American [291]. among [126]. Amusing [756, 364]. analogy [193]. Analysis [1117, 811, 772, 877, 821, 768, 619, 183, 240, 848, 835, 849, 655]. analytic [127]. analytical [157]. Andrew [1117, 572, 699]. Andrews [39, 39]. Angle [919, 559, 377]. anniversary [280, 448]. annulus [181]. Answer [1005]. Answers [504]. Anthropomorphic [113]. Aperiodic [301]. Apery [1110]. Appearance [1109]. Application [766, 785, 452, 182, 151, 333]. Applications [869, 649, 715, 535]. appreciation [381]. Approach [722, 978, 281, 450, 143, 601, 288, 192, 150]. Approximate [702]. approximately [189]. Approximation [479, 981, 342, 117, 718, 155, 96, 891, 229]. approximations [897, 375]. April [673]. Apropos [441]. Arbitrary [982, 453, 344, 81]. Archer [317, 329, 337, 343, 351, 359]. Archimedes [797]. Archimedian [419]. Arcs [121]. Area [1099, 323, 573, 304]. Areas [720, 520, 634, 114, 342, 304]. argument [603]. Arguments [737]. arising [498]. Aristotle [830]. Arithmetic [1080, 256, 622, 507, 405, 543, 533, 54, 674, 506]. arithmetic-geometric [543, 533]. arithmetic-mean [405]. Arrangements [975, 451]. Arthur [209]. aspect [488]. aspects [215, 668]. Assertions [826]. assistance [719]. associated [114, 480]. Association [746]. Astronomy [759]. asymptotic [8, 86]. atonal [565]. attracting [687]. Augustus [977]. authorship [838]. automata [569, 24, 19]. automorphisms [315, 523]. avatars [712]. average [540, 304]. Avoiding [505, 995]. Award [379, 209, 572, 771, 290, 656, 17, 900]. Awards [1030]. Away [732]. axiomatic [281, 480]. Axiomatization [1012, 1070]. axioms [901, 645]. axis [35, 344].

B [462, 382, 837, 70, 319]. Bôcher [207]. Babylonian [258]. Back [940, 972, 1000, 1031, 1059, 1088, 1120, 802]. Baire [168]. ball [596]. balloon [461]. ballot [165]. balls [651]. Banach [686, 698].

Banchoff [442]. Barbeau [837]. Baroque [565]. Barrow [1118]. Barrow-Green [1118]. base [137]. baseball [643]. based [90]. Basel [623]. Bases [859, 1025, 638, 842]. basis [578]. Baxter [39]. be [641, 228, 889, 394, 753]. Beauzamy [455]. became [417]. Beckman [731]. Beckman-Quarles [731]. Beeblebrox [740, 231]. beginners [371]. behaviour [157]. Behind [879, 437]. Berlin [337]. Bernoulli [276, 154, 31, 460]. Bernstein [981]. Bessel [205]. best [187]. Beta [1109]. Betting [868]. between [590, 1083, 362, 559, 620, 430, 82, 661, 40, 503]. Beverly [851]. Beyond [776]. Bicycle [795]. Bieberbach [909, 650]. big [298]. Bill [673]. Billiards [244, 715]. Binary [934, 174, 10, 14, 278, 511, 239]. Binet [906]. Binomial [965, 564, 79, 8, 11, 86, 15, 254, 548]. bipartite [419, 661]. Birkhäuser [793, 623]. birth [899, 648]. Birthdates [747]. birthday [61]. bisection [4, 67]. bisector [377]. Bisectors [338]. bitangent [18]. Blatter [1117]. Block [1097]. Blows [1092]. Body [1118, 28]. Bombieri [455]. Bonnet [606]. Book [1117, 998, 745, 997, 938, 1118, 1057, 1028, 1086, 937, 969, 970, 836]. Boolean [98, 216]. Bored [701]. Borel [127]. Borromean [144, 350]. Borsuk [742]. Boundary [877]. Bounded [1004, 218]. bounds [392]. Box [180, 1117]. Box-spline [180]. boxes [475]. brachistochrone [501, 586]. Brain [1028, 231, 740]. Brainteasers [837]. Brams [878]. branches [107, 369, 368]. Branching [174]. Bredon [782]. breezes [526]. Bricklayer [774]. Bricklaying [302]. bridges [891]. Brouwer [742]. Bruce [1028]. bubbles [402]. Buffon [446, 422, 383]. Building [669]. Bull [766]. Bunyakowski [73]. Burnside [194, 844]. bursts [132]. Burtin [918]. Butterfly [222, 858].

C [1117, 462, 382, 970]. Cake [878, 471, 686]. Cake-Cutting [878]. Cakes [779]. Calculating [476, 255, 611]. Calculus [763, 728, 1004, 1062, 727, 725, 849, 734, 1035, 884, 252, 690, 715, 436, 457, 649, 672, 671, 392]. Calinger [688]. Cam [375]. cameras [840]. Cameron [970]. campanological [907]. Can [753, 619, 394]. cannot [641]. canonical [594]. Cantor [869, 138, 447, 430, 854, 943, 912]. Cantor-type [912]. Cantwell [851]. Carathéodory [400, 115]. cards [77]. Care [996]. Careers [699]. Carlyle [119]. Carries [662]. Cartwright [640]. Carus [902]. Case [977, 642, 240, 669, 529]. cases [157]. casino [470]. Casti [889]. cat [251]. Catalan [1096, 614, 749]. catches [643]. Categories [758]. catenoid [90]. Cattle [797]. Cauchy [73, 719, 136, 150, 562]. Cauchy-Bunyakowski-Schwarz [73]. Cayley [98, 871, 204, 169, 984]. ceiling [77]. celebration [280]. Celestial [709, 648]. cellular [569, 19]. Center [1085, 992]. centers [693]. central [517]. century [436, 457]. Certain [819, 198, 743, 81]. chain [403]. challenge [110]. champagne [644]. Chan [1117]. Change [1004, 974, 845]. Chaos [1036, 424, 943, 232, 509, 219, 546, 502, 625, 691, 404]. Chaotic [332, 132, 301, 157]. character [58]. Characteristic [920, 1093, 574]. Characteristics [865]. Characterization [465, 138, 133, 77, 303, 168, 22]. Characterizations [933, 413]. Characterizing [682]. Charles [1117, 379, 290, 656, 572, 771, 900, 612]. Charting [583]. Chatters [970]. Chebychev [502, 341]. Chebyshev [961, 191]. chemistry [173]. Chernyak [837]. Chessboard [867, 482, 591]. Chicken [837]. Choice [726]. Christian [1117]. Christmas [166]. Chu [455]. Chui [1117]. Circle [800, 390, 109, 199, 22, 520, 355]. Circles [1083, 689, 119, 621, 144, 360, 153]. circulants [216]. circular [406, 442]. Circumscribed [153]. Clarke [803]. class [499, 467, 134]. Classes [865, 804, 493, 692, 141, 514]. Classical [565, 849, 848]. classification [213]. classifications [364]. Clinton [673]. Closed [947]. Closed-Form [947]. Cluster [895]. Codes [747]. Coefficient [952, 564, 8, 11, 86, 15]. coefficients [79, 206, 21, 179, 548]. coins [282]. Colaw [291]. collaboration [640]. Color [673, 1057, 515]. colorful [666]. Coloring [68, 3, 1067, 162, 538]. COMAP [769]. combinations [506]. combinatorial [227]. combinatorics [662, 570]. come [385]. Comment [462, 408, 39, 663]. Commentary [672]. Common [1050, 541, 696, 122, 529]. Commutative [970, 43]. Commuting [1050, 1097]. Commuting-Block [1097]. compact [133]. compactness [221]. Competition [729, 857, 1066]. competitive [456]. complement [354]. complementary [559]. complements [122]. complete [831, 834, 661, 458]. Complex [835, 1023, 866, 464, 132, 270, 171, 137, 28]. Complication [1036]. Composite [1038]. compositions [106, 250]. comprehension [516, 536]. compression [328]. Computation [883, 1122]. Computational [821, 89]. computations [558, 660]. Computer [775, 645, 717, 703]. computer-aided [717]. Computing [1086]. concave [833]. concept [605]. Conceptual [758, 719]. Concerned [753]. concerning [537, 504, 236]. concise [578]. Concurrent [1051]. condition [62, 228, 677]. Conditional [1006]. cone [478]. cones [637]. confidence [72]. configurations [316]. conflict [657]. conflict-free [657]. Confronting [754]. Congruence [755, 443, 983, 493, 695]. Congruences [514]. Congruent [1083, 600, 360]. congruential [223]. Conics [856, 114].

Conjecture [783, 817, 277, 567, 909, 481, 22, 666, 614, 319, 584, 673]. conjugacies [892]. Conjugacy [804, 514]. Connected [930, 659, 1079, 56]. Connections [240]. consecutive [411]. Constance [888]. Constant [952, 1110, 949, 454, 321, 283, 297, 521]. constants [1122]. Constrained [1052, 477, 346, 553]. Construction [1068, 217, 742, 532]. constructions [119]. Constructivism [938]. contained [506]. containing [464]. continental [690]. Continued [219, 502, 951, 111, 83, 124, 898]. continuity [196, 682]. Continuous [56, 151, 981, 454, 201, 687]. contour [855, 171]. contraction [151]. Control [1036, 1033, 1063]. Controllability [1033]. converge [619]. Convergence [186, 1006, 321, 178, 413, 263]. Convergence-preserving [186]. Convergent [936]. Converse [873, 751, 950, 599, 148, 562]. Converses [234]. Convex [720, 881, 833, 571]. Convexity [87, 1084, 788]. Convolution [1024]. Conway [745, 942, 110, 92]. corner [30]. Coronas [805]. correct [685]. Correction [740, 382]. Correlated [963]. Cosine [539]. Count [498]. Count-wheels [498]. Countable [724, 335]. counterexample [63, 25, 384]. Counting [305, 822, 976, 423, 681]. Course [970]. Covered [931]. covering [276, 177, 277]. Covers [724, 5, 66, 236]. Cramer [1078]. Creates [890]. credit [470]. criterion [635, 116, 272, 40]. critical [485, 305, 346, 477]. Cross [453, 190, 283]. crossing [661, 458]. cryptography [352]. crystallographic [650]. cube [271, 315]. cubes [311]. Cubic [385]. Cup [1056]. curiosity [333]. curious [423]. currencies [373, 408]. current [617]. Curtis [809]. Curvature [766, 91, 1065, 297]. curve [544, 121, 40]. Curves [521, 1065, 406, 62, 625, 442, 566, 554].

Cutting [779, 878, 686]. Cycle [774]. cyclic [288, 247]. cyclotomic [676, 618].

D [744, 441, 68, 878, 408, 808, 795]. Damped [1036]. Dan [795]. dark [102]. Data [328, 150]. dating [657]. Daubechies [327]. David [732, 744]. Day [673]. days [329]. dead [690]. Deadly [883]. Deborah [656]. decimal [347, 426]. Decision [910, 352]. decline [359]. Decomposition [78, 24, 403, 577, 439]. Dedekind [952, 257]. deducible [432]. Defining [1096]. definite [116]. definition [232, 546, 449, 691]. Degenerate [723]. Dégot [455]. degree [518]. Dehaene [890]. dense [64]. Densest [360]. Denton [793]. dependence [903]. Depends [1065]. deranged [12, 16]. Derangements [166]. Derek [850]. derivation [75, 96, 551]. Derivative [798, 472, 1039, 45, 115]. Derivatives [858, 215, 602]. Descartes [816, 1068, 876]. Descent [800]. descents [418]. description [157]. Design [1086]. designer [851]. desk [509]. Determinant [1023]. determinantal [666]. Determinants [495, 1097, 1, 484, 601]. determination [208]. determine [507]. Devaney [232, 546]. development [623, 697, 719]. developments [842]. deviations [718]. Devlin [876]. diagonalizability [60]. Diagonalizable [736]. Diagonalization [43]. Diagonalize [867]. diagram [364]. Diameters [813]. dice [916, 410]. Dickson [1003]. Did [795, 508, 385]. Difference [733, 654, 430]. Differences [935, 917]. differentiable [151, 201]. Differential [851, 812, 807, 570]. differentiate [175]. Differentiating [604]. Differentiation [920, 189, 139]. digit [452, 500]. digital [192]. digits [685, 489]. Digraphs [1079]. Dilemma [233]. Dilworth [403]. dimension [453, 573]. Dimensional [773, 779, 845]. Dimensions [975, 819, 592, 218, 743]. Dinitz [584].

Diophantine [655, 65, 293, 7, 846]. Direct [742]. Directions [982, 244]. Dirichlet [432, 1064, 42, 356]. Disc [1037, 374]. Discontinuous [1111]. discovered [575]. Discovering [744, 710]. Discovery [777, 813]. Discrete [731, 251, 416]. discriminant [635]. discussion [186]. disk [42]. Disks [129, 931, 591]. Dispute [878]. disputed [838]. Dissections [787]. Dissertation [953]. Distance [167, 1065, 1067, 908]. distant [843]. distinct [493]. Distinguish [819, 743]. Distinguished [17, 379, 572, 771, 290, 209, 656, 900]. Distributed [1096, 781]. distribution [124, 150]. divergence [651, 413]. Diversions [837]. Divide [964]. Divisibility [1081, 279]. Division [878, 776, 1094, 471]. divisor [696]. Divisors [1027, 54, 541, 503]. Dixon [136]. Dmitri [689]. DNA [609]. Do [728, 435, 753, 252, 703]. Does [619, 853, 901, 507, 743, 819, 607]. Doing [786]. Domains [720, 275, 665]. Don’t [980, 802, 1092]. Doob [762]. door [912]. dot [437]. Double [802]. Down [484, 84]. Dr [379, 290, 572, 771, 900]. Dr. [209, 656]. drag [903]. dragging [358]. dream [574]. drops [418]. Drums [483]. Dual [826, 217]. Duality [1033]. dynamical [77]. Dynamics [773, 132, 170]. Dynkin [364].

E. [808, 768]. E3062 [366]. E3267 [6, 310]. E3303 [10, 14]. E3309 [11, 15]. E3335 [12, 16]. E3411 [106, 250]. E3415 [289, 118]. E3429 [108, 200]. E3463 [109, 199]. Early [759, 716, 170]. Easy [1112]. Economics [821]. Ed [837]. edited [899, 699, 688, 864, 795]. edition [837]. editor [50, 825, 30]. editorial [719]. editors [284]. Education [1034, 753]. Effect [953]. Effects [1062]. Eglash [1086]. Ehrenpreis [159]. Ehrhart [952]. eiconal [487]. eigenvalues [464, 23, 26]. Eigenvector [1050]. eigenvectors [26]. Eighth [865, 857]. eighties [91]. eikonal [487]. Einstein [382, 243]. elastic [576]. electricity [508]. Elegant [760, 951]. Elementary [289, 944, 325, 818, 118, 772, 6, 12, 10, 11, 106, 16, 14, 15, 108, 109, 200, 199, 310, 440, 250, 20, 884, 524, 416, 143, 444, 78, 350, 308, 159, 28, 906, 646, 245, 150]. Elements [964, 495, 1]. Elimination [785]. Ellipse [1056, 393, 545]. ellipses [908]. Ellipsoids [1064]. Elliptic [554, 544, 249]. Elyadi [733]. embeddable [357]. Embedding [335, 222]. Emerging [1090]. Empty [1095]. enclose [634]. end [690]. Endnotes [825]. Energy [737]. Engineering [781]. Engineers [1117]. enigmatic [904]. enough [48, 481]. Enrico [612]. Enrique [877]. Entire [561, 202, 487]. envelope [407]. envy [471]. envy-free [471]. Equal [680, 161, 323]. equality [184]. Equation [812, 1082, 846, 7, 65, 205, 487, 420, 204, 150]. Equations [760, 1012, 1070, 807, 881, 75, 385, 553, 518, 81, 851, 733]. Equidissections [610]. equilateral [360]. equivalent [475]. Erdos [1028, 287, 783, 817, 653, 1028]. ergodic [203, 58]. Ernest [938]. Errata [66, 67, 68, 69, 70]. Erratum [442]. essay [508]. essays [589]. Essentially [1111]. estimates [46]. Euclidean [702, 266, 665, 598, 30]. Euler [462, 432, 944, 392, 321, 380, 395, 462, 574, 949, 376, 663, 674, 593]. Euler-Gergonne-Soddy [593]. Eulerian [1024]. evaluating [855]. evaluations [20]. Évariste [793]. eversion [478]. Every [388, 454, 555, 34]. Everyday [710]. evidence [609]. Evolution [697, 719, 953, 767, 791, 743, 496, 463, 378, 571]. exact [75, 5, 66, 276, 558]. Examination [764, 385]. Example [818, 1011, 275, 298]. Examples [5, 66]. Exceptional [875]. exchange [505]. excursion [912]. Exercise [1053, 212]. Existence [914, 831, 859, 377, 628]. expansion [522, 8, 86, 511]. Expansions [934]. Experience [689]. Experimental [1090]. Experimentation [481]. explanation [345]. explicit [696]. exploding [132]. Exploratory [722]. Exploring [501]. exponential [132, 551]. Exponentiation [474]. exponents [896]. exposition [221, 583]. expository [589]. expressible [411, 388]. Expressing [636]. expressions [660, 438]. extended [37]. Extending [35]. Extension [982, 45]. extensions [52, 94, 288, 60]. extrema [469].

F. [758]. Fabian [633]. Factorial [862]. Factoring [862]. factorization [275, 440]. Fail [873]. fails [579]. Fair [878, 1094]. fall [490, 830]. Family [773]. Farey [913]. Farkas [886]. Farmer [744]. Fast [794, 342, 540]. Faster [1095, 174]. favorite [544]. Federalist [838]. Feedback [1063]. Fermat [673, 1007, 370, 856, 896]. few [589]. Fibonacci [913, 97, 541, 535]. Fibonacci-like [541]. Field [1012, 1070, 778]. fields [831, 266, 35, 676, 344, 612]. fifth [280, 518]. Fifty [729, 231, 740, 857, 1066]. Fifty-Eighth [857]. fifty-ninth [231, 740, 1066]. Fifty-Seventh [729]. figures [210]. filled [625]. filling [121]. Finance [821]. Financial [811]. Find [992, 485, 573]. Finding [55, 152, 26, 696]. Finite [859, 778, 884, 177, 638, 24, 235, 140, 242, 104, 112, 55, 488]. First [728, 826, 758, 134, 436, 457, 633]. First-class [134]. fit [187]. five [647]. Fixed [812, 374, 111, 687]. fixing [891]. flashlights [598]. flatness [62]. float [131]. floor [77]. fluxions [690]. focusing [393]. folding [491, 585, 530]. Fomin [689]. Fool [673]. Football [525]. Force [809, 292]. Forced [1036]. forces [372]. forcing [336, 332]. Ford [1030]. Form [947, 731, 302, 143, 594, 553]. forms [636, 239]. Formula [741, 944, 976, 1115, 307, 427, 512, 845, 467, 38, 696, 254, 906]. Foucault [519]. Four [1057, 480, 307, 458, 560, 673]. Four-Color [1057]. four-vertex [560]. Fourier [877, 794, 590, 452, 240, 539, 409, 785]. Fourier-Motzkin [785]. fractal [137]. Fractals [1086, 839]. Fraction [951, 111]. Fractional [438, 884]. fractions [135, 219, 502, 347, 83, 426, 124, 898]. Frame [379]. France [343]. Francois [809]. Frank [734]. fraud [253]. Fréchet [400]. Fred [584]. Free [576, 775, 880, 471, 470, 657]. French [937]. Fresh [526]. friend [604]. Fritsch [1057]. Frobenius [235]. Front [941, 973, 1001, 1032, 1060, 1089, 1121]. Fubini [309]. Function [1038, 767, 791, 1021, 1109, 627, 289, 454, 138, 512, 118, 507, 564, 450, 897, 107, 369, 368, 201, 674, 582, 906, 399]. Functional [75, 183]. functionals [909]. Functions [634, 841, 930, 1054, 1111, 981, 1115, 704, 626, 561, 652, 56, 205, 915, 286, 603, 77, 839, 167, 202, 218, 189, 151, 195, 206, 141, 134, 186, 529, 264, 713]. Fundamental [1033, 1063, 353, 488]. Further [979].

G [863, 462, 39]. G. [768]. gadfly [164]. Gale [732, 497, 595]. Gale-Ryser [595]. Galileo [800]. Galois [793, 288, 25, 371]. Galvin [584]. Game [872, 732, 77, 882, 324, 274, 525, 6, 587, 310, 19]. Games [922, 181]. gamma [627, 450, 107, 369, 368, 906]. Gandt [809]. Gardner [837]. Gauss [606]. Gaussian [799, 896]. Gear [83]. gears [419]. general [568, 677]. Generalization [856, 707, 227, 635, 223, 376, 584]. Generalizations [856, 529]. Generalized [722, 788, 798, 778, 1039, 274, 553, 107, 369, 368, 588, 211]. Generating [542, 884]. generation [123]. generator [335]. Generically [873]. genesis [605]. Genkin [689]. geometer [462]. Geometric [627, 763, 622, 568, 59, 601, 405, 543, 533, 80, 458]. geometric-mean [405]. geometrical [696]. geometries [100]. Geometry [809, 519, 852, 1037, 406, 655, 88, 490, 840, 89, 76, 442, 318, 510, 417, 782]. Georg [447]. George [904, 508]. Gerda [1057]. Gergonne [593]. Germain [384]. German [719, 516, 536]. Germany [329]. Gerstein [849]. Giants [265]. Gibbs [117]. Giblin [442]. Gilbert [744]. Gillman [900]. Gin [379, 290, 656, 572, 900, 771]. Giuga [567]. given [917]. gives [413]. Glance [752]. Gleason [572]. Glen [782]. Glimpses [655]. globally [687]. GM [1114]. Go [795]. Golay [642]. Goldbach [269]. Golden [854, 910]. gons [530]. González [877]. González-Velasco [877]. Good [802, 750, 479]. Goodbye [876]. Göttingen [337]. Gottschalk [1009]. Grade [865]. Grant [719]. granular [372]. Graph [324, 549, 834, 364, 458]. Graphical [777]. graphics [187]. Graphs [814, 56, 419, 661, 217]. Gravity [1085]. Great [977]. greatest [541, 696]. Greed [48]. Green [904, 1118, 508]. grid [683]. Grossman [709]. Group [1041, 870, 52, 247, 538, 547, 514, 633]. Groups [859, 880, 638, 216, 524, 133, 335, 892, 235, 542, 242, 645, 55, 488, 92, 650]. grow [174]. Growth [1008]. Grünbaum [462]. Guide [744]. Gung [379, 290, 572, 771, 656, 900]. Guy [745].

H [68]. H. [1007]. Hadamard [894]. Hadwiger [277]. Haimo [656]. Hairer [768]. Hajarnavis [970]. Half [867, 436, 457]. Half-Chessboard [867]. Halley [510, 866]. Halmos [466, 836]. Hamilton [871, 204]. Handbook [863]. Hanging [764]. Hanoi [431, 246]. Hard [1112, 801, 479]. Hardy [719, 319]. Harm [728]. Harmonic [1054, 884, 652, 603, 76]. Harmony [1094]. Hedlund [1009]. Hendrik [518]. Henkin [17]. her [391]. Hereditary [692]. Hermite [302]. Heron [658]. Hersh [938]. heuristic [319]. Hex [181]. hexagon [361]. hexahedra [475]. High [779, 253, 300]. High-Dimensional [779]. Higham [863]. higher [218]. Hijab [849]. Hilbert [654, 578, 416, 172, 308, 644]. Hilton [850]. Hipparchus [684]. Hirsch [808]. Hirst [317, 329, 337, 343, 351, 359]. Historical [688, 697, 719]. History [655, 1003, 490, 848, 768]. Hoffman [1028]. Hohn [851]. Holder [981, 494]. holes [355]. Holton [850]. Homomorphisms [789]. Hoop [1002, 664]. Hoops [980]. Hop [980]. Hopping [980, 1002, 664]. Horn [646]. horology [498]. Horton [745]. Hough [684]. hourglass [372]. Howard [1117]. Hu [379, 290, 572, 771, 656, 900]. Hubert [851]. Hugh [759]. human [677, 645]. hundred [630]. Hunting [858]. Hypatia [391]. Hyperbolic [318, 1037, 120, 538, 249, 417, 529]. Hyperbolicity [943]. hyperboloid [318]. hypergenerating [289, 118]. hypergeometric [420]. hyperplanes [838]. Hypocycloids [124]. hypothesis [398].

Ian [969]. icosahedron [575]. Idea [879]. Ideal [818, 665]. Ideals [750, 49]. Ideas [852]. Identities [765, 13, 300, 453, 261, 39]. Identity [993, 777, 1079, 1026, 995, 260, 366, 47, 326, 666, 386, 327, 455]. II [392, 329, 457, 719, 791, 819, 1063, 536]. III [337]. Ilia [689]. illuminable [555]. Illustrated [584]. Imaginary [997]. Impartial [882]. implicative [645]. Implies [742, 647, 455]. impossibility [687]. impossible [491, 532, 585, 144]. Improving [204, 229]. Incenters [1108]. Incircles [1098]. including [402]. inclusion [188, 145]. Incommensurability [813]. incongruent [5, 66]. Indecomposable [962]. indeed [905]. independence [902]. independent [107, 369, 368]. index [346]. India [746]. Indigenous [1086]. Induction [801]. inductive [533]. Inequalities [114, 1022, 714, 59, 163]. Inequality [1116, 966, 921, 622, 220, 325, 73, 675, 188, 405, 11, 15, 543, 494, 533, 345, 308, 425, 552, 570, 375, 191, 53, 415, 562, 455]. inequivalent [52]. inference [80]. Infinite [1006, 5, 66, 177, 255, 658, 628, 440, 236, 245, 619]. infinitesimal [697, 719]. infinitesimals [641]. infinity [561, 998]. influence [690]. influences [436, 457]. Information [1119, 1117]. Injective [523]. inner [303]. Input [1063]. Inscribed [919, 520]. Integer [475, 307, 842, 896]. integers [581, 829, 613, 149, 182, 636, 397]. Integrable [1038, 195]. Integral [104, 749, 454, 392, 624, 275, 281, 171, 136, 81]. Integrals [974, 979, 884, 894, 845, 2, 195, 911, 713, 70, 708]. integrate [264]. Integrating [1069]. Integration [651, 688, 785, 855, 130, 349, 378]. Interactive [851]. interior [374]. Intermediate [1008, 208]. international [669]. interpretation [568, 171]. Intersections [720, 230]. interval [72]. Intervals [506, 101, 404]. Intractable [882]. Introduction [915, 370, 849, 758, 733, 170, 655, 970]. Introductory [970]. invariant [515, 95]. Inverse [892, 616, 670, 553, 206]. inverse-similarity [670]. Inversion [788]. Invertible [781, 679]. Inverting [654]. Investing [978]. Involving [1082, 1026, 47]. irrational [347, 426]. Irrationality [142, 860, 105, 694]. Isogonal [316]. Isometries [414, 531, 702, 94]. Isoperimetric [946]. Isoptic [22]. issues [480]. Italian [793]. Italy [343]. Itenberg [689]. Iteration [626]. Itself [1081, 800]. IV [343].

J [379, 1007, 997, 863, 495, 1057, 849, 878, 408, 669, 39, 970]. J. [640, 669]. Jacobi [307, 777, 898]. Jacobson [1011, 70]. James [1117]. jazz [565]. Jean [850, 851]. jeep [497]. Jensen [345]. John [745, 793, 851, 889, 291]. Johnson [864]. Jones [408]. Jordan [354, 577, 143, 40, 213]. Jordan-Alexander [354]. Jordan-Schönflies [213]. Joseph [795]. Journeys [1028]. Joy [709, 710]. Jr. [1117, 808]. Juggling [418]. Julia [888]. June [1118].

Kac [902]. Kaluza [698]. Kantorovich [552]. Kaplansky [722, 214]. Keith [876]. Kelly [272]. Kempe [780]. Kepler [205, 362]. key [352]. Klamkin [22]. Knaster [686]. knight [367]. knockouts [834]. knot [632]. Knots [744, 515]. Know [1092, 607, 435]. Knuth [441]. Konhauser [795]. König [222]. Kotz [864]. Krantz [863]. Kreiss [383]. Kronecker [522]. Krylov [879]. Kummer [420, 550, 413]. Kuratowski [40].

L [1117, 864, 889, 640]. Lambert [694]. Lanczos [798, 1039]. Large [1055, 230, 891, 296, 51, 718, 473, 127]. Larry [849]. Lasker [439]. Last [673, 868, 370, 896]. late [495]. Latin [700, 666, 169]. lattice [225]. Laura [793]. Law [952, 1004, 818, 473, 127]. laws [362]. Lawvere [758]. Leading [864]. Leads [722]. Leaning [998]. Learning [826, 727]. Lebesgue [749]. Lecturing [701]. Left [1003]. Legacy [760]. Legendre [787]. Lehmer [313, 68]. Lemma [722, 886, 774, 1094, 918, 194, 505, 147, 212]. Lemoine [119]. length [573]. lengths [377]. Lengthy [752]. lenses [249]. Leon [17]. Leonard [900]. Lester [1030]. Letter [50]. Level [778]. Lexell [919]. L’Hôpital [75, 130]. Life [698, 888, 710, 969]. like [541]. Limit [967, 994, 517]. Lindemann [36]. Lindemann-Weierstrass [36]. line [454, 187, 687]. Linear [736, 882, 173, 193, 836, 27, 117, 156, 288, 78, 649, 353]. Linearization [1063]. Lines [187, 259, 1051, 634]. link [40]. links [515].

Liouville [198, 148]. Lite [734]. literal [899]. little [298]. Littlewood [640, 319]. Liu [1117]. Local [1062, 469]. logarithm [312]. Logarithmic [1084, 254]. logic [182]. logical [717]. Logistic [943]. London [351]. long [681]. look [380, 462, 402, 663]. looking [918]. Lost [748]. Loved [1028]. low [204]. low-rank [204]. lowbrows [340]. Lowell [729, 857, 1066]. Löwner [206]. Lucas [37, 313, 223]. Lucas-Lehmer [313]. Luxembourg [623]. Lynn [209].

M [837]. M. [640, 22]. Maclaurin [690, 1113]. Magic [905, 27, 916, 156, 945, 401]. Magical [969]. magnetism [508]. Major [221, 693, 436, 457, 612]. make [839, 330]. Makes [977]. making [910]. Malgrange [159]. Malgrange-Ehrenpreis [159]. Man [1028]. Mandelbrot [913]. Many [685, 850, 712, 248, 388]. map [551]. mapping [251, 183, 103]. mappings [151, 687]. Maps [773, 943, 181, 434, 523]. Marie [851]. Mark [902]. market [272]. Markov [352]. Marriage [945]. Marriages [814]. Martin [837]. marvelous [389]. Marvin [291]. Mascheroni [444]. master [648]. Match [872]. matching [459]. Matchings [932]. material [372]. Math [66, 67, 65, 441, 740, 462, 495, 382, 68, 69, 408, 442, 39, 837, 70]. Mathematica [688, 821]. Mathematical [729, 850, 609, 863, 857, 1066, 849, 291, 1028, 837, 689, 770, 795, 642, 717, 623, 240, 362, 267, 640, 912, 498, 516, 536]. Mathematician [1034, 643, 904, 317, 329, 337, 343, 351, 359, 677, 937]. Mathematicians [402, 525, 903]. Mathematics [783, 1090, 735, 710, 823, 938, 843, 758, 953, 865, 615, 819, 977, 786, 753, 391, 623, 901, 606, 84, 565, 526, 463, 436, 457, 697, 719, 677, 557, 743, 565, 888, 699, 744, 938, 890, 769]. Mathieu [524]. Matrices [1050, 1097, 1078, 750, 679, 185, 314, 197, 116, 577, 670, 123, 692, 534, 26, 95, 438, 551, 415]. Matrix [522, 1022, 867, 949, 464, 63, 662, 163, 844, 57, 331, 383]. Matter [940, 972, 1000, 1031, 1059, 1088, 1120, 941, 973, 1001, 1032, 1060, 1089, 1121]. max [94, 74, 365]. max-min [74]. Maxima [885]. Maximization [1052]. maximum [331]. May [873]. Maze [969]. McDill [851]. McPeak [794]. Mean [873, 921, 979, 1010, 622, 72, 188, 2, 405, 543, 533, 128, 672, 671, 70, 708, 751, 950]. mean-geometric [622]. Means [788, 81]. measure [428, 121]. measures [628]. Mechanical [852]. mechanics [648, 709]. mechanisms [558]. medalist [612]. Medial [1099, 933]. medians [658]. Medium [616]. Mercedes [632]. Merton [812]. Method [785, 4, 67, 499, 286, 696, 179, 510, 263, 584]. Methods [879, 866, 571]. Metric [723, 120, 94, 915]. metric-preserving [915]. middle [8, 86]. Milin [909]. min [74]. Mind [890]. mini [490, 589]. mini-essays [589]. mini-history [490]. Minimal [237, 90, 676, 322]. minimalist [565]. minimization [586]. Minimum [800, 908]. Minkowski [494]. Minsk [837]. Mira [421]. Miranda [706]. mirror [828]. mirrors [249, 850]. Mixed [921, 858, 405, 533]. Mixed-Mean [921]. Mixing [103]. Mixtilinear [1098]. Mixtures [224]. Möbius [348, 507]. mod [412]. Model [1037]. Modeling [811, 821]. Models [861, 100, 781]. Modern [1061, 725, 1086, 615]. modified [499, 258]. Modular [1041]. modules [667]. Modulo [755, 493, 695, 124]. Moebius [102]. Mohr [444]. Mohr-Mascheroni [444]. moment [178]. Moments [961]. Monograph [902]. Monotonicity [885]. Monthly [66, 67, 65, 441, 740, 563, 462, 495, 382, 68, 69, 408, 757, 442, 39, 70, 284, 291, 703]. Moore [499]. Mordell [287, 653]. Morgan [977, 734]. Morphisms [431]. Mosaic [770].

Moscow [655]. most [452]. motion [332, 553]. Motivated [979, 13, 39]. Motzkin [785]. MR [655, 66, 67, 65, 623, 441, 740, 462, 495, 382, 68, 69, 442, 39, 70]. multidimensional [492]. Multiple [158, 974, 713]. multiplicative [507]. Multivariable [1035]. Multivision [882]. Musings [54]. My [544, 1028]. mylar [461].

N. [744]. Nahin [997]. Napierian [312]. Napoleon [234]. Nathaniel [709]. National [1062]. Nauka [655]. Need [1092, 901, 808]. Needham [835]. needle [383]. negative [827]. Negatively [963]. Nelson [1067]. Nested [176]. nets [273]. Neugebauer [381]. Never [495]. Newman [721]. Newton [809, 382, 286, 690, 362, 261, 292, 648, 243, 866]. Newtonian [28]. Nicholas [863]. Niels [518]. Nim [191]. ninth [231, 740, 1066]. No [728, 66, 67, 65, 441, 740, 462, 382, 68, 69, 408, 442, 39, 70]. Nobel [557]. nod [392]. Noether [439]. Non [859, 1078, 778, 454, 827, 198, 842, 440, 598]. non-constant [454]. Non-Euclidean [598]. Non-existence [859]. non-integer [842]. non-Liouville [198]. non-negative [827]. Non-Square [1078]. non-unique [440]. none [619]. Nonlinear [1063]. nonmeasurable [226]. nonnegativity [433]. nonperiodicity [424]. nonsplittable [350]. Nonstandard [848]. nontransitive [410]. nonunique [275]. Nonzero [860]. noodles [446, 422]. Norm [163, 57, 414, 21, 455]. norm-inequality [455]. normal [476, 72, 302, 902]. Norman [864]. norms [63]. notation [441, 238]. Note [859, 979, 803, 861, 293, 152, 309, 675, 585, 686, 487, 136, 596, 426, 326, 708, 898]. Notebook [748]. Notes [718, 765, 730, 711, 731, 441, 238]. Nothing [887]. novice [459]. nowhere [151, 201]. nowhere-differentiable [151]. Number [947, 874, 137, 887, 815, 890, 164, 1008, 804, 1113, 789, 721, 307, 142, 558, 158, 182, 467, 388, 198, 514, 460, 458, 503]. Numbers [1096, 1003, 710, 962, 1024, 1028, 749, 296, 630, 276, 154, 262, 51, 31, 902, 447, 388, 97, 279, 473, 127, 64, 535, 661, 165, 745]. numerators [31]. Numerical [883].

O [1117, 290]. Objects [875]. Observability [1033]. obvious [194]. occurrence [165]. oceans [198]. Odd [401]. ODE [818]. old [604, 74, 434, 49, 891, 560]. oldest [448]. Omar [849]. One [773, 1112, 532, 607, 124, 34]. One-Dimensional [773]. one-sentence [34]. Only [1028]. opaque [271]. Open [1028]. operation [278]. Operator [736, 165]. Operators [788, 692, 81]. Optimal [274]. optimization [571]. orbits [301]. Order [509, 224, 247, 140, 538, 259, 514, 80, 81]. order-restricted [80]. ordered [831]. Orderly [373, 408]. Ordinary [807]. origami [491]. Origin [1065]. orthogonal [522, 670, 123, 146]. oscillations [891]. oscillators [76]. oscillatory [332]. Other [969, 132, 35]. Otto [381]. our [436, 457]. outer [315]. Overview [267].

P [442]. Packability [647]. packings [360]. Paper [1056, 491, 274, 585]. paperfolding [893]. papers [838]. Parabolic [249]. Paradox [764, 711, 407, 410]. paradoxes [463]. Parallel [834, 784]. Parallelogram [1099, 855]. parallels [576]. Parameters [1096]. Parametric [139]. Part [120, 819, 767, 791]. Partial [858, 812, 135]. partials [469]. parties [657]. Partition [983]. partitioning [268]. Partitions [724, 772, 668, 503]. Pascal [740, 37, 314, 231, 246, 361]. Path [800, 486]. Pathological [286]. Paths [802, 870]. Patterns [770]. Paul [997, 938, 1028, 783, 836]. pay [680]. Pebbling [482]. pedal [58, 103]. Pedersen [850]. Pendulum [1036, 332, 519]. Percy [612]. Perfect [932, 475]. perimeter [125]. Period [251]. Periodic [930, 244, 399]. Periodicity [650]. Periods [922, 994]. permanent [427]. permanents [166]. permutahedron [53]. Permutations [513, 111]. Personalities [864]. Perspective [725, 840]. Peschke [1057]. Peter [970, 850]. Peters [1117]. Petsinis [937]. phenomenon [117, 500]. phi [398]. Philosophy [938, 526]. Physically [979]. Pi [1042, 229]. Picard [202]. pick [182, 220, 512, 342, 295]. Pick-type [342]. pieces [600]. piecewise [406, 117, 442]. piecewise-linear [117]. pigeonhole [434, 214]. pills [108, 200]. Pinhole [840]. Pizza [398]. Place [786]. planar [419]. planarity [40]. Plane [975, 931, 454, 464, 445, 100, 140, 538, 128, 530, 294, 531]. planes [451]. planet [843]. Platonic [68, 3]. Pleasures [822]. Plus [1035]. Plutarch [815, 684]. Poincaré [706, 1037, 1118]. Poincaré-Miranda [706]. Point [812, 374, 346, 208, 458, 555]. points [485, 225, 613, 111, 305, 109, 199, 112, 687, 477, 399]. Poisson [75, 195, 375]. Poker [868]. Polar [304]. Pollak [290]. Pólya [574, 1053]. Polygon [787, 119, 600]. Polygonal [870, 555, 388]. polygons [244, 342, 520, 409, 341, 113]. polyhedra [380, 462, 663, 386]. Polyhedron [781, 785]. polylogarithmic [1122]. Polynomial [358, 1113, 881, 1115, 485, 568, 228, 618, 625, 21, 523, 322]. Polynomialrecurrencology [678]. Polynomials [787, 961, 51, 777, 968, 981, 1069, 502, 305, 99, 363, 827, 472, 676, 433, 341, 306]. Polytope [631]. Polytopes [881, 631]. Pompeiu [320, 387]. pools [525]. popular [467]. Porter [744]. Poset [1081]. Positive [788, 116, 613, 121, 413]. possible [427, 490]. Post [352]. Post-Markov [352]. Postcards [365]. posters [85]. Power [700, 710, 590]. Powers [934, 581, 33, 604, 839, 163, 401].

Practices [769]. precession [521]. precision [253]. prescribed [377]. present [61]. presents [166]. Preserve [750]. preserving [915, 186]. primality [567]. primary [439]. Prime [705, 778, 296, 493, 630, 51, 54, 34, 721]. Prime-producing [705]. Primer [1117, 863, 903, 1117]. Primes [799, 752, 895, 527, 248, 285, 548, 636, 239, 504]. Primitive [976]. Principal [665, 714, 35, 344]. Principia [809]. principle [348, 603, 434]. Principles [769]. prisoners [680]. Prize [557]. Pro [726]. Probabilistic [683, 668]. probabilities [476]. Probability [872, 762, 558, 623, 628, 458, 448]. Problem [926, 925, 1064, 927, 616, 856, 946, 928, 279, 278, 367, 924, 836, 923, 929, 774, 1067, 770, 797, 289, 452, 271, 466, 227, 74, 118, 162, 18, 670, 629, 497, 501, 6, 172, 632, 366, 8, 12, 10, 11, 106, 86, 16, 14, 15, 47, 107, 108, 109, 200, 199, 369, 368, 320, 42, 310, 912, 644, 429, 28, 458, 908, 448, 250, 387, 236, 269, 383, 498, 150, 1118]. Problems [959, 988, 1077, 1074, 1107, 792, 806, 820, 1103, 1016, 1044, 732, 985, 1020, 1101, 700, 954, 1019, 1076, 1045, 1072, 1017, 1105, 1046, 1015, 987, 958, 1047, 955, 1102, 990, 757, 1100, 991, 989, 1013, 1049, 1104, 1073, 1075, 1071, 1048, 1043, 957, 1018, 1014, 960, 795, 986, 1106, 956, 352, 556, 563, 428, 479, 537, 703, 877]. processes [174]. producing [705]. Product [880, 453, 197, 303, 628]. Product-Free [880]. Products [1054, 1042, 1006, 522, 190, 894, 215, 161, 47, 513, 539, 245]. program [660]. progress [84]. Progressions [1080, 283, 674]. Projecting [982]. projection [631]. projections [654]. projective [840, 140]. Proof [755, 942, 983, 172, 405, 886, 790, 884, 721, 622, 13, 307, 287, 36, 194, 325, 222, 313, 578, 524, 273, 845, 354, 609, 403, 183, 363, 577, 389, 416, 444, 198, 653, 595, 694, 528, 586, 147, 155, 136, 494, 533, 356, 350, 22, 308, 349, 39, 159, 214, 319, 628, 906, 32, 907, 646, 361, 34, 584].

Proofs [786, 914, 849]. properties

[627, 59, 245]. Property

[920, 423, 374, 393, 279, 223, 323, 103].

proportional [634]. protocol [471]. Prove [783, 127]. Proved [800]. Proving [786, 831]. public [352]. public-key [352]. punk [565]. Pure [1034]. pursuits [683]. Putnam [729, 857, 1066, 485, 227]. Puzzle [1061, 431, 29]. Puzzles [837]. pyramid [29]. Pythagorean [158, 588, 1037].

Quadratic

[1003, 730, 266, 636, 239, 32, 399]. Quadratics [527, 705]. Quadrilateral [1085]. quantum [173]. Quarles [731]. quartic [568]. quasiperiodicity [650]. quaternion [626]. Quaternions [776]. queens [429]. Queenwood [343]. Question [1005, 385]. questions [504]. Quick [871]. quicker [321]. Quine [738]. quintics [465]. Quotients [285, 504].

R [1030, 836, 39, 970]. Rabinowitz [795]. Radicals [746, 176]. radius [57]. Ramanujan [39, 13, 748, 340, 746, 755, 61, 765, 983, 326, 747]. Random [975, 592, 101, 968, 932, 44, 390, 278, 539, 294, 437]. randomness [608]. rank [204]. Rankin [907]. ranks [185, 534]. rapid [1122]. Rates [861]. Ratio [1084, 613, 283]. Rational [813, 860, 399, 244, 658, 594, 897, 123, 264]. Rationals [1041, 182]. Ratios [965]. Ravi [770]. Raymond [1117]. Reading [648, 580]. Real [772, 968, 491, 305, 670, 228]. Really [938, 313, 619]. rearrangement [53]. reasoning [717, 645]. rebel [639]. Reciprocal [993, 175]. Reciprocity [952, 1003, 32]. rectangles [339, 212]. Rectangular [679]. rectilinear [458]. Recurrence [1112, 967, 294, 12, 16, 537]. Recurrent [1021]. recursive [172]. redefined [841]. Reflected [1051]. reflection [828]. Reflections [850, 421, 828, 394, 372]. Reform [1062, 754, 1091, 753]. Regions [464]. Regular [859, 297, 660, 870, 638, 445, 341]. regularity [80]. Reid [888]. related [37, 466]. relating [514]. relation [216, 503]. Relations [1112, 82, 661]. relationship [362, 620]. relatives [277, 202]. reliability [609]. Remainder [46, 392, 38]. remark [72]. Remarks [739, 196]. Remez [325]. reminiscences [381]. removal [594]. Rental [1094]. repeatedly [175]. Replacement [861]. Replication [203]. reply [39]. Representation [756, 949, 41, 104]. representations [307, 293, 829]. Representing [79, 239, 527]. reputation [669]. Require [996]. Research [688, 642]. Resistive [616]. Resnikoff [1117]. Resolution [878]. respectable [417]. respond [462]. response [462]. restricted [80]. result [363]. Resultant [722]. results [37]. Rethinking [727, 671, 672]. Return [624]. Reuben [938]. Reverend [612]. reverse [425]. reversing [892]. Review [1117, 998, 997, 938, 1118, 1057, 1028, 1086, 937, 969, 970]. Reviews [698, 745, 699, 709, 744, 710, 782, 688, 733, 796, 810, 824, 939, 971, 999, 1029, 1058, 1087, 781, 769, 734, 689, 770, 768, 795, 794]. revised [837]. Revisited [816, 780, 1002, 1053, 685, 139, 129, 73, 74, 847, 97, 891, 439, 560]. revolution [576]. Rewriteability [242]. Richard [745, 794]. Riemann [624, 398, 183, 195, 953]. Rigatelli [793]. rigor [623, 672, 671]. Ring [789, 605, 269]. rings [844, 350, 474, 43, 440]. rise [490]. road [625]. roads [516, 536]. Robert [837]. Rogers [39, 13]. role [546, 463, 672, 671]. Rolle [270, 579, 492]. Roman [575, 698]. Romantic [565]. Romberg [349]. Ron [1086]. Ronald [688]. Rook [867]. Room [850]. rooms [555]. root [358, 4, 67]. Rooted [1026]. Roots [961, 968, 866, 897, 228, 412]. rose [125, 837]. Rota [666]. Roth [594]. Rouché [99]. Round [868, 855, 497]. round-trip [497]. row [231, 740]. Rudin [642]. Rudolf [1057]. Rule [816, 1068, 1078, 75, 130, 594, 910]. Russian [655, 689, 899]. Ryser [595].

S [1117, 1007, 22]. Saber [733].

saddlepoint [152]. same [483]. sampler [608]. Samuel [864]. Sandra [794]. Scalable [1117]. scenes [437]. Schafer [771]. Schanuel [758]. Schechter [1028]. Scheduling [657]. Schönflies [213]. School [786, 300]. Schools [808]. Schröder [684]. Schur [790]. Schwarz [220, 73, 212]. Schwarz-Pick [220]. science [617]. sciences [267, 863, 864]. scientist [639]. Scientists [1117]. scissors [274]. sculptures [85]. Search [775, 1028, 428, 140]. Searching [529]. Secant [1069, 472, 263]. Second [815, 826, 469]. second-partials [469]. Secret [969]. secrets [652]. sections [445]. seeing [283]. seen [641]. Self [580, 573, 217]. self-dual [217]. Self-reading [580]. self-similarity [573]. Semigroups [1008]. semisimple [667]. Sense [890]. sensitive [569]. sentence [34]. Separable [60]. Separating [838]. Separation [761, 306]. Sequence [1021, 642, 913, 110]. Sequences [296, 248, 423, 396, 97, 58, 54, 541, 537, 580, 681]. Sequential [268]. Sergey [689]. Series [763, 811, 832, 936, 884, 590, 135, 392, 253, 452, 255, 619, 154, 41, 240, 409, 413]. Service [379, 572, 771, 290, 209, 656, 657, 17, 900]. Set [869, 94, 700, 854, 658, 913, 917, 912, 668, 298]. Sets [930, 943, 613, 621, 336, 230, 430, 506]. Seven [883]. Seventh [729]. seventy [280]. seventy-fifth [280]. Shannon [587]. shape [461]. Shapiro [642]. share [677]. sharing [448]. Sharkovsky [599, 739]. Sheer [709]. shells [129]. Shenitzer [655, 899, 719, 743, 516, 536]. Shephard [462, 462]. Sherman [710]. Short [983, 721, 307, 287, 363, 444, 653, 486, 349, 681].

Shorter [755]. shortest [72, 486]. Should [753]. side [102]. Sierpiński [432]. significance [677]. significant [452, 500, 480]. significant-digit [500]. Signs [816, 1068, 963]. Silo [766]. Similarities [590]. similarity [670, 573]. Simple [741, 805, 886, 790, 275, 273, 712, 354, 183, 171, 577, 695, 595, 147, 155, 538, 494, 26, 319, 294, 907, 157, 298, 361]. simplices [297]. Simplicity [748, 524, 119]. Simplifying [356]. Simson [982]. Simultaneous [1080]. Simultaneously [704]. Sine [935]. SING [556]. SING-SING [556]. Single [1063, 645]. Singular [71, 23]. Sins [883]. Six [918, 647]. sleeping [233]. slicing [398]. slowed [84]. small [108, 200]. Smallest [948]. Smith [1007]. smooth [829]. Smoothness [112]. Snyder [781]. soap [402]. social [267, 938]. sociology [606]. Soddy [593]. software [851]. solids [419, 68, 3]. Solitaire [945, 6, 310]. Solution [289, 366, 86, 16, 14, 15, 47, 200, 199, 369, 368, 948, 310, 250, 568, 81]. Solutions [792, 806, 820, 795, 833, 235, 487, 479, 420]. solvable [465]. Solving [760, 770]. Some [558, 215, 196, 714, 996, 100, 865, 537, 668, 1033, 1063, 245, 37, 619, 347, 426, 480, 535, 157]. sometimes [48]. Somewhat [752]. Sorting [784]. sound [483]. sources [440]. space [654, 121]. space-filling [121]. Spaces [819, 723, 702, 303, 168, 9, 69, 297, 743]. specialization [584]. specified [189]. spectral [71, 57]. spectrum [611]. Sperner [1094]. sphere [647, 18, 55, 583]. spheres [647, 651]. spherical [828, 566]. spigot [489]. spike [104]. spline [180]. Splitting [922, 600, 174]. sprouts [324, 659]. Square [1078, 412, 29, 416, 897]. squarefree [431]. Squares [411, 1007, 1080, 307, 905, 27, 262, 79, 156, 161, 666, 169, 34]. Squaring [355]. Stability [59, 1040, 95]. stacking [203]. Stalking [545]. Stan [795]. Stanford [744].

Stanislas [890]. Stanley [795]. starlike [206]. state [617]. Statistical [902, 864, 80]. Statistics [735, 224]. stay [456]. Stedman [633]. Steen [209]. Stefan [698]. Stein [710, 236]. stepping [126]. stereogram [437]. Sterrett [699]. Steve [1117]. Steven [851, 863, 758, 878]. Stewart [969, 969]. Stirling [154, 1084, 279, 96]. stochastic [450, 547]. stock [272]. stockbroker [233]. Stolarsky [425]. stone [274]. stones [126]. Story [997, 1028, 548]. Strang [210]. Strange [253, 210]. Strategies [587, 274]. Strategy [882]. Strength [710]. strengthening [220]. strict [152]. strikes [334, 30]. String [459]. strings [431]. strip [102, 304]. Strogatz [851]. Stroll [799]. Strong [774, 127]. Strongly [1079]. Structure [1117, 717, 146]. Structures [849]. Student [329]. Study [865, 642]. Subdifferential [803]. subgroups [226]. submatrices [714]. subordinate [63]. Subset [732]. Subsets [880]. Subspaces [761, 559, 95, 122]. successive [634]. Sue [998]. sufficient [228]. Sum [964, 307, 411, 619, 388, 908, 34]. summability [416]. Summation [944, 993, 10, 14]. Sums [581, 952, 996, 1027, 255, 262, 827, 79, 161, 106, 47, 539, 311, 250]. superlinear [263]. Supremum [935]. surface [576, 597]. Surfaces [744, 946, 775, 237, 90, 213, 744]. Surjectivity [869]. Surprise [748, 764, 367]. surveyor [317]. suspension [891]. Sutherland [379]. Swiftest [800]. switching [587]. Sylvester [116, 669, 458]. symbols [898]. symmetric [704, 542, 433, 713]. Symmetries [315, 55]. Symmetry [858, 141, 348, 892]. Synthesis [1091]. Systems [1063, 276, 177, 77, 137, 480]. Szeged [299]. Szego [1053].

T [442]. T. [744]. Table [984]. tables [169]. Tagged [772]. Take [732]. Take-Away [732]. Taking [922]. Tale [997, 544, 911]. Tangent [1055, 1069, 634, 472]. Tanks [1095]. Tantrices [566]. Tarski [300]. Tartaglia [616]. Taxicabs [747, 311]. Taylor [46, 878, 208, 38]. teach [252, 499]. Teacher [977]. Teaching [735, 826, 137, 78, 688]. TEAM [978]. technique [127]. technology [615]. Telegraphic [796, 810, 824, 939, 971, 999, 1029, 1058, 1087]. Temple [382]. tennis [596]. tenth [172]. Tepper [656]. Tessellations [241]. test [313, 469, 550, 413]. Tetrahedra [976, 338, 445]. Tetrahedron [1099]. Texts [807]. th [934]. Their [964, 94]. them [480]. theme [485, 560]. Theodore [744]. Theorem [873, 1007, 919, 1009, 790, 739, 1040, 979, 1010, 731, 1037, 884, 432, 549, 468, 352, 287, 207, 651, 707, 36, 98, 37, 222, 578, 273, 24, 370, 309, 354, 599, 270, 579, 99, 46, 257, 35, 188, 492, 403, 183, 202, 31, 577, 148, 295, 380, 718, 395, 462, 71, 444, 235, 2, 653, 595, 706, 844, 206, 155, 136, 596, 208, 376, 356, 344, 663, 588, 582, 159, 128, 214, 353, 488, 32, 907, 672, 560, 178, 40, 213, 671, 650, 646, 383, 234, 70, 361, 517, 708, 896, 751, 673, 950, 721, 982, 1057]. Theorems [874, 221, 85, 23]. theorist [633]. Theory [737, 1003, 887, 1012, 1070, 812, 1008, 1033, 1063, 142, 558, 324, 182, 203, 288, 25, 78, 164, 460, 165, 448, 371]. There [700]. thin [621]. things [131]. Thinking [727, 719]. Thomas [317, 329, 337, 343, 351, 359]. Three [975, 556, 652, 1083, 1118, 765, 428, 161, 141, 912]. three-door [912]. Thurston [759]. Tie [872]. tiled [339]. tiling [92]. tilings [180]. Time [811, 800, 619]. TIMSS [865]. Toetjes [93]. Tom [937]. too [495]. tool [137]. Tools [1090]. topological [451, 192, 9, 69, 128]. topologies [167, 49]. topologist [462]. Topology [782, 1035, 516, 536, 137, 192]. Toricelli [818]. Torricelli [856]. torsional [891].

Totally [491]. Toti [793]. Totient [1082]. tour [89]. tournament [185, 534]. Tower [431, 246]. trace [415]. Tracking [1108]. Tracy [382]. Tradition [1091]. trains [83]. transcendental [447, 198]. transcendentals [716]. transfiguration [490]. Transform [794, 416]. transformations [348, 146, 204]. Transforming [530]. transforms [539]. Transitive [569]. transitivity [546, 404]. Translated [793, 516, 536, 809, 655, 899, 719, 743]. transmutation [292]. transpositions [513]. trapezoidal [611]. trapezoids [610]. trapped [394]. traversal [396]. treatise [690]. Treatment [1061]. tree [913, 126]. Trees [1026, 434]. trials [460]. Triangle [966, 933, 490, 231, 740, 246, 360, 377, 593]. Triangles [225, 1083, 723, 338, 37, 658, 592, 445, 443, 544, 693, 58, 44]. Triangular [874, 262, 715]. Triangulating [390]. tricks [434]. Trigonometric [894, 827, 438]. Trio [874, 730]. trip [497]. triples [158]. Tristan [835]. trivial [313]. Truth [1028]. Tuples [964]. Turán [549]. Turing [517]. Turner [771]. twist [181]. Twists [747]. Two [1112, 1096, 1007, 441, 85, 839, 1083, 202, 715, 826, 765, 161, 238, 23, 364, 778, 658, 197, 305, 600, 544, 106, 407, 311, 504, 250, 560, 911, 516, 536, 34]. Tychonoff [273]. type [342, 912]. types [544].

Ulam [742]. undergraduates [402, 583]. undetermined [179, 584]. Unexpected [764]. unified [221]. Uniform [1004]. uniformization [212]. Uniformly [1096]. Union [1042]. Unions [122]. Unique [842, 440]. uniqueness [247, 488]. Unit [1067]. unitary [415]. Unity [724]. Univalent [827]. Universal [869, 396, 9, 69]. Universally [226]. Unsettled [1116]. unsolvability [172]. Unsolved [732, 700, 757, 1100, 563]. unstable [687]. Urn [861]. Use [772, 570]. Using [573, 992].

V [351]. Vakil [770]. Value [873, 751, 950, 877, 979, 1010, 188, 2, 128, 672, 671, 70, 708]. Values [1055, 296, 51, 71, 23]. Vandermonde [334]. vanish [561]. Variable [861, 845]. Variables [974, 305, 141, 28]. Variation [1025, 1114, 485, 218]. Variations [468, 436, 457, 560]. various [1122]. vector [63]. Vectors [963]. Velasco [877]. Velleman [795]. velocity [903]. Versa [826]. Versatile [282, 366]. version [492, 183]. vertex [435, 560]. vertices [225]. Very [667, 159]. VI [359]. via [952, 558, 273, 512, 130, 49, 919, 28, 455]. Vice [826]. Videotape [865]. Vieta [1042]. View [944]. visit [28]. Visitor [795]. Visual [835, 345]. Visualizing [149]. Vita [688]. Volume [1119, 761, 588]. Volumes [637]. vs [400].

W [744, 970]. Wagon [795]. Wald [187]. walk [294]. Walker [1117]. Walks [975]. Wallace [982]. Wallet [711]. Wallis [1042]. Walls [992]. Wanner [768]. want [607]. Was [800, 690]. wave [150]. Wavelet [1117]. Wavelets [1117, 330, 1117, 1117]. Way [1092, 801, 795, 606, 435]. ways [918]. weak [196]. Weatherman [1092]. Weierstrass [36, 512, 718, 155]. Weighted [921]. Wells [1117]. West [851]. Weyl [675]. Weyr [1093]. wheels [498]. Where [485, 385]. Which [445, 805, 1095, 1092, 723, 169, 795, 4, 67, 561, 532]. whirlwind [89]. Who [1028]. Whose [964, 1065, 454, 634, 839, 495, 1]. Wielandt [582]. wild [545]. William [729, 857, 1066, 758]. Williamson [495]. Wilson [809]. Wind [1092]. Winning [882]. without [142, 424, 259]. Wolstenholme [468, 707, 847]. Woolfe [998]. Work [800]. Workable [419]. works [179]. Workshop [794]. Worlds [889]. worth [470]. Would [889]. Would-be [889]. write [508, 528]. Writing [863, 863]. wrong [449]. Wythoff [191].

xtravagant [317, 329, 337, 343, 351, 359].

years [284, 630, 738, 359]. Yorkshire [317]. Yueh [379, 290, 656, 572, 900, 771]. Yueh-Gin [379, 290, 656, 572, 900, 771]. Yuri [837].

Zaphod [740, 231]. zero [454, 602]. Zeros [1113, 198, 160, 306]. Zeta [1021]. ZIP [942, 747]. zones [323]. Zonohedra [211]. Zorn [147].

References

Williamson:1946:DWE

[1] J. Williamson. Determinants whose elements are 0 and 1. American Mathematical Monthly, 53(??):427–434, ???? 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See addendum [495].

Jacobson:1982:MVT

[2] Bernard Jacobson. On the mean value theorem for integrals. American Mathematical Monthly, 89(5):300–301, ???? 1982. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See errata [70].

Lehmer:1986:CPS

[3] D. H. Lehmer. Coloring the Platonic solids. American Mathematical Monthly, 93(4):288–292, ???? 1986. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See errata [68].

Benjamin:1987:BMW

[4] Arthur Benjamin. The bisection method: which root. American Mathematical Monthly, 94(9):861–863, ???? 1987. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See errata [67].

Beebee:1988:EII

[5] John Beebee. Examples of infinite, incongruent exact covers. American Mathematical Monthly, 95(2):121– 123, ???? 1988. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See errata [66].

Hayes:1988:EPE

[6] Barry Hayes, Donald E. Knuth, and Carlos Subi. Elementary problem E3267, a solitaire game. American Mathematical Monthly, 95(5):456–457, May 1988. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [310].

Boyd:1988:DE

[7] David W. Boyd. The Diophantine equation x^{2} + y^{m} = z^{2n}. American Mathematical Monthly, 95(6):544– 547, ???? 1988. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See addendum [65].

Knuth:1988:APA

[8] Donald E. Knuth and Ilan Vardi. Advanced problem 6581, the asymptotic expansion of the middle binomial coefficient. American Mathematical Monthly, 95(8):774, October 1988. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [86].

Magill:1988:UTS

[9] K. D. Magill, Jr. Universal topological spaces. American Mathematical Monthly, 95(10):942–946, ???? 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See errata [69].

Knuth:1989:EPEb

[10] Donald E. Knuth. Elementary problem E3303, a binary summation. American Mathematical Monthly, 96(??):54, 1989. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [14].

Knuth:1989:EPEc

[11] Donald E. Knuth. Elementary problem E3309, a binomial coefficient inequality. American Mathematical Monthly, 96(??):154, 1989. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [15].

Knuth:1989:EPEa

[12] Donald E. Knuth. Elementary problem E3335, a deranged recurrence. American Mathematical Monthly, 96(??):525, 1989. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [16].

Andrews:1989:MPR

[13] George E. Andrews and R. J. Baxter. A motivated proof of the Rogers– Ramanujan identities. American Mathematical Monthly, 96(5):401–409, ???? 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See comment [39].

Knuth:1990:SEPb

[14] Donald E. Knuth. Solution to elementary problem E3303, a binary summation. American Mathematical Monthly, 97(??):348–349, 1990. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [10].

Knuth:1990:SEPc

[15] Donald E. Knuth. Solution to elementary problem E3309, a binomial coefficient inequality. American Mathematical Monthly, 97(??):614, 1990. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [11].

Knuth:1990:SEPa

[16] Donald E. Knuth. Solution to elementary problem E3335, a deranged recurrence. American Mathematical Monthly, 97(??):927, 1990. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [12].

Alder:1990:ADS

[17] Henry L. Alder and Ivan Niven. Award for Distinguished Service to Leon Henkin. American Mathematical Monthly, 97(1):3–4, January 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Giblin:1990:BSP

[18] Peter J. Giblin and Donal B. O’Shea. The bitangent sphere problem. American Mathematical Monthly, 97(1):5– 23, January 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sutner:1990:GCA

[19] Klaus Sutner. The σ-game and cellular automata. American Mathematical Monthly, 97(1):24–34, January 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Weinstock:1990:EE

[20] Robert Weinstock. Elementary evaluations of ∫_{0}^{∞} e^{−x^{2}} dx and ∫_{0}^{∞} cos x^{2} dx, and ∫_{0}^{∞} sin x^{2} dx. American Mathematical Monthly, 97(1):39–42, January 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Newman:1990:NPC

[21] Donald J. Newman and J. S. Byrnes. The L^{4} norm of a polynomial with coefficients ±1. American Mathematical Monthly, 97(1):42–45, January 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nitsche:1990:ICC

[22] Johannes C. C. Nitsche. Isoptic characterization of a circle (proof of a conjecture of M. S. Klamkin). American Mathematical Monthly, 97(1):45– 47, January 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mathias:1990:TTS

[23] Roy Mathias. Two theorems on singular values and eigenvalues. American Mathematical Monthly, 97(1):47– 50, January 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Coloma:1990:DTF

[24] Teresa L. Santa Coloma and Ralph P. Tucci. A decomposition theorem for finite automata. American Mathematical Monthly, 97(1):51–54, January 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hungerford:1990:CGT

[25] Thomas W. Hungerford. A counterexample in Galois theory. American Mathematical Monthly, 97(1):54– 57, January 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Newton:1990:SAF

[26] Tyre A. Newton. A simple algorithm for finding eigenvalues and eigenvectors for 2 × 2 matrices. American Mathematical Monthly, 97(1):57–60, January 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Essen:1990:MSL

[27] Arno van den Essen. Magic squares and linear algebra. American Mathematical Monthly, 97(1):60–62, January 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Saari:1990:VNB

[28] Donald G. Saari. A visit to the Newtonian N-body problem via elementary complex variables. American Mathematical Monthly, 97(2):105– 119, February 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anglin:1990:SPP

[29] W. S. Anglin. The square pyramid puzzle. American Mathematical Monthly, 97(2):120–124, February 1990. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wagon:1990:ECE

[30] Stan Wagon. Editor’s corner: the Euclidean algorithm strikes again. American Mathematical Monthly, 97(2):125– 129, February 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Girstmair:1990:TNB

[31] Kurt Girstmair. A theorem on the numerators of the Bernoulli numbers. American Mathematical Monthly, 97 (2):136–138, February 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Swan:1990:APQ

[32] Richard G. Swan. Another proof of the quadratic reciprocity theorem. American Mathematical Monthly, 97(2):138– 139, February 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bender:1990:AP

[33] Edward A. Bender, Fred Kochman, and Douglas B. West. Adding up to powers. American Mathematical Monthly, 97(2):139–143, February 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zagier:1990:OSP

[34] D. Zagier. A one-sentence proof that every prime p ≡ 1 (mod 4) is a sum of two squares. American Mathematical Monthly, 97(2):144, February 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Friedberg:1990:EPA

[35] Stephen H. Friedberg. Extending the principal axis theorem to fields other than R. American Mathematical Monthly, 97(2):147–149, February 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beukers:1990:APL

[36] F. Beukers, J.-P. Bézivin, and P. Robba. An alternative proof of the Lindemann-Weierstrass theorem. American Mathematical Monthly, 97 (3):193–197, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bollinger:1990:LTS

[37] Richard C. Bollinger and Charles L. Burchard. Lucas’s theorem and some related results for extended Pascal triangles. American Mathematical Monthly, 97(3):198–204, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Poffald:1990:RTF

[38] Esteban I. Poffald. The remainder in Taylor’s formula. American Mathematical Monthly, 97(3):205–213, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Robinson:1990:CMP

[39] Raphael M. Robinson. Comment to: “A motivated proof of the Rogers– Ramanujan identities” [Amer. Math. Monthly 96 (1989), no. 5, 401–409, MR 90e:11147] by G. E. Andrews and R. J. Baxter. With a reply by Andrews. American Mathematical Monthly, 97 (3):214–215, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [13].

Thomassen:1990:LBJ

[40] Carsten Thomassen. A link between the Jordan curve theorem and the Kuratowski planarity criterion. American Mathematical Monthly, 97(3):216– 218, March 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ewell:1990:NSR

[41] John A. Ewell. A new series representation for ζ(3). American Mathematical Monthly, 97(3):219–220, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Minda:1990:DPD

[42] David Minda. The Dirichlet problem for a disk. American Mathematical Monthly, 97(3):220–223, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Richter:1990:DCR

[43] R. Bruce Richter and William P. Wardlaw. Diagonalization over commutative rings. American Mathematical Monthly, 97(3):223–227, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Whittaker:1990:RT

[44] Joe Whittaker. Random triangles. American Mathematical Monthly, 97 (3):228–230, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kopel:1990:NED

[45] Daniel Kopel and Michael Schramm. A new extension of the derivative. American Mathematical Monthly, 97 (3):230–233, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Folland:1990:RET

[46] G. B. Folland. Remainder estimates in Taylor’s theorem. American Mathematical Monthly, 97(3):233–235, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1990:SPI

[47] Donald E. Knuth. Solution to problem 6575, an identity involving sums and products. American Mathematical Monthly, 97(3):256, March 1990. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beck:1990:GSE

[48] Anatole Beck. Greed is (sometimes) not enough. American Mathematical Monthly, 97(4):289–294, April 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jankovic:1990:NTO

[49] Dragan Janković and T. R. Hamlett. New topologies from old via ideals. American Mathematical Monthly, 97 (4):295–310, April 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Calderon:1990:LE

[50] Calixto P. Calderón. Letter to the editor. American Mathematical Monthly, 97(4):311–312, April 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Garrison:1990:PLN

[51] Betty Garrison. Polynomials with large numbers of prime values. American Mathematical Monthly, 97(4):316–317, April 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cannonito:1990:IGE

[52] F. B. Cannonito. On inequivalent group extensions. American Mathematical Monthly, 97(4):317–319, April 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Vince:1990:RIP

[53] A. Vince. A rearrangement inequality and the permutahedron. American Mathematical Monthly, 97(4):319–323, April 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morton:1990:MPD

[54] Patrick Morton. Musings on the prime divisors of arithmetic sequences. American Mathematical Monthly, 97 (4):323–328, April 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Senechal:1990:FFG

[55] Marjorie Senechal. Finding the finite groups of symmetries of the sphere. American Mathematical Monthly, 97 (4):329–335, April 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Burgess:1990:CFC

[56] C. E. Burgess. Continuous functions and connected graphs. American Mathematical Monthly, 97(4):337–339, April 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leslie:1990:MNS

[57] Robert A. Leslie. The l_{2} matrix norm and the spectral radius. American Mathematical Monthly, 97(4):339–341, April 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lax:1990:ECS

[58] Peter D. Lax. The ergodic character of sequences of pedal triangles. American Mathematical Monthly, 97 (5):377–381, May 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Groemer:1990:SPG

[59] H. Groemer. Stability properties of geometric inequalities. American Mathematical Monthly, 97(5):382–394, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Richman:1990:SED

[60] Fred Richman. Separable extensions and diagonalizability. American Mathematical Monthly, 97(5):395–398, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hirschhorn:1990:BPR

[61] M. D. Hirschhorn. A birthday present for Ramanujan. American Mathematical Monthly, 97(5):398–400, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kupitz:1990:CFC

[62] Yaakov S. Kupitz and Micha A. Perles. A condition for flatness of curves in R^{n}. American Mathematical Monthly, 97(5):401–405, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Groen:1990:CVN

[63] Pieter P. N. de Groen. A counterexample on vector norms and the subordinate matrix norms. American Mathematical Monthly, 97(5):406–407, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Petkovsek:1990:AND

[64] Marko Petkovšek. Ambiguous numbers are dense. American Mathematical Monthly, 97(5):408–411, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boyd:1990:ADE

[65] David W. Boyd. Addendum: “The Diophantine equation x^{2} + y^{m} = z^{2n} ” [Amer. Math. Monthly 95 (1988), no. 6, 544–547, MR 89f:11046]. American Mathematical Monthly, 97(5):411–412, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [7].

Beebee:1990:EEI

[66] John Beebee. Errata: “Examples of infinite, incongruent exact covers” [Amer. Math. Monthly 95 (1988), no. 2, 121–123, MR 89g:11013]. American Mathematical Monthly, 97(5):412, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [5].

Benjamin:1990:EBM

[67] Arthur T. Benjamin. Errata: “The bisection method: which root?” [Amer. Math. Monthly 94 (1987), no. 9, 861– 863, MR 89c:26016]. American Mathematical Monthly, 97(5):412, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [4].

Kleiner:1990:ECP

[68] Hans-J. Kleiner. Errata: “Coloring the Platonic solids” [Amer. Math. Monthly 93 (1986), no. 4, 288–292, MR 87e:05072] by D. H. Lehmer. American Mathematical Monthly, 97(5):412, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [3].

Magill:1990:EUT

[69] K. D. Magill, Jr. Errata: “Universal topological spaces” [Amer. Math. Monthly 95 (1988), no. 10, 942–946, MR 89m:54018]. American Mathematical Monthly, 97(5):412, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [9].

Wong:1990:EMV

[70] Jingcheng Wong. Errata: “On the mean value theorem for integrals” [Amer. Math. Monthly 89 (1982), no. 5, 300–301, MR 83g:26002] by B. Jacobson. American Mathematical Monthly, 97(5):412, May 1990. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [2].

Hoechsmann:1990:SVS

[71] K. Hoechsmann. Singular values and the spectral theorem. American Mathematical Monthly, 97(5):413–414, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bartoszynski:1990:RSC

[72] Robert Bartoszyński and Wai Chan. A remark on the shortest confidence interval of a normal mean. American Mathematical Monthly, 97(5):415–417, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dubeau:1990:CBS

[73] François Dubeau. Cauchy-Bunyakowski-Schwarz inequality revisited. American Mathematical Monthly, 97(5):419–421, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Embry-Wardrop:1990:OMM

[74] Mary Embry-Wardrop. An old maxmin problem revisited. American Mathematical Monthly, 97(5):421–423, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Aczel:1990:FEL

[75] J. Aczél. Functional equations and L’Hôpital’s rule in an exact Poisson derivation. American Mathematical Monthly, 97(5):423–426, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Meyer:1990:GHO

[76] Kenneth R. Meyer. The geometry of harmonic oscillators. American Mathematical Monthly, 97(6):457– 465, June/July 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eisele:1990:GCD

[77] P. Eisele and K. P. Hadeler. Game of cards, dynamical systems, and a characterization of the floor and ceiling functions. American Mathematical Monthly, 97(6):466–477, June/July 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

London:1990:DTT

[78] R. R. London and H. P. Rogosinski. Decomposition theory in the teaching of elementary linear algebra. American Mathematical Monthly, 97(6):478– 485, June/July 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Granville:1990:RBC

[79] Andrew Granville and Yiliang Zhu. Representing binomial coefficients as sums of squares. American Mathematical Monthly, 97(6):486–493, June/July 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Robertson:1990:GRO

[80] Tim Robertson and Carolyn Pillers. A geometric regularity from orderrestricted statistical inference. American Mathematical Monthly, 97(6):496– 498, June/July 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ross:1990:SCI

[81] Bertram Ross and Baldev K. Sachdeva. The solution of certain integral equations by means of operators of arbitrary order. American Mathematical Monthly, 97(6):498–503, June/July 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mercier:1990:RB

[82] Armel Mercier. Relations between ω(n) and Ω(n). American Mathematical Monthly, 97(6):503–505, June/July 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mathews:1990:GTC

[83] Jerold Mathews. Gear trains and continued fractions. American Mathematical Monthly, 97(6):505–510, June/July 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Halmos:1990:PMS

[84] Paul R. Halmos. Has progress in mathematics slowed down. American Mathematical Monthly, 97(7):561–588, August/September 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ferguson:1990:TTT

[85] Helaman Ferguson. Two theorems, two sculptures, two posters. American Mathematical Monthly, 97(7):589–610, August/September 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1990:SAP

[86] Donald E. Knuth. Solution to advanced problem 6581, the asymptotic expansion of the middle binomial coefficient. American Mathematical Monthly, 97(7):629–630, August/September 1990. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [8].

Berger:1990:C

[87] Marcel Berger. Convexity. American Mathematical Monthly, 97(8):650– 678, October 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chern:1990:WG

[88] Shiing Shen Chern. What is geometry. American Mathematical Monthly, 97(8):679–686, October 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Graham:1990:WTC

[89] Ron Graham and Frances Yao. A whirlwind tour of computational geometry. American Mathematical Monthly, 97(8):687–701, October 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hoffman:1990:MSB

[90] David Hoffman and William H. Meeks, III. Minimal surfaces based on the catenoid. American Mathematical Monthly, 97(8):702–730, October 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Osserman:1990:CE

[91] Robert Osserman. Curvature in the eighties. American Mathematical Monthly, 97(8):731–756, October 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Thurston:1990:CTG

[92] William P. Thurston. Conway’s tiling groups. American Mathematical Monthly, 97(8):757–773, October 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Feder:1990:T

[93] Tomás Feder.  Toetjes.  Ameri-

can Mathematical Monthly, 97(9):785– 794, November 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cohen:1990:SIT

[94] Regina B. Cohen. Set isometries and their extensions in the max metric. American Mathematical Monthly, 97 (9):795–808, November 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ran:1990:SIS

[95] A. C. M. Ran and L. Rodman. On stability of invariant subspaces of matrices. American Mathematical Monthly, 97(9):809–823, November 1990. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Marsaglia:1990:NDS

[96] George Marsaglia and John C. W. Marsaglia. A new derivation of Stirling’s approximation to n!. American Mathematical Monthly, 97(9):826– 829, November 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jamieson:1990:FNA

[97] M. J. Jamieson. Fibonacci numbers and Aitken sequences revisited. American Mathematical Monthly, 97(9):829– 831, November 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bloom:1990:CTB

[98] Stephen L. Bloom, Zoltán Ésik, and Ernest G. Manes. A Cayley theorem for Boolean algebras. American Mathematical Monthly, 97(9):831– 833, November 1990. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Filaseta:1990:RTP

[99] Michael Filaseta. Rouché’s theorem for polynomials. American Mathematical Monthly, 97(9):834–835, November 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grunbaum:1990:SMP

[100] Branko Grünbaum and Jan Mycielski. Some models of plane geometries. American Mathematical Monthly, 97 (9):839–846, November 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Justicz:1990:RI

[101] Joyce Justicz, Edward R. Scheinerman, and Peter M. Winkler. Random intervals. American Mathematical Monthly, 97(10):881–889, December 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Schwarz:1990:DSM

[102] Gideon E. Schwarz. The dark side of the Moebius strip. American Mathematical Monthly, 97(10):890–897, December 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ungar:1990:MPP

[103] Peter Ungar. Mixing property of the pedal mapping. American Mathematical Monthly, 97(10):898–900, December 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mirollo:1990:IRF

[104] R. E. Mirollo and S. H. Strogatz. Integral representation of a finite spike. American Mathematical Monthly, 97 (10):901–903, December 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Desbrow:1990:I

[105] D. Desbrow. On the irrationality of π^{2}. American Mathematical Monthly, 97 (10):903–906, December 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1990:EPEb

[106] Donald E. Knuth and Boris Pittel. Elementary problem E3411, two sums over compositions. American Mathematical Monthly, 97(10):916–917, December 1990. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [250].

Knuth:1991:APG

[107] Donald E. Knuth. Advanced problem 6649, a generalized gamma function with independent branches. American Mathematical Monthly, 98(??):168, 1991. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [369, 368].

Knuth:1991:EPEb

[108] Donald E. Knuth and John McCarthy. Elementary problem E3429, small pills. American Mathematical Monthly, 98 (??):264, 1991. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1991:EPEc

[109] Donald E. Knuth. Elementary problem E3463, points in a circle. American Mathematical Monthly, 98(??):852, 1991. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mallows:1991:CCS

[110] Colin L. Mallows. Conway’s challenge sequence. American Mathematical Monthly, 98(1):5–20, January 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Darmon:1991:CFP

[111] Henri Darmon and John McKay. A continued fraction and permutations with fixed points. American Mathematical Monthly, 98(1):25–27, January 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Neves:1991:SFP

[112] Vítor Neves. Smoothness from finite points. American Mathematical Monthly, 98(1):27–31, January 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Toussaint:1991:AP

[113] Godfried Toussaint. Anthropomorphic polygons. American Mathematical Monthly, 98(1):31–35, January 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Day:1991:IAA

[114] W. A. Day. Inequalities for areas associated with conics. American Mathematical Monthly, 98(1):36–39, January 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kuhn:1991:DC

[115] Stephen Kuhn. The derivative à la carathéodory. American Mathematical Monthly, 98(1):40–44, January 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gilbert:1991:PDM

[116] George T. Gilbert. Positive definite matrices and Sylvester’s criterion. American Mathematical Monthly, 98 (1):44–46, January 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Foster:1991:GPP

[117] J. Foster and F. B. Richards. The Gibbs phenomenon for piecewise-linear approximation. American Mathematical Monthly, 98(1):47–49, January 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Flajolet:1991:EPE

[118] Philippe Flajolet and Donald E. Knuth. Elementary problem E3415, a hypergenerating function. American Mathematical Monthly, 98(1):54, January 1991. ISSN 0002-9890 (print), 1930-0972 (electronic). Solutions in [289].

DeTemple:1991:CCL

[119] Duane W. DeTemple. Carlyle circles and the Lemoine simplicity of polygon constructions. American Mathematical Monthly, 98(2):97–108, February 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bear:1991:PMH

[120] H. S. Bear. Part metric and hyperbolic metric. American Mathematical Monthly, 98(2):109–123, February 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lance:1991:APM

[121] Timothy Lance and Edward Thomas. Arcs with positive measure and a space-filling curve. American Mathematical Monthly, 98(2):124–127, February 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rao:1991:UCC

[122] K. P. S. Bhaskara Rao and A. Ramachandra Rao. Unions and common complements of subspaces. American Mathematical Monthly, 98(2):127– 131, February 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Liebeck:1991:GAR

[123] Hans Liebeck and Anthony Osborne. The generation of all rational orthogonal matrices. American Mathematical Monthly, 98(2):131–133, February 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Richert:1991:HCF

[124] Norman Richert. Hypocycloids, continued fractions, and distribution modulo one. American Mathematical Monthly, 98(2):133–139, February 1991. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Butler:1991:PR

[125] James P. Butler. The perimeter of a rose. American Mathematical Monthly, 98(2):139–143, February 1991. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Meijer:1991:TAS

[126] A. R. Meijer. A tree among the stepping stones. American Mathematical Monthly, 98(2):144–146, February 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Liu:1991:ATP

[127] Wen Liu. An analytic technique to prove Borel’s strong law of large numbers. American Mathematical Monthly, 98(2):146–148, February 1991. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rosenholtz:1991:TMV

[128] Ira Rosenholtz. A topological mean value theorem for the plane. American Mathematical Monthly, 98(2):149– 154, February 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Carlip:1991:DSR

[129] Walter Carlip. Disks and shells revisited. American Mathematical Monthly, 98(2):154–156, February 1991. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hartig:1991:LRI

[130] Donald Hartig. L’Hôpital’s rule via integration. American Mathematical Monthly, 98(2):156–157, February 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gilbert:1991:HTF

[131] E. N. Gilbert. How things float. American Mathematical Monthly, 98 (3):201–216, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Devaney:1991:EEO

[132] Robert L. Devaney and Marilyn B. Durkin. The exploding exponential and other chaotic bursts in complex dynamics. American Mathematical Monthly, 98(3):217–233, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Deshpande:1991:ACC

[133] M. V. Deshpande and Sudhir R. Ghorpade. Algebraic characterization of compact abelian groups. American Mathematical Monthly, 98(3):235–237, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Myerson:1991:FCF

[134] G. Myerson. First-class functions. American Mathematical Monthly, 98 (3):237–240, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Barry:1991:SPF

[135] P. D. Barry and D. J. Hurley. On series of partial fractions. American Mathematical Monthly, 98(3):240–242, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Loeb:1991:NDP

[136] Peter A. Loeb. A note on Dixon’s proof of Cauchy’s integral theorem. American Mathematical Monthly, 98 (3):242–244, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goffinet:1991:NSC

[137] Daniel Goffinet. Number systems with a complex base: a fractal tool for teaching topology. American Mathematical Monthly, 98(3):249–255, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chalice:1991:CCF

[138] Donald R. Chalice. A characterization of the Cantor function. American Mathematical Monthly, 98(3):255– 258, March 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bragg:1991:PDR

[139] L. R. Bragg. Parametric differentiation revisited. American Mathematical Monthly, 98(3):259–262, March 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lam:1991:SFP

[140] C. W. H. Lam. The search for a finite projective plane of order 10. American Mathematical Monthly, 98 (4):305–318, April 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Metropolis:1991:SCF

[141] N. Metropolis and Gian-Carlo Rota. Symmetry classes: functions of three variables. American Mathematical Monthly, 98(4):328–332, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beigel:1991:INT

[142] Richard Beigel. Irrationality without number theory. American Mathematical Monthly, 98(4):332–335, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hall:1991:AEA

[143] J. I. Hall. Another elementary approach to the Jordan form. American Mathematical Monthly, 98(4):336–340, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lindstrom:1991:BCI

[144] Bernt Lindström and Hans-Olov Zetterström. Borromean circles are impossible. American Mathematical Monthly, 98(4):340–341, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Miamee:1991:I

[145] A. G. Miamee. The inclusion L^{p}(μ) ⊆ L^{q}(ν). American Mathematical Monthly, 98(4):342–345, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rudolph:1991:SOT

[146] Lee Rudolph. The structure of orthogonal transformations. American Mathematical Monthly, 98(4):349–352, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lewin:1991:SPZ

[147] Jonathan Lewin. A simple proof of Zorn’s lemma. American Mathematical Monthly, 98(4):353–354, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grant:1991:CLT

[148] Richard Grant. The converse of Liouville’s theorem. American Mathematical Monthly, 98(4):354, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cuoco:1991:VAI

[149] Albert A. Cuoco. Visualizing the padic integers. American Mathematical Monthly, 98(4):355–364, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wilcox:1991:CPW

[150] Calvin H. Wilcox. The Cauchy problem for the wave equation with distribution data: an elementary approach. American Mathematical Monthly, 98 (5):401–410, May 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Katsuura:1991:CND

[151] Hidefumi Katsuura. Continuous nowhere-differentiable functions — an application of contraction mappings. American Mathematical Monthly, 98 (5):411–416, May 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bienstock:1991:NFS

[152] Daniel Bienstock, Fan Chung, Michael L. Fredman, Alejandro A. Schäffer, Peter W. Shor, and Subhash Suri. A note on finding a strict saddlepoint. American Mathematical Monthly, 98 (5):418–419, May 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Osserman:1991:CC

[153] Robert Osserman. Circumscribed circles. American Mathematical Monthly, 98(5):419–422, May 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Deeba:1991:SSB

[154] Elias Y. Deeba and Dennis M. Rodriguez. Stirling’s series and Bernoulli numbers. American Mathematical Monthly, 98(5):423–426, May 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lindsey:1991:SPW

[155] J. H. Lindsey, II. A simple proof of the Weierstrass approximation theorem. American Mathematical Monthly, 98(5):429–430, May 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Henrich:1991:MSL

[156] Christopher J. Henrich. Magic squares and linear algebra. American Mathematical Monthly, 98(6):481–488, June/ July 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Whittaker:1991:ADS

[157] James V. Whittaker. An analytical description of some simple cases of chaotic behaviour. American Mathematical Monthly, 98(6):489–504, June/ July 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fassler:1991:MPN

[158] Albert Fässler. Multiple Pythagorean number triples. American Mathematical Monthly, 98(6):505–517, June/July 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rosay:1991:VEP

[159] Jean-Pierre Rosay. A very elementary proof of the Malgrange-Ehrenpreis theorem. American Mathematical Monthly, 98(6):518–523, June/ July 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shen:1991:Z

[160] Li-Chien Shen. On the zeros of tan^{(k)} z. American Mathematical Monthly, 98 (6):524–527, June/July 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kelly:1991:TES

[161] John B. Kelly. Two equal sums of three squares with equal products. American Mathematical Monthly, 98(6):527– 529, June/July 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gessel:1991:CP

[162] Ira M. Gessel. A coloring problem. American Mathematical Monthly, 98 (6):530–533, June/July 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kwong:1991:NIP

[163] Man Kam Kwong and A. Zettl. Norm inequalities for the powers of a matrix. American Mathematical Monthly, 98 (6):533–538, June/July 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mazur:1991:NTG

[164] B. Mazur. Number theory as gadfly. American Mathematical Monthly, 98 (7):593–610, August/September 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Riddle:1991:OBN

[165] Lawrence H. Riddle. An occurrence of the ballot numbers in operator theory. American Mathematical Monthly, 98 (7):613–617, August/September 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Penrice:1991:DPC

[166] Stephen G. Penrice. Derangements, permanents, and Christmas presents. American Mathematical Monthly, 98 (7):617–620, August/September 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Galvin:1991:DFT

[167] Fred Galvin and S. D. Shore. Distance functions and topologies. American Mathematical Monthly, 98(7):620– 623, August/September 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jankovic:1991:CBS

[168] Dragan Janković, Maximilian Ganster, and Ivan Reilly. A characterization of Baire spaces. American Mathematical Monthly, 98(7):624–625, August/September 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Siu:1991:WLS

[169] Man Keung Siu. Which Latin squares are Cayley tables. American Mathematical Monthly, 98(7):625–627, August/September 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Weiss:1991:EID

[170] Marysia T. Weiss. An early introduction to dynamics. American Mathematical Monthly, 98(7):635–641, August/September 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gluchoff:1991:SIC

[171] Alan Gluchoff. A simple interpretation of the complex contour integral. American Mathematical Monthly, 98(7):641– 644, August/September 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jones:1991:PRU

[172] J. P. Jones and Y. V. Matijasevič. Proof of recursive unsolvability of Hilbert’s tenth problem. American Mathematical Monthly, 98(8):689– 709, October 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goldstein:1991:LAQ

[173] Jerome A. Goldstein and Mel Levy. Linear algebra and quantum chemistry. American Mathematical Monthly, 98 (8):710–718, October 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dekking:1991:BPG

[174] F. M. Dekking. Branching processes that grow faster than binary splitting. American Mathematical Monthly, 98 (8):728–731, October 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leslie:1991:HRD

[175] Robert A. Leslie. How not to repeatedly differentiate a reciprocal. American Mathematical Monthly, 98(8):732– 735, October 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:1991:NR

[176] Jonathan M. Borwein and G. de Barra. Nested radicals. American Mathematical Monthly, 98(8):735–739, October 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berger:1991:IFC

[177] M. A. Berger, A. Felzenbaum, A. S. Fraenkel, and R. Holzman. On infinite and finite covering systems. American Mathematical Monthly, 98(8):739– 742, October 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Takacs:1991:MCT

[178] Lajos Takács. A moment convergence theorem. American Mathematical Monthly, 98(8):742–746, October 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ross:1991:WMU

[179] Clay C. Ross, Jr. Why the method of undetermined coefficients works.

American Mathematical Monthly, 98 (8):747–749, October 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boor:1991:BST

[180] Carl de Boor and Klaus Höllig. Boxspline tilings. American Mathematical Monthly, 98(9):793–802, November 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Alpern:1991:HGT

[181] Steve Alpern and Anatole Beck. Hex games and twist maps on the annulus. American Mathematical Monthly, 98 (9):803–811, November 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Flath:1991:HPI

[182] Dan Flath and Stan Wagon. How to pick out the integers in the rationals: an application of number theory to logic. American Mathematical Monthly, 98(9):812–823, November 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Garabedian:1991:SPS

[183] P. R. Garabedian. A simple proof of a simple version of the Riemann mapping theorem by simple functional analysis. American Mathematical Monthly, 98 (9):824–826, November 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Villani:1991:E

[184] Alfonso Villani. On the equality L^{1}(μ) = L_{loc}^{1}(μ). American Mathematical Monthly, 98(9):826–828, November 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Caen:1991:RTM

[185] D. de Caen. The ranks of tournament matrices. American Mathematical Monthly, 98(9):829–831, November 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Smith:1991:CPF

[186] Arthur Smith. Convergence-preserving functions: an alternative discussion. American Mathematical Monthly, 98 (9):831–833, November 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Duffin:1991:LBF

[187] R. J. Duffin. Lines of best fit by graphics and the Wald line. American Mathematical Monthly, 98(9):835– 839, November 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Furi:1991:MVT

[188] M. Furi and M. Martelli. On the mean value theorem, inequality, and inclusion. American Mathematical Monthly, 98(9):840–846, November 1991. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Groetsch:1991:DAS

[189] C. W. Groetsch. Differentiation of approximately specified functions. American Mathematical Monthly, 98(9):847– 850, November 1991. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Binding:1991:MCP

[190] Paul Binding. More on cross products. American Mathematical Monthly, 98 (9):850–851, November 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stolarsky:1991:WNC

[191] Kenneth B. Stolarsky. From Wythoff’s Nim to Chebyshev’s inequality. American Mathematical Monthly, 98(10): 889–900, December 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kong:1991:TAD

[192] T. Yung Kong, Ralph Kopperman, and Paul R. Meyer. A topological approach to digital topology. American Mathematical Monthly, 98(10): 901–917, December 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hochwald:1991:LAA

[193] Scott H. Hochwald. Linear algebra by analogy. American Mathematical Monthly, 98(10):918–926, December 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bogart:1991:OPB

[194] Kenneth P. Bogart. An obvious proof of Burnside’s lemma. American Mathematical Monthly, 98(10): 927–928, December 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kral:1991:PIR

[195] J. Král and W. F. Pfeffer. Poisson integrals of Riemann integrable functions.

American Mathematical Monthly, 98 (10):929–931, December 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chew:1991:SRW

[196] James Chew and Jing Cheng Tong. Some remarks on weak continuity. American Mathematical Monthly, 98 (10):931–934, December 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dokovic:1991:PTA

[197] Dragomir Ž. – Doković. On the product of two alternating matrices. American Mathematical Monthly, 98(10): 935–936, December 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knight:1991:OZP

[198] M. J. Knight. An “oceans of zeros” proof that a certain non-liouville number is transcendental. American Mathematical Monthly, 98(10): 947–949, December 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1993:SEPc

[199] Donald E. Knuth. Solution to elementary problem E3463, points in a circle. American Mathematical Monthly, 100(??):693–694, 1993. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1992:SEPb

[200] Donald E. Knuth and John McCarthy. Solution to elementary problem E3429, small pills. American Mathematical Monthly, 99(??):684, 1992. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lynch:1992:CND

[201] Mark Lynch. A continuous, nowhere differentiable function. American Mathematical Monthly, 99(1):8–9, January 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gethner:1992:TRP

[202] Robert M. Gethner. Two relatives of Picard’s theorem on entire functions. American Mathematical Monthly, 99 (1):13–19, January 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Friedman:1992:RSE

[203] Nathaniel A. Friedman. Replication and stacking in ergodic theory. American Mathematical Monthly, 99(1):31– 41, January 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Segercrantz:1992:ICH

[204] J. Segercrantz. Improving the Cayley– Hamilton equation for low-rank transformations. American Mathematical Monthly, 99(1):42–44, January 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Colwell:1992:BFK

[205] Peter Colwell. Bessel functions and Kepler’s equation. American Mathematical Monthly, 99(1):45–48, January 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Libera:1992:LIC

[206] Richard J. Libera and Eligiusz J. Złotkiewicz. Löwner’s inverse coefficients theorem for starlike functions. American Mathematical Monthly, 99 (1):49–50, January 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Axler:1992:BT

[207] Sheldon Axler, Paul Bourdon, and Wade Ramey. Bôcher’s theorem. American Mathematical Monthly, 99 (1):51–55, January 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mera:1992:DIP

[208] Ruben Mera. On the determination of the intermediate point in Taylor’s theorem. American Mathematical Monthly, 99(1):56–58, January 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hoffman:1992:ADS

[209] Kenneth M. Hoffman and James R. C. Leitzel. Award for distinguished service to Dr. Lynn Arthur Steen. American Mathematical Monthly, 99(2):99– 100, February 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Richert:1992:SSF

[210] Norman Richert. Strang’s strange figures. American Mathematical Monthly, 99(2):101–107, February 1992. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taylor:1992:ZGZ

[211] Jean E. Taylor. Zonohedra and generalized zonohedra. American Mathematical Monthly, 99(2):108–111, February 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Velling:1992:URE

[212] John A. Velling. The uniformization of rectangles, an exercise in Schwarz’s lemma. American Mathematical Monthly, 99(2):112–115, February 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Thomassen:1992:JST

[213] Carsten Thomassen. The Jordan-Schönflies theorem and the classification of surfaces. American Mathematical Monthly, 99(2):116–130, February 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rosenholtz:1992:PPK

[214] Ira Rosenholtz. A pigeonhole proof of Kaplansky’s theorem. American Mathematical Monthly, 99(2):132– 133, February 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bruckner:1992:SAP

[215] A. M. Bruckner, J. Mařík, and C. E. Weil. Some aspects of products of derivatives. American Mathematical Monthly, 99(2):134–145, February 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brink:1992:BCG

[216] Chris Brink and Jan Pretorius. Boolean circulants, groups, and relation algebras. American Mathematical Monthly, 99(2):146–152, February 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Servatius:1992:CSD

[217] Brigitte Servatius and Peter R. Christopher. Construction of selfdual graphs. American Mathematical Monthly, 99(2):153–158, February 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gora:1992:FBV

[218] Paweł Góra and Abraham Boyarsky. On functions of bounded variation in higher dimensions. American Mathematical Monthly, 99(2):159–160, February 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Corless:1992:CFC

[219] R. M. Corless. Continued fractions and chaos. American Mathematical Monthly, 99(3):203–215, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beardon:1992:SSP

[220] A. F. Beardon and T. K. Carne. A strengthening of the Schwarz-Pick inequality. American Mathematical Monthly, 99(3):216–217, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dydak:1992:MTC

[221] Jerzy Dydak and Nathan Feldman. Major theorems on compactness: a unified exposition. American Mathematical Monthly, 99(3):220–227, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brualdi:1992:BEP

[222] R. A. Brualdi and J. Csima. Butterfly embedding proof of a theorem of König. American Mathematical Monthly, 99 (3):228–230, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McIntosh:1992:GCP

[223] Richard J. McIntosh. A generalization of a congruential property of Lucas. American Mathematical Monthly, 99(3):231–238, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Huff:1992:MOS

[224] Barthel W. Huff. Mixtures and order statistics. American Mathematical Monthly, 99(3):239–242, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beeson:1992:TVL

[225] Michael J. Beeson. Triangles with vertices on lattice points. American Mathematical Monthly, 99(3):243– 252, March 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stromberg:1992:UNS

[226] Karl R. Stromberg. Universally nonmeasurable subgroups of R. American Mathematical Monthly, 99(3):253– 255, March 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Egecioglu:1992:CGP

[227] Ömer Eğecioğlu. A combinatorial generalization of a Putnam problem.

American Mathematical Monthly, 99 (3):256–258, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kurtz:1992:SCA

[228] David C. Kurtz. A sufficient condition for all the roots of a polynomial to be real. American Mathematical Monthly, 99(3):259–263, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shanks:1992:IAP

[229] Daniel Shanks. Improving an approximation for pi. American Mathematical Monthly, 99(3):263, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Halmos:1992:LIL

[230] Paul R. Halmos. Large intersections of large sets. American Mathematical Monthly, 99(4):307–312, April 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Granville:1992:ZBB

[231] Andrew Granville. Zaphod Beeblebrox’s brain and the fifty-ninth row of Pascal’s triangle. American Mathematical Monthly, 99(4):318–331, April 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See correction [740].

Banks:1992:DDC

[232] J. Banks, J. Brooks, G. Cairns, G. Davis, and P. Stacey. On Devaney’s definition of chaos. American Mathematical Monthly, 99(4):332–334, April 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

King:1992:DSS

[233] Jonathan L. King. Dilemma of the sleeping stockbroker. American Mathematical Monthly, 99(4):335–338, April 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wetzel:1992:CNT

[234] John E. Wetzel. Converses of Napoleon’s theorem. American Mathematical Monthly, 99(4):339–351, April 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Isaacs:1992:TFS

[235] I. M. Isaacs and G. R. Robinson. On a theorem of Frobenius: solutions of x^{n} = 1 in finite groups. American Mathematical Monthly, 99 (4):352–354, April 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

VandenEynden:1992:PSC

[236] Charles Vanden Eynden. On a problem of Stein concerning infinite covers. American Mathematical Monthly, 99(4):355–358, April 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cliborn:1992:MS

[237] James H. Cliborn and Blake Jordan. Minimal surfaces. American Mathematical Monthly, 99(4):376, April 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1992:TNN

[238] Donald E. Knuth. Two notes on notation. American Mathematical Monthly, 99(5):403–422, May 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also remarks in [441].

Spearman:1992:RPB

[239] Blair K. Spearman and Kenneth S. Williams. Representing primes by binary quadratic forms. American Mathematical Monthly, 99(5):423–426, May 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gonzalez-Velasco:1992:CMA

[240] Enrique A. González-Velasco. Connections in mathematical analysis: the case of Fourier series. American Mathematical Monthly, 99(5):427–441, May 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fulton:1992:T

[241] Chandler Fulton. Tessellations. American Mathematical Monthly, 99(5):442– 445, May 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leavitt:1992:RFG

[242] J. L. Leavitt, G. J. Sherman, and M. E. Walker. Rewriteability in finite groups. American Mathematical Monthly, 99 (5):446–452, May 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Temple:1992:NE

[243] Blake Temple and Craig A. Tracy. From Newton to Einstein. American Mathematical Monthly, 99(6):507– 521, June/July 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See correction [382].

Boshernitzan:1992:BRP

[244] Michael D. Boshernitzan. Billiards and rational periodic directions in polygons. American Mathematical Monthly, 99(6):522–529, June/July 1992. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wermuth:1992:SEP

[245] Edgar M. E. Wermuth. Some elementary properties of infinite products. American Mathematical Monthly, 99 (6):530–537, June/July 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hinz:1992:PTT

[246] Andreas M. Hinz. Pascal’s triangle and the Tower of Hanoi. American Mathematical Monthly, 99(6):538– 544, June/July 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jungnickel:1992:UCG

[247] Dieter Jungnickel. On the uniqueness of the cyclic group of order n. American Mathematical Monthly, 99(6):545– 547, June/July 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forman:1992:SMP

[248] Robin Forman. Sequences with many primes. American Mathematical Monthly, 99(6):548–557, June/July 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maesumi:1992:PME

[249] Mohsen Maesumi. Parabolic mirrors, elliptic and hyperbolic lenses. American Mathematical Monthly, 99(6):558– 560, June/July 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Strong:1992:SEP

[250] Richard Strong. Solution to elementary problem E3411, two sums over compositions. American Mathematical Monthly, 99(6):578–579, June/July 1992. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [106].

Dyson:1992:PDC

[251] Freeman J. Dyson and Harold Falk. Period of a discrete cat mapping. American Mathematical Monthly, 99(7):603– 614, August/September 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bressoud:1992:WDW

[252] David M. Bressoud. Why do we teach calculus. American Mathematical Monthly, 99(7):615–617, August/September 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:1992:SSH

[253] J. M. Borwein and P. B. Borwein. Strange series and high precision fraud. American Mathematical Monthly, 99 (7):622–640, August/September 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Roman:1992:LBF

[254] Steven Roman. The logarithmic binomial formula. American Mathematical Monthly, 99(7):641–648, August/September 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Braden:1992:CSI

[255] Bart Braden. Calculating sums of infinite series. American Mathematical Monthly, 99(7):649–655, August/ September 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Alvarez:1992:A

[256] Sergio A. Alvarez. L^{p} arithmetic. American Mathematical Monthly, 99 (7):656–662, August/September 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fowler:1992:DT

[257] David Fowler. Dedekind’s theorem: \sqrt{2} × \sqrt{3} = \sqrt{6}. American Mathematical Monthly, 99(8):725–733, October 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knill:1992:MBA

[258] Ronald J. Knill. A modified Babylonian algorithm. American Mathematical Monthly, 99(8):734–737, October 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Marchisotto:1992:LO

[259] E. A. Marchisotto. Lines without order. American Mathematical Monthly, 99(8):738–745, October 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Golomb:1992:I

[260] Solomon W. Golomb. An identity for \binom{2n}{n}. American Mathematical Monthly, 99(8):746–748, October 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mead:1992:NI

[261] D. G. Mead. Newton’s identities. American Mathematical Monthly, 99 (8):749–751, October 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ewell:1992:STN

[262] John A. Ewell. On sums of triangular numbers and sums of squares. American Mathematical Monthly, 99(8):752– 757, October 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Vianello:1992:SCS

[263] Marco Vianello and Renato Zanovello. On the superlinear convergence of the secant method. American Mathematical Monthly, 99(8):758–761, October 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Subramaniam:1992:HIR

[264] T. N. Subramaniam and Donald E. G. Malm. How to integrate rational functions. American Mathematical Monthly, 99(8):762–772, October 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morawetz:1992:G

[265] Cathleen S. Morawetz. Giants. American Mathematical Monthly, 99(9):819– 828, November 1992. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eggleton:1992:EQF

[266] R. B. Eggleton, C. B. Lacampagne, and J. L. Selfridge. Euclidean quadratic fields. American Mathematical Monthly, 99(9):829–837, November 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kim:1992:OMS

[267] K. H. Kim, F. W. Roush, and M. D. Intriligator. Overview of mathematical social sciences. American Mathematical Monthly, 99(9):838–844, November 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Schilling:1992:SP

[268] Mark F. Schilling. Sequential partitioning. American Mathematical Monthly, 99(9):846–855, November 1992. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wang:1992:GPR

[269] Jun Wang. Goldbach’s problem in the ring M_{n}(Z). American Mathematical Monthly, 99(9):856–857, November 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Evard:1992:CRT

[270] J.-Cl. Evard and F. Jafari. A complex Rolle’s theorem. American Mathematical Monthly, 99(9):858–861, November 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brakke:1992:OCP

[271] Kenneth A. Brakke. The opaque cube problem. American Mathematical Monthly, 99(9):866–871, November 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rotando:1992:KCS

[272] Louis M. Rotando and Edward O. Thorp. The Kelly criterion and the stock market. American Mathematical Monthly, 99(10):922–931, December 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chernoff:1992:SPT

[273] Paul R. Chernoff. A simple proof of Tychonoff’s theorem via nets. American Mathematical Monthly, 99(10): 932–934, December 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fisher:1992:OSG

[274] David C. Fisher and Jennifer Ryan. Optimal strategies for a generalized “scissors, paper, and stone” game. American Mathematical Monthly, 99 (10):935–942, December 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chapman:1992:SEN

[275] Scott T. Chapman. A simple example of nonunique factorization in integral domains. American Mathematical Monthly, 99(10):943–945, December 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beebee:1992:BNE

[276] John Beebee. Bernoulli numbers and exact covering systems. American Mathematical Monthly, 99(10): 946–948, December 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bezdek:1992:HCC

[277] Károly Bezdek. Hadwiger’s covering conjecture and its relatives. American Mathematical Monthly, 99(10): 954–956, December 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1993:PRB

[278] Donald E. Knuth. Problem 10280, a random binary operation. American Mathematical Monthly, 100(??):76, 1993. ISSN 0002-9890 (print), 1930-0972 (electronic). Solution to appear.

Knuth:1993:PDP

[279] Donald E. Knuth. Problem 10298, a divisibility property of Stirling numbers. American Mathematical Monthly, 100 (??):400, 1993. ISSN 0002-9890 (print), 1930-0972 (electronic). Solution to appear.

Price:1993:SFA

[280] G. Baley Price. The seventy-fifth anniversary celebration. American Mathematical Monthly, 100(1):4–15, January 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gillman:1993:AAI

[281] Leonard Gillman. An axiomatic approach to the integral. American Mathematical Monthly, 100(1):16–25, January 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Szalkai:1993:VC

[282] István Szalkai and Dan Velleman. Versatile coins. American Mathematical Monthly, 100(1):26–33, January 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Duffin:1993:SPC

[283] R. J. Duffin. On seeing progressions of constant cross ratio. American Mathematical Monthly, 100(1):38– 47, January 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:1993:YME

[284] Anonymous. 100 years of Monthly editors. American Mathematical Monthly, 100(1):48–49, January 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hobby:1993:QP

[285] David Hobby and D. M. Silberger. Quotients of primes. American Mathematical Monthly, 100(1):50–52, January 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Donovan:1993:PFN

[286] George C. Donovan, Arnold R. Miller, and Timothy J. Moreland. Pathological functions for Newton’s method. American Mathematical Monthly, 100 (1):53–58, January 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Avez:1993:SPT

[287] André Avez. A short proof of a theorem of Erdős and Mordell. American Mathematical Monthly, 100(1):60– 62, January 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Houston:1993:LAA

[288] Evan G. Houston. A linear algebra approach to cyclic extensions in Galois theory. American Mathematical Monthly, 100(1):64–66, January 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Alvis:1993:SEP

[289] Dean Alvis. Solution to elementary problem E3415, a hypergenerating function. American Mathematical Monthly, 100(1):84–85, January 1993. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [118].

Gnanadesikan:1993:YGG

[290] Ramanathan Gnanadesikan and Henry J. Landau. Yueh-Gin Gung and Dr. Charles Y. Hu award for distinguished service to Henry O. Pollak. American Mathematical Monthly, 100(2): 115–116, February 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maxwell:1993:JMC

[291] John D. Maxwell. John Marvin Colaw: and the American Mathematical Monthly. American Mathematical Monthly, 100(2):117–118, February 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Needham:1993:NTF

[292] Tristan Needham. Newton and the transmutation of force. American Mathematical Monthly, 100(2):119– 137, February 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Baxa:1993:NDR

[293] Christoph Baxa. A note on Diophantine representations. American Mathematical Monthly, 100(2): 138–143, February 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shore:1993:RSR

[294] Terence R. Shore and Douglas B. Tyler. Recurrence of simple random walk in the plane. American Mathematical Monthly, 100(2):144–149, February 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grunbaum:1993:PT

[295] Branko Grünbaum and G. C. Shephard. Pick’s theorem. American Mathematical Monthly, 100(2): 150–161, February 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Abel:1993:SLN

[296] Ulrich Abel and Hartmut Siebert. Sequences with large numbers of prime values. American Mathematical Monthly, 100(2):167–169, February 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Martini:1993:RSS

[297] Horst Martini. Regular simplices in spaces of constant curvature. American Mathematical Monthly, 100(2): 169–171, February 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Williams:1993:SEL

[298] John K. Williams. A simple example of little big set. American Mathematical Monthly, 100(2):172–174, February 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lorch:1993:S

[299] Edgar R. Lorch. Szeged in 1934. American Mathematical Monthly, 100 (3):219–230, March 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Burris:1993:THS

[300] Stanley Burris and Simon Lee. Tarski’s high school identities. American Mathematical Monthly, 100(3):231– 236, March 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

MacEachern:1993:ACO

[301] Steven N. MacEachern and L. Mark Berliner. Aperiodic chaotic orbits. American Mathematical Monthly, 100 (3):237–241, March 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gilbert:1993:BHN

[302] William J. Gilbert. Bricklaying and the Hermite normal form. American Mathematical Monthly, 100(3): 242–245, March 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Falkner:1993:CIP

[303] Neil Falkner. A characterization of inner product spaces. American Mathematical Monthly, 100(3):

246–249, March 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Strang:1993:PAA

[304] Gilbert Strang. Polar area is the average of strip areas. American Mathematical Monthly, 100(3):250–254, March 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Durfee:1993:CCP

[305] Alan Durfee, Nathan Kronenfeld, Heidi Munson, Jeff Roy, and Ina Westby. Counting critical points of real polynomials in two variables. American Mathematical Monthly, 100(3): 255–271, March 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Walker:1993:SZP

[306] Peter Walker. Separation of the zeros of polynomials. American Mathematical Monthly, 100(3):272–273, March 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Andrews:1993:SPJ

[307] George E. Andrews, Shalosh B. Ekhad, and Doron Zeilberger. A short proof of Jacobi’s formula for the number of representations of an integer as a sum of four squares. American Mathematical Monthly, 100(3):274–276, March 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Oleszkiewicz:1993:EPH

[308] Krzysztof Oleszkiewicz. An elementary proof of Hilbert’s inequality. American Mathematical Monthly, 100(3):

276–280, March 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Debieve:1993:NFT

[309] Camille Debiève. A note on Fubini’s theorem. American Mathematical Monthly, 100(3):281, March 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nijenhuis:1993:SEP

[310] Albert Nijenhuis. Solution to elementary problem E3267, a solitaire game. American Mathematical Monthly, 100 (3):292–294, March 1993. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [6].

Silverman:1993:TST

[311] Joseph H. Silverman. Taxicabs and sums of two cubes. American Mathematical Monthly, 100(4):331–340, April 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ayoub:1993:WNL

[312] Raymond Ayoub. What is a Napierian logarithm. American Mathematical Monthly, 100(4):351–364, April 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bruce:1993:RTP

[313] J. W. Bruce. A really trivial proof of the Lucas-Lehmer test. American Mathematical Monthly, 100(4): 370–371, April 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Call:1993:PM

[314] Gregory S. Call and Daniel J. Velleman. Pascal’s matrices. American Mathematical Monthly, 100(4): 372–376, April 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fournelle:1993:SCO

[315] Thomas A. Fournelle. Symmetries of the cube and outer automorphisms of S_{6}. American Mathematical Monthly, 100(4):377–380, April 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Murdoch:1993:IC

[316] Timothy A. Murdoch. Isogonal configurations. American Mathematical Monthly, 100(4):381–384, April 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gardner:1993:TAHa

[317] J. Helen Gardner and Robin J. Wilson. Thomas Archer Hirst—mathematician xtravagant. I. A Yorkshire surveyor. American Mathematical Monthly, 100 (5):435–441, May 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Reynolds:1993:HGH

[318] William F. Reynolds. Hyperbolic geometry on a hyperboloid. American Mathematical Monthly, 100(5): 442–455, May 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rubinstein:1993:SHP

[319] Michael Rubinstein. A simple heuristic proof of Hardy and Littlewood’s conjecture B. American Mathematical Monthly, 100(5):456–460, May 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Laquer:1993:PP

[320] H. Turner Laquer. The Pompeiu problem. American Mathematical Monthly, 100(5):461–467, May 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

DeTemple:1993:QCE

[321] Duane W. DeTemple. A quicker convergence to Euler’s constant. American Mathematical Monthly, 100(5): 468–470, May 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Watkins:1993:MP

[322] William Watkins and Joel Zeitlin. The minimal polynomial of cos(2π/n). American Mathematical Monthly, 100 (5):471–474, May 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Richmond:1993:EAZ

[323] B. Richmond and T. Richmond. The equal area zones property. American Mathematical Monthly, 100(5): 475–477, May 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Copper:1993:GTG

[324] Mark Copper. Graph theory and the game of sprouts. American Mathematical Monthly, 100(5):478–482, May 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bojanov:1993:EPR

[325] Borislav Bojanov. Elementary proof of the Remez inequality. American Mathematical Monthly, 100(5):483–485, May 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nanjundiah:1993:NIR

[326] T. S. Nanjundiah. A note on an identity of Ramanujan. American Mathematical Monthly, 100(5):485–487, May 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zeilberger:1993:ID

[327] Doron Zeilberger. On an identity of Daubechies. American Mathematical Monthly, 100(5):487, May 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McGeoch:1993:DC

[328] Catherine C. McGeoch. Data compression. American Mathematical Monthly, 100(5):493–497, May 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gardner:1993:TAHb

[329] J. Helen Gardner and Robin J. Wilson. Thomas Archer Hirst— mathematician xtravagant. II. Student days in Germany. American Mathematical Monthly, 100(6):531–538, June/ July 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Strichartz:1993:HMW

[330] Robert S. Strichartz. How to make wavelets. American Mathematical Monthly, 100(6):539–556, June/July 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Waterhouse:1993:MM

[331] William C. Waterhouse. A matrix maximum. American Mathematical Monthly, 100(6):557–562, June/July 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hastings:1993:CMP

[332] S. P. Hastings and J. B. McLeod. Chaotic motion of a pendulum with oscillatory forcing. American Mathematical Monthly, 100(6):563–572, June/ July 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wagstaff:1993:AC

[333] David A. Wagstaff, Theodore A. Norman, and Douglas M. Campbell. An application for the curiosity (log_{x} N)′. American Mathematical Monthly, 100 (6):573–574, June/July 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grosof:1993:VSA

[334] Miriam Schapiro Grosof and Geraldine Taiani. Vandermonde strikes again. American Mathematical Monthly, 100 (6):575–577, June/July 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Galvin:1993:ECG

[335] Fred Galvin. Embedding countable groups in 2-generator groups. American Mathematical Monthly, 100(6): 578–580, June/July 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gallian:1993:AFS

[336] Joseph A. Gallian and Michael Reid. Abelian forcing sets. American Mathematical Monthly, 100(6):580– 582, June/July 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gardner:1993:TAHc

[337] J. Helen Gardner and Robin J. Wilson. Thomas Archer Hirst—mathematician xtravagant. III. Göttingen and Berlin. American Mathematical Monthly, 100 (7):619–625, August/September 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beyer:1993:BTT

[338] W. A. Beyer and Blair Swartz. Bisectors of triangles and tetrahedra. American Mathematical Monthly, 100 (7):626–640, August/September 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mead:1993:MRT

[339] D. G. Mead and S. K. Stein. More on rectangles tiled by rectangles. American Mathematical Monthly, 100(7): 641–643, August/September 1993. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berndt:1993:RL

[340] Bruce C. Berndt and S. Bhargava. Ramanujan — for lowbrows. American Mathematical Monthly, 100(7): 644–656, August/September 1993. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Savio:1993:CPR

[341] D. Y. Savio and E. R. Suryanarayan. Chebychev polynomials and regular polygons. American Mathematical Monthly, 100(7):657–661, August/ September 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ding:1993:FPT

[342] Ren Ding, Krzysztof Kołodziejczyk, Grattan Murphy, and John Reay. A fast Pick-type approximation for areas of H-polygons. American Mathematical Monthly, 100(7):669–673, August/September 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gardner:1993:TAHd

[343] J. Helen Gardner and Robin J. Wilson. Thomas Archer Hirst—mathematician xtravagant. IV. Queenwood, France and Italy. American Mathematical Monthly, 100(8):723–731, October 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mornhinweg:1993:PAT

[344] David Mornhinweg, Daniel B. Shapiro, and K. G. Valente. The principal axis theorem over arbitrary fields. American Mathematical Monthly, 100(8):

749–754, October 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Needham:1993:VEJ

[345] Tristan Needham. A visual explanation of Jensen’s inequality. American Mathematical Monthly, 100(8): 768–771, October 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hassell:1993:ICC

[346] Catherine Hassell and Elmer Rees. The index of a constrained critical point. American Mathematical Monthly, 100 (8):772–778, October 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hegyvari:1993:SID

[347] Norbert Hegyvári. On some irrational decimal fractions. American Mathematical Monthly, 100(8):779–780, October 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brickman:1993:SPM

[348] Louis Brickman. The symmetry principle for Möbius transformations. American Mathematical Monthly, 100(8): 781–782, October 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Petersdorff:1993:SPR

[349] T. von Petersdorff. A short proof for Romberg integration. American Mathematical Monthly, 100(8):783–785, October 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nanyes:1993:EPB

[350] Ollie Nanyes. An elementary proof that the Borromean rings are nonsplittable. American Mathematical Monthly, 100 (8):786–789, October 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gardner:1993:TAHe

[351] J. Helen Gardner and Robin J. Wilson. Thomas Archer Hirst—mathematician xtravagant. V. London in the 1860s. American Mathematical Monthly, 100 (9):827–834, November 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anshel:1993:PMT

[352] Iris Lee Anshel and Michael Anshel. From the Post-Markov theorem through decision problems to public-key cryptography. American Mathematical Monthly, 100(9):835– 844, November 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Strang:1993:FTL

[353] Gilbert Strang. The fundamental theorem of linear algebra. American Mathematical Monthly, 100(9): 848–855, November 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dold:1993:SPJ

[354] Albrecht Dold. A simple proof of the Jordan-Alexander complement theorem. American Mathematical Monthly, 100(9):856–857, November 1993. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rummler:1993:SCH

[355] Hansklaus Rummler. Squaring the circle with holes. American Mathematical Monthly, 100(9):858–860, November 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Monsky:1993:SPD

[356] Paul Monsky. Simplifying the proof of Dirichlet’s theorem. American Mathematical Monthly, 100(9):861– 862, November 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maehara:1993:WE

[357] Hiroshi Maehara. Why is P^{2} not embeddable in R^{3}. American Mathematical Monthly, 100(9):862–864, November 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anderson:1993:PRD

[358] Bruce Anderson. Polynomial root dragging. American Mathematical Monthly, 100(9):864–866, November 1993. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gardner:1993:TAHf

[359] J. Helen Gardner and Robin J. Wilson. Thomas Archer Hirst—mathematician xtravagant. VI. Years of decline. American Mathematical Monthly, 100(10): 907–915, December 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Melissen:1993:DPC

[360] Hans Melissen. Densest packings of congruent circles in an equilateral triangle. American Mathematical Monthly, 100(10):916–925, December 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Yzeren:1993:SPP

[361] Jan van Yzeren. A simple proof of Pascal’s hexagon theorem. American Mathematical Monthly, 100(10): 930–931, December 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hyman:1993:MRB

[362] Andrew T. Hyman. The mathematical relationship between Kepler’s laws and Newton’s laws. American Mathematical Monthly, 100(10):932– 936, December 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gelca:1993:SPR

[363] Răzvan Gelca. A short proof of a result on polynomials. American Mathematical Monthly, 100(10):936– 937, December 1993. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Proctor:1993:TAD

[364] Robert A. Proctor. Two amusing Dynkin diagram graph classifications. American Mathematical Monthly, 100 (10):937–941, December 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Halmos:1993:PM

[365] Paul Halmos. Postcards from Max. American Mathematical Monthly, 100 (10):942–944, December 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1987:SPEb

[366] Donald E. Knuth. Solution to problem E3062, A versatile identity. American Mathematical Monthly, 94(3):376–377, March 1987. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1994:PKS

[367] Donald E. Knuth. Problem 103xx, a knight’s surprise. American Mathematical Monthly, 101(??):(to appear), 1994. ISSN 0002-9890 (print), 1930-0972 (electronic). Solution to appear.

Knuth:1994:SAPb

[368] Donald E. Knuth. Solution to advanced problem 6649, a generalized gamma function with independent branches. American Mathematical Monthly, 101 (??):(to appear), 1994. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [107], and other solutions in [369].

Knuth:1994:SAPa

[369] Donald E. Knuth. Solution to advanced problem 6649, a generalized gamma function with independent branches. American Mathematical Monthly, 101 (??):77–78, 1994. ISSN 0002-9890 (print), 1930-0972 (electronic). Original problem in [107], and other solutions in [368].

Cox:1994:IFL

[370] David A. Cox. Introduction to Fermat’s last theorem. American Mathematical Monthly, 101(1):3–14, January 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stillwell:1994:GTB

[371] John Stillwell. Galois theory for beginners. American Mathematical Monthly, 101(1):22–27, January 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pitman:1994:HRF

[372] E. Bruce Pitman. Into the hourglass: reflections on the forces acting on a granular material. American Mathematical Monthly, 101(1):28– 35, January 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jones:1994:OC

[373] John Dewey Jones. Orderly currencies. American Mathematical Monthly, 101 (1):36–38, January 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brown:1994:IFP

[374] Robert F. Brown and Robert E. Greene. An interior fixed point property of the disc. American Mathematical Monthly, 101(1):39–47, January 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Steele:1994:CIP

[375] J. Michael Steele. Le Cam’s inequality and Poisson approximations. American Mathematical Monthly, 101(1):48– 54, January 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mitrea:1994:GTE

[376] Dorina Mitrea and Marius Mitrea. A generalization of a theorem of Euler.

American Mathematical Monthly, 101 (1):55–58, January 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mironescu:1994:ETP

[377] Petru Mironescu and Laurenţiu Panaitopol. The existence of a triangle with prescribed angle bisector lengths. American Mathematical Monthly, 101 (1):58–60, January 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shenitzer:1994:EI

[378] A. Shenitzer and J. Steprāns. The evolution of integration. American Mathematical Monthly, 101(1):66–72, January 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ballew:1994:YGG

[379] David W. Ballew. Yueh-Gin Gung and Dr. Charles Y. Hu Award for Distinguished Service to J. Sutherland Frame. American Mathematical Monthly, 101(2):107–108, February 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grunbaum:1994:NLE

[380] Branko Grünbaum and G. C. Shephard. A new look at Euler’s theorem for polyhedra. American Mathematical Monthly, 101(2):109–128, February 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See comment and response [462].

Davis:1994:ONR

[381] Philip J. Davis. Otto Neugebauer: reminiscences and appreciation. American Mathematical Monthly, 101(2): 129–131, February 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kavinoky:1994:CNE

[382] Richard Kavinoky. Correction to: “From Newton to Einstein” [Amer. Math. Monthly 99 (1992), no. 6, 507– 521, MR 93c:01001] by B. Temple and C. A. Tracy. American Mathematical Monthly, 101(2):131, February 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [243].

Wegert:1994:BNP

[383] Elias Wegert and Lloyd N. Trefethen. From the Buffon needle problem to the Kreiss matrix theorem. American Mathematical Monthly, 101(2): 132–139, February 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Waterhouse:1994:CG

[384] William C. Waterhouse. A counterexample for Germain. American Mathematical Monthly, 101(2): 140–150, February 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Griffiths:1994:CEW

[385] H. B. Griffiths and A. E. Hirst. Cubic equations, or Where did the examination question come from? American Mathematical Monthly, 101(2): 151–161, February 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stachel:1994:IP

[386] Hellmuth Stachel. On the identity of polyhedra. American Mathematical Monthly, 101(2):162–165, February 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ullrich:1994:MPP

[387] David C. Ullrich. More on the Pompeiu problem. American Mathematical Monthly, 101(2):165–168, February 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Guy:1994:ENE

[388] Richard K. Guy. Every number is expressible as the sum of how many polygonal numbers. American Mathematical Monthly, 101(2): 169–172, February 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gouvea:1994:MP

[389] Fernando Q. Gouvêa. A marvelous proof. American Mathematical Monthly, 101(3):203–222, March 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Aldous:1994:TCR

[390] David Aldous. Triangulating the circle, at random. American Mathematical Monthly, 101(3):223–233, March 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Deakin:1994:HHM

[391] Michael A. B. Deakin. Hypatia and her mathematics. American Mathematical Monthly, 101(3):234–243, March 1994.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Barshinger:1994:CIE

[392] Richard Barshinger. Calculus II and Euler also (with a nod to series integral remainder bounds). American Mathematical Monthly, 101(3): 244–249, March 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Frantz:1994:FPE

[393] Marc Frantz. A focusing property of the ellipse. American Mathematical Monthly, 101(3):250–258, March 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Peirone:1994:RCT

[394] Roberto Peirone. Reflections can be trapped. American Mathematical Monthly, 101(3):259–260, March 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Heinrich:1994:ET

[395] Katherine Heinrich and Peter Horák. Euler’s theorem. American Mathematical Monthly, 101(3):260–261, March 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Feigenbaum:1994:UTS

[396] Joan Feigenbaum and Nick Reingold. Universal traversal sequences. American Mathematical Monthly, 101(3): 262–265, March 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stillwell:1994:WAI

[397] John Stillwell. What are algebraic integers and what are they for. American Mathematical Monthly, 101(3): 266–270, March 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bender:1994:PSP

[398] Edward A. Bender, Oren Patashnik, and Howard Rumsey, Jr. Pizza slicing, phi’s, and the Riemann hypothesis. American Mathematical Monthly, 101 (4):307–317, April 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Walde:1994:RPP

[399] Ralph Walde and Paula Russo. Rational periodic points of the quadratic function Q_{c}(x) = x^{2} + c. American Mathematical Monthly, 101(4): 318–331, April 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

G:1994:FVC

[400] Ernesto Acosta G. and Cesar Delgado G. Fréchet vs. carathéodory. American Mathematical Monthly, 101 (4):332–338, April 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Thompson:1994:OMP

[401] A. C. Thompson. Odd magic powers. American Mathematical Monthly, 101 (4):339–342, April 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morgan:1994:MIU

[402] Frank Morgan. Mathematicians, including undergraduates, look at soap bubbles. American Mathematical Monthly, 101(4):343–351, April 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Galvin:1994:PDC

[403] Fred Galvin. A proof of Dilworth’s chain decomposition theorem. American Mathematical Monthly, 101(4): 352–353, April 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Vellekoop:1994:ITC

[404] Michel Vellekoop and Raoul Berglund. On intervals, transitivity = chaos. American Mathematical Monthly, 101 (4):353–355, April 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kedlaya:1994:PMA

[405] Kiran Kedlaya. Proof of a mixed arithmetic-mean, geometric-mean inequality. American Mathematical Monthly, 101(4):355–357, April 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Banchoff:1994:GPC

[406] Thomas Banchoff and Peter Giblin. On the geometry of piecewise circular curves. American Mathematical Monthly, 101(5):403–416, May 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See errata [442].

Linzer:1994:TEP

[407] Elliot Linzer. The two envelope paradox. American Mathematical Monthly, 101(5):417–419, May 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maurer:1994:COC

[408] Stephen B. Maurer. Comment: “Orderly currencies” [Amer. Math. Monthly 101 (1994), no. 1, 36–38] by J. D. Jones. American Mathematical Monthly, 101(5):419, May 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Robert:1994:FSP

[409] Alain Robert. Fourier series of polygons. American Mathematical Monthly, 101(5):420–428, May 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Savage:1994:PND

[410] Richard P. Savage, Jr. The paradox of nontransitive dice. American Mathematical Monthly, 101(5):429–436, May 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beeckmans:1994:SES

[411] Laurent Beeckmans. Squares expressible as sum of consecutive squares. American Mathematical Monthly, 101 (5):437–442, May 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Turner:1994:SRM

[412] Stephen M. Turner. Square roots mod p. American Mathematical Monthly, 101(5):443–449, May 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tong:1994:KTG

[413] Jing Cheng Tong. Kummer’s test gives characterizations for convergence or divergence of all positive series. American Mathematical Monthly, 101 (5):450–452, May 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Li:1994:IN

[414] Chi-Kwong Li and Wasin So. Isometries of l_{p}-norm. American Mathematical Monthly, 101(5):452–453, May 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wang:1994:TIU

[415] Bo Ying Wang and Fu Zhen Zhang. A trace inequality for unitary matrices. American Mathematical Monthly, 101 (5):453–455, May 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grafakos:1994:EPS

[416] Loukas Grafakos. An elementary proof of the square summability of the discrete Hilbert transform. American Mathematical Monthly, 101(5): 456–458, May 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shenitzer:1994:HHG

[417] Abe Shenitzer. How hyperbolic geometry became respectable. American Mathematical Monthly, 101(5): 464–470, May 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Buhler:1994:JDD

[418] Joe Buhler, David Eisenbud, Ron Graham, and Colin Wright. Juggling drops and descents. American Mathematical Monthly, 101(6):507–519, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gordon:1994:WGA

[419] Gary Gordon. Workable gears, Archimedian solids and planar bipartite graphs. American Mathematical Monthly, 101(6):527–534, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Prosser:1994:KSH

[420] Reese T. Prosser. On the Kummer solutions of the hypergeometric equation. American Mathematical Monthly, 101 (6):535–543, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Emert:1994:RM

[421] John W. Emert, Kay I. Meeks, and Roger B. Nelson. Reflections on a Mira. American Mathematical Monthly, 101 (6):544–549, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Waymire:1994:BN

[422] Ed Waymire. Buffon noodles. American Mathematical Monthly, 101(6): 550–559, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bronstein:1994:CPC

[423] Victor Bronstein and Aviezri S. Fraenkel. On a curious property of counting sequences. American Mathematical Monthly, 101(6):560– 563, June/July 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knudsen:1994:CN

[424] Carsten Knudsen. Chaos without nonperiodicity. American Mathematical Monthly, 101(6):563–565, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pecaric:1994:RSI

[425] Josip Pečarić. A reverse Stolarsky’s inequality. American Mathematical Monthly, 101(6):565–567, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mercer:1994:NSI

[426] A. McD. Mercer. A note on some irrational decimal fractions. American Mathematical Monthly, 101(6): 567–568, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Callan:1994:PPF

[427] David Callan. A possible permanent formula. American Mathematical Monthly, 101(6):571–573, June/July 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

King:1994:TPS

[428] Jonathan L. King. Three problems in search of a measure. American Mathematical Monthly, 101(7):609–628, August/September 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rivin:1994:QP

[429] Igor Rivin, Ilan Vardi, and Paul Zimmermann. The n-queens problem. American Mathematical Monthly, 101 (7):629–639, August/September 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kraft:1994:WDB

[430] Roger L. Kraft. What’s the difference between Cantor sets. American Mathematical Monthly, 101(7):640–650, August/September 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Allouche:1994:MSS

[431] Jean-Paul Allouche, Dan Astoorian, Jim Randall, and Jeffrey Shallit. Morphisms, squarefree strings, and the Tower of Hanoi puzzle. American Mathematical Monthly, 101(7): 651–658, August/September 1994. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ageev:1994:STD

[432] A. A. Ageev. Sierpiński’s theorem is deducible from Euler and Dirichlet. American Mathematical Monthly, 101 (7):659–660, August/September 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Matus:1994:NSP

[433] F. Matúš. On nonnegativity of symmetric polynomials. American Mathematical Monthly, 101(7):661–664, August/September 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Graham:1994:NTO

[434] N. Graham, R. C. Entringer, and L. A. Székely. New tricks for old trees: maps and the pigeonhole principle. American Mathematical Monthly, 101 (7):664–667, August/September 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ondich:1994:DYK

[435] Jeffrey Ondich. Do you know the way to vertex A. American Mathematical Monthly, 101(7):668–673, August/September 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kreyszig:1994:CVMa

[436] Erwin Kreyszig. On the calculus of variations and its major influences on the mathematics of the first half of our century. I. American Mathematical Monthly, 101(7):674–678, August/September 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Terrell:1994:BSR

[437] Maria S. Terrell and Robert E. Terrell. Behind the scenes of a random dot stereogram. American Mathematical Monthly, 101(8):715–724, October 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shimura:1994:FTE

[438] Goro Shimura. Fractional and trigonometric expressions for matrices. American Mathematical Monthly, 101(8): 744–758, October 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Osofsky:1994:NLP

[439] Barbara L. Osofsky. Noether Lasker primary decomposition revisited. American Mathematical Monthly, 101(8):759–768, October 1994. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stein:1994:EIS

[440] S. Stein and S. Szabó. Elementary infinite sources of non-unique factorization rings. American Mathematical Monthly, 101(8):769–770, October 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fekete:1994:ATN

[441] Antal E. Fekete. Apropos “Two notes on notation” [Amer. Math. Monthly 99 (1992), no. 5, 403–422, MR 93f:05001] by D. E. Knuth. American Mathematical Monthly, 101(8):771–778, October 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [238].

Ramshaw:1994:EGP

[442] Lyle Ramshaw. Erratum: “On the geometry of piecewise circular curves” [Amer. Math. Monthly 101 (1994), no. 5, 403–416, MR 95k:51030a] by T. Banchoff and P. Giblin. American Mathematical Monthly, 101(8):778, October 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [406].

Gillman:1994:CT

[443] Leonard Gillman. Congruence of triangles. American Mathematical Monthly, 101(8):782–783, October 1994. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hungerbuhler:1994:SEP

[444] Norbert Hungerbühler. A short elementary proof of the Mohr-Mascheroni theorem. American Mathematical Monthly, 101(8):784–787, October 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eriksson:1994:WTP

[445] Folke Eriksson. Which triangles are plane sections of regular tetrahedra. American Mathematical Monthly, 101 (8):788–789, October 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Waymire:1994:ABN

[446] Edward C. Waymire. Addendum: “Buffon noodles”. American Mathematical Monthly, 101(8):791, October 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gray:1994:GCT

[447] Robert Gray. Georg Cantor and transcendental numbers. American Mathematical Monthly, 101(9):819– 832, November 1994. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sobel:1994:ASP

[448] Milton Sobel and Krzysztof Frankowski. The 500th anniversary of the sharing problem (the oldest problem in the theory of probability). American Mathematical Monthly, 101(9): 833–847, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Thurston:1994:WWD

[449] Hugh Thurston. What is wrong with the definition of dy/dx. American Mathematical Monthly, 101(9): 855–857, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gordon:1994:SAG

[450] Louis Gordon. A stochastic approach to the gamma function. American Mathematical Monthly, 101(9): 858–865, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goodman:1994:ATP

[451] Jacob E. Goodman, Richard Pollack, Rephael Wenger, and Tudor Zamfirescu. Arrangements and topological planes. American Mathematical Monthly, 101(9):866–878, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boyle:1994:AFS

[452] Jeff Boyle. An application of Fourier series to the most significant digit problem. American Mathematical Monthly, 101(9):879–886, November 1994. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dittmer:1994:CPI

[453] Andrew Dittmer. Cross product identities in arbitrary dimension. American Mathematical Monthly, 101(9): 887–891, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Armitage:1994:NCC

[454] D. H. Armitage. A non-constant continuous function on the plane whose integral on every line is zero. American Mathematical Monthly, 101(9): 892–894, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zeilberger:1994:CII

[455] Doron Zeilberger. Chu’s 1303 identity implies Bombieri’s 1990 norminequality (via an identity of Beauzamy and Dégot). American Mathematical Monthly, 101(9):894–896, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McGeoch:1994:HSC

[456] Catherine C. McGeoch. How to stay competitive. American Mathematical Monthly, 101(9):897–901, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kreyszig:1994:CVMb

[457] Erwin Kreyszig. On the calculus of variations and its major influences on the mathematics of the first half of our century. II. American Mathematical Monthly, 101(9):902–908, November 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Scheinerman:1994:RCN

[458] Edward R. Scheinerman and Herbert S. Wilf. The rectilinear crossing number of a complete graph and Sylvester’s “four point problem” of geometric probability. American Mathematical Monthly, 101(10):939–943, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Percus:1994:SMN

[459] Ora E. Percus and Jerome K. Percus. String matching for the novice. American Mathematical Monthly, 101 (10):944–947, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rawlings:1994:BTN

[460] Don Rawlings. Bernoulli trials and number theory. American Mathematical Monthly, 101(10):948–952, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Paulsen:1994:WSM

[461] William H. Paulsen. What is the shape of a mylar balloon. American Mathematical Monthly, 101(10): 953–958, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hilton:1994:ETP

[462] Peter Hilton and Jean Pedersen. Euler’s theorem for polyhedra: a topologist and geometer respond. comment to: “A new look at Euler’s theorem for polyhedra” [Amer. Math. Monthly 101 (1994), no. 2, 109–128, MR 95c:52032] by B. Grünbaum and G. C. Shephard. With a response by Grünbaum and Shephard. American Mathematical Monthly, 101(10):959–962, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [380].

Kleiner:1994:RPE

[463] I. Kleiner and N. Movshovitz-Hadar. The role of paradoxes in the evolution of mathematics. American Mathematical Monthly, 101(10):963–974, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brualdi:1994:RCP

[464] Richard A. Brualdi and Stephen Mellendorf. Regions in the complex plane containing the eigenvalues of a matrix. American Mathematical Monthly, 101 (10):975–985, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Spearman:1994:CSQ

[465] Blair K. Spearman and Kenneth S. Williams. Characterization of solvable quintics x^{5} + ax + b. American Mathematical Monthly, 101(10): 986–992, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cosgrave:1994:HPR

[466] John B. Cosgrave. A Halmos problem and a related problem. American Mathematical Monthly, 101(10): 993–996, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Girstmair:1994:PCN

[467] Kurt Girstmair. A “popular” class number formula. American Mathematical Monthly, 101(10):997–1001, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Alkan:1994:VWT

[468] Emre Alkan. Variations on Wolstenholme’s theorem. American Mathematical Monthly, 101(10):1001–1004, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gillman:1994:SPT

[469] Leonard Gillman. The second-partials test for local extrema of f(x, y). American Mathematical Monthly, 101(10): 1004–1006, December 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Orkin:1995:WWF

[470] Michael Orkin and Richard Kakigi. What is the worth of free casino credit. American Mathematical Monthly, 102 (1):3–8, January 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brams:1995:EFC

[471] Steven J. Brams and Alan D. Taylor. An envy-free cake division protocol. American Mathematical Monthly, 102(1):9–18, January 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hoffman:1995:DPT

[472] Michael E. Hoffman. Derivative polynomials for tangent and secant. American Mathematical Monthly, 102(1):23– 30, January 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Liggett:1995:LLN

[473] Thomas M. Liggett and Peter Petersen. The law of large numbers and \sqrt{2}. American Mathematical Monthly, 102 (1):31–35, January 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Redfield:1995:ER

[474] R. H. Redfield. Exponentiation in rings. American Mathematical Monthly, 102(1):36–40, January 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Peterson:1995:IHE

[475] Blake E. Peterson and James H. Jordan. Integer hexahedra equivalent to perfect boxes. American Mathematical Monthly, 102(1):41–45, January 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bagby:1995:CNP

[476] Richard J. Bagby. Calculating normal probabilities. American Mathematical Monthly, 102(1):46–48, January 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shutler:1995:CCP

[477] Paul Shutler. Constrained critical points. American Mathematical Monthly, 102(1):49–52, January 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tabachnikov:1995:CE

[478] S. Tabachnikov. A cone eversion. American Mathematical Monthly, 102 (1):52–56, January 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Libeskind-Hadas:1995:AAG

[479] Ran Libeskind-Hadas. Approximation algorithms: good solutions to hard problems. American Mathematical Monthly, 102(1):57–61, January 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mykytiuk:1995:FSA

[480] Stefan Mykytiuk and Abe Shenitzer. Four significant axiomatic systems and some of the issues associated with them. American Mathematical Monthly, 102(1):62–67, January 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Haimo:1995:ECE

[481] Deborah Tepper Haimo. Experimentation and conjecture are not enough. American Mathematical Monthly, 102 (2):102–112, February 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chung:1995:PC

[482] Fan Chung, Ron Graham, John Morrison, and Andrew Odlyzko. Pebbling a chessboard. American Mathematical Monthly, 102(2):113–123, February 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chapman:1995:DSS

[483] S. J. Chapman. Drums that sound the same. American Mathematical Monthly, 102(2):124–138, February 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Axler:1995:D

[484] Sheldon Axler. Down with determinants. American Mathematical Monthly, 102(2):139–154, February 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Andrews:1995:WFC

[485] Peter Andrews. Where not to find the critical points of a polynomial— variation on a Putnam theme. American Mathematical Monthly, 102(2): 155–158, February 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lax:1995:SPS

[486] Peter D. Lax. A short path to the shortest path. American Mathematical Monthly, 102(2):158–159, February 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Khavinson:1995:NES

[487] Dmitry Khavinson. A note on entire solutions of the eiconal [eikonal] equation. American Mathematical Monthly, 102 (2):159–161, February 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Surowski:1995:UAF

[488] David B. Surowski. The uniqueness aspect of the fundamental theorem of finite abelian groups. American Mathematical Monthly, 102(2): 162–163, February 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rabinowitz:1995:SAD

[489] Stanley Rabinowitz and Stan Wagon. A spigot algorithm for the digits of π. American Mathematical Monthly, 102(3):195–203, March 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Davis:1995:RFP

[490] Philip J. Davis. The rise, fall, and possible transfiguration of triangle geometry: a mini-history. American Mathematical Monthly, 102(3): 204–214, March 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Auckly:1995:TRO

[491] David Auckly and John Cleveland. Totally real origami and impossible paper folding. American Mathematical Monthly, 102(3):215–226, March 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Furi:1995:MVR

[492] Massimo Furi and Mario Martelli. A multidimensional version of Rolle’s theorem. American Mathematical Monthly, 102(3):243–249, March 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Alon:1995:ADC

[493] Noga Alon, Melvyn B. Nathanson, and Imre Ruzsa. Adding distinct congruence classes modulo a prime. American Mathematical Monthly, 102(3): 250–255, March 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maligranda:1995:SPH

[494] Lech Maligranda. A simple proof of the Hölder and the Minkowski inequality. American Mathematical Monthly, 102(3):256–259, March 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kaplansky:1995:NTL

[495] Irving Kaplansky. Never too late. addendum to: “Determinants whose elements are 0 and 1” [Amer. Math. Monthly 53 (1946), 427–434, MR 8, 128] by J. Williamson. American Mathematical Monthly, 102(3):259, March 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [1].

Bashmakova:1995:EA

[496] I. G. Bashmakova and A. N. Rudakov. The evolution of algebra 1800–1870. American Mathematical Monthly, 102 (3):266–270, March 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hausrath:1995:GRT

[497] Alan Hausrath, Bradley Jackson, John Mitchem, and Edward Schmeichel. Gale’s round-trip jeep problem. American Mathematical Monthly, 102(4): 299–309, April 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Weintraub:1995:CWM

[498] Steven H. Weintraub. Count-wheels: a mathematical problem arising in horology. American Mathematical Monthly, 102(4):310–316, April 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chalice:1995:HTC

[499] Donald R. Chalice. How to teach a class by the modified Moore method. American Mathematical Monthly, 102 (4):317–321, April 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hill:1995:SDP

[500] Theodore P. Hill. The significant-digit phenomenon. American Mathematical Monthly, 102(4):322–327, April 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Haws:1995:EBP

[501] LaDawn Haws and Terry Kiser. Exploring the brachistochrone problem. American Mathematical Monthly, 102 (4):328–336, April 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Derrick:1995:CFC

[502] William Derrick and Jack Eidswick. Continued fractions, Chebychev polynomials, and chaos. American Mathematical Monthly, 102(4):337–344, April 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wang:1995:RBP

[503] Zheng Bing Wang, Robert Fokkink, and Wan Fokkink. A relation between partitions and the number of divisors. American Mathematical Monthly, 102 (4):345–346, April 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Starni:1995:ATQ

[504] Paolo Starni. Answers to two questions concerning quotients of primes. American Mathematical Monthly, 102 (4):347–349, April 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ford:1995:AEL

[505] James W. M. Ford. Avoiding the exchange lemma. American Mathematical Monthly, 102(4):350–351, April 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Silverman:1995:ICA

[506] Stephen Silverman. Intervals contained in arithmetic combinations of sets. American Mathematical Monthly, 102(4):351–353, April 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Flath:1995:DMF

[507] D. Flath and A. Zulauf. Does the Möbius function determine multiplicative arithmetic. American Mathematical Monthly, 102(4):354–356, April 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grattan-Guinness:1995:WDG

[508] I. Grattan-Guinness. Why did George Green write his essay of 1828 on electricity and magnetism. American Mathematical Monthly, 102(5): 387–396, May 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bassein:1995:OCY

[509] Susan Bassein. Order and chaos on your desk. American Mathematical Monthly, 102(5):409–416, May 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Scavo:1995:GHM

[510] T. R. Scavo and J. B. Thoo. On the geometry of Halley’s method. American Mathematical Monthly, 102(5): 417–426, May 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Meijer:1995:BE

[511] A. R. Meijer. The binary expansion of 1/p. American Mathematical Monthly, 102(5):427–430, May 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Diaz:1995:PFW

[512] Ricardo Diaz and Sinai Robins. Pick’s formula via the Weierstrass ℘-function. American Mathematical Monthly, 102 (5):431–437, May 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mackiw:1995:PPT

[513] George Mackiw. Permutations as products of transpositions. American Mathematical Monthly, 102(5):438–440, May 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Poonen:1995:CRO

[514] Bjorn Poonen. Congruences relating the order of a group to the number of conjugacy classes. American Mathematical Monthly, 102(5):440–442, May 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Andersson:1995:CIK

[515] Peter Andersson. The color invariant for knots and links. American Mathematical Monthly, 102(5):442–448, May 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Weyl:1995:TAAa

[516] Hermann Weyl. Topology and abstract algebra as two roads of mathematical comprehension. I. translated from the German by Abe Shenitzer. American Mathematical Monthly, 102 (5):453–460, May 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zabell:1995:ATC

[517] S. L. Zabell. Alan Turing and the central limit theorem. American Mathematical Monthly, 102(6): 483–494, June/July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rosen:1995:NHA

[518] Michael I. Rosen. Niels Hendrik Abel and equations of the fifth degree. American Mathematical Monthly, 102 (6):495–505, June/July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Oprea:1995:GFP

[519] John Oprea. Geometry and the Foucault pendulum. American Mathematical Monthly, 102(6):515–522, June/ July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Robbins:1995:API

[520] David P. Robbins. Areas of polygons inscribed in a circle. American Mathematical Monthly, 102(6): 523–530, June/July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Scofield:1995:CCP

[521] Paul D. Scofield. Curves of constant precession. American Mathematical Monthly, 102(6):531–537, June/July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Allen:1995:MEO

[522] Jeffery C. Allen. Matrix expansion by orthogonal Kronecker products. American Mathematical Monthly, 102(6): 538–540, June/July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rudin:1995:IPM

[523] Walter Rudin. Injective polynomial maps are automorphisms. American Mathematical Monthly, 102(6): 540–543, June/July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chapman:1995:EPS

[524] Robin J. Chapman. An elementary proof of the simplicity of the Mathieu groups M_{11} and M_{23}. American Mathematical Monthly, 102(6): 544–545, June/July 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamalainen:1995:FPG

[525] Heikki Hämäläinen, Iiro Honkala, Simon Litsyn, and Patric Östergård. Football pools — a game for mathematicians. American Mathematical Monthly, 102(7):579–588, August/ September 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hersh:1995:FBP

[526] Reuben Hersh. Fresh breezes in the philosophy of mathematics. American Mathematical Monthly, 102(7): 589–594, August/September 1995. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boston:1995:QRP

[527] Nigel Boston and Marshall L. Greenwood. Quadratics representing primes. American Mathematical Monthly, 102 (7):595–599, August/September 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lamport:1995:HWP

[528] Leslie Lamport. How to write a proof. American Mathematical Monthly, 102 (7):600–608, August/September 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stolarsky:1995:SCG

[529] Kenneth B. Stolarsky. Searching for common generalizations: the case of hyperbolic functions. American Mathematical Monthly, 102(7):609–619, August/September 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sabinin:1995:TGF

[530] P. Sabinin and M. G. Stone. Transforming n-gons by folding the plane. American Mathematical Monthly, 102 (7):620–627, August/September 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Singer:1995:IP

[531] David A. Singer. Isometries of the plane. American Mathematical Monthly, 102(7):628–631, August/ September 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Geyler:1995:OMC

[532] V. A. Geyler. One more construction which is impossible. American Mathematical Monthly, 102(7):632–634, August/September 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Matsuda:1995:IPM

[533] Takashi Matsuda. An inductive proof of a mixed arithmetic-geometric mean inequality. American Mathematical Monthly, 102(7):634–637, August/September 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Michael:1995:RTM

[534] T. S. Michael. The ranks of tournament matrices. American Mathematical Monthly, 102(7):637–639, August/September 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ranum:1995:SAF

[535] David L. Ranum. On some applications of Fibonacci numbers. American Mathematical Monthly, 102(7):640–645, August/September 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Weyl:1995:TAAb

[536] Hermann Weyl. Topology and abstract algebra as two roads of mathematical comprehension. II. translated from the German by Abe Shenitzer. American Mathematical Monthly, 102 (7):646–651, August/September 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Myerson:1995:SPC

[537] G. Myerson and A. J. van der Poorten. Some problems concerning recurrence sequences. American Mathematical Monthly, 102(8):698–705, October 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mackenzie:1995:HPC

[538] Dana Mackenzie. A hyperbolic plane coloring and the simple group of order 168. American Mathematical Monthly, 102(8):706–715, October 1995. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morrison:1995:CPF

[539] Kent E. Morrison. Cosine products, Fourier transforms, and random sums. American Mathematical Monthly, 102 (8):716–724, October 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Schay:1995:HAF

[540] Geza Schay. How to add fast — on average. American Mathematical Monthly, 102(8):725–730, October 1995. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morton:1995:FLS

[541] H. R. Morton. Fibonacci-like sequences and greatest common divisors. American Mathematical Monthly, 102(8): 731–734, October 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Isaacs:1995:GSG

[542] I. M. Isaacs and Thilo Zieschang. Generating symmetric groups. American Mathematical Monthly, 102(8): 734–739, October 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lucht:1995:AGM

[543] Lutz G. Lucht. On the arithmeticgeometric mean inequality. American Mathematical Monthly, 102(8): 739–740, October 1995. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Guy:1995:MFE

[544] Richard K. Guy. My favorite elliptic curve: a tale of two types of triangles. American Mathematical Monthly, 102 (9):771–781, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kendig:1995:SWE

[545] Keith M. Kendig. Stalking the wild ellipse. American Mathematical Monthly, 102(9):782–787, November 1995. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Crannell:1995:RTD

[546] Annalisa Crannell. The role of transitivity in Devaney’s definition of chaos.

American Mathematical Monthly, 102 (9):788–797, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Poole:1995:SG

[547] David G. Poole. The stochastic group. American Mathematical Monthly, 102 (9):798–801, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sander:1995:SBC

[548] J. W. Sander. A story of binomial coefficients and primes. American Mathematical Monthly, 102(9): 802–807, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Aigner:1995:TGT

[549] Martin Aigner. Turán’s graph theorem. American Mathematical Monthly, 102 (9):808–816, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Samelson:1995:MKT

[550] Hans Samelson. More on Kummer’s test. American Mathematical Monthly, 102(9):817–818, November 1995. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tuynman:1995:DEM

[551] G. M. Tuynman. The derivation of the exponential map of matrices. American Mathematical Monthly, 102(9): 818–820, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ptak:1995:KI

[552] Vlastimil Pták. The Kantorovich inequality. American Mathematical Monthly, 102(9):820–821, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kalaba:1995:GIF

[553] Robert Kalaba and Rong Xu. On the generalized inverse form of the equations of constrained motion. American Mathematical Monthly, 102(9): 821–825, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stillwell:1995:EC

[554] John Stillwell. Elliptic curves. American Mathematical Monthly, 102(9): 831–837, November 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tokarsky:1995:PRI

[555] George W. Tokarsky. Polygonal rooms not illuminable from every point. American Mathematical Monthly, 102 (10):867–879, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Blom:1995:TSS

[556] Gunnar Blom, Lars Holst, and Dennis Sandell. Three SING-SING problems. American Mathematical Monthly, 102 (10):880–887, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morrill:1995:NPM

[557] John E. Morrill. A Nobel Prize in mathematics. American Mathematical Monthly, 102(10):888–891, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Blecksmith:1995:SEN

[558] Richard Blecksmith and Purushottam W. Laud. Some exact number theory computations via probability mechanisms. American Mathematical Monthly, 102(10):893–903, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ipsen:1995:ABC

[559] Ilse C. F. Ipsen and Carl D. Meyer. The angle between complementary subspaces. American Mathematical Monthly, 102(10):904–911, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tabachnikov:1995:FVT

[560] Serge Tabachnikov. The four-vertex theorem revisited — two variations on the old theme. American Mathematical Monthly, 102(10):912–916, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Burckel:1995:EFW

[561] R. B. Burckel. Entire functions which vanish at infinity. American Mathematical Monthly, 102(10):916–918, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zagier:1995:CCI

[562] D. Zagier. A converse to Cauchy’s inequality. American Mathematical Monthly, 102(10):919–920, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Guy:1995:MUP

[563] Richard K. Guy and Richard J. Nowakowski. Monthly unsolved problems, 1969–1995. American Mathematical Monthly, 102(10):921–926, December 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fowler:1996:BCF

[564] David Fowler. The binomial coefficient function. American Mathematical Monthly, 103(1):1–17, January 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Henle:1996:CMB

[565] Jim Henle. Classical mathematics. Baroque mathematics. Romantic mathematics? Mathematics jazz! also atonal, New Age, minimalist, and punk mathematics. American Mathematical Monthly, 103(1):18–29, January 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Solomon:1996:TSC

[566] Bruce Solomon. Tantrices of spherical curves. American Mathematical Monthly, 103(1):30–39, January 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:1996:GCP

[567] D. Borwein, J. M. Borwein, P. B. Borwein, and R. Girgensohn. Giuga’s conjecture on primality. American Mathematical Monthly, 103(1):40–50, January 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Faucette:1996:GIS

[568] William M. Faucette. A geometric interpretation of the solution of the general quartic polynomial. American Mathematical Monthly, 103(1):51– 57, January 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Codenotti:1996:TCA

[569] Bruno Codenotti and Luciano Margara. Transitive cellular automata are sensitive. American Mathematical Monthly, 103(1):58–62, January 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Redheffer:1996:UDI

[570] Ray Redheffer. Use of a differential inequality in combinatorics. American Mathematical Monthly, 103(1):62– 64, January 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tikhomirov:1996:EMC

[571] V. M. Tikhomirov. The evolution of methods of convex optimization. American Mathematical Monthly, 103 (1):65–71, January 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pollak:1996:YGG

[572] H. O. Pollak. Yueh-Gin Gung and Dr. Charles Y. Hu Award for Distinguished Service to Andrew Gleason. American Mathematical Monthly, 103 (2):105–106, February 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sandefur:1996:USS

[573] James T. Sandefur. Using selfsimilarity to find length, area, and dimension. American Mathematical Monthly, 103(2):107–120, February 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hilton:1996:ECP

[574] Peter Hilton and Jean Pedersen. The Euler characteristic and Pólya’s dream. American Mathematical Monthly, 103 (2):121–131, February 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Artmann:1996:RID

[575] Benno Artmann. A Roman icosahedron discovered. American Mathematical Monthly, 103(2):132–133, February 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Barros:1996:FEP

[576] Manuel Barros and Oscar J. Garay. Free elastic parallels in a surface of revolution. American Mathematical Monthly, 103(2):149–156, February 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gohberg:1996:SPJ

[577] Israel Gohberg and Seymour Goldberg. A simple proof of the Jordan decomposition theorem for matrices. American Mathematical Monthly, 103 (2):157–159, February 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Caruth:1996:CPH

[578] A. Caruth. A concise proof of Hilbert’s basis theorem. American Mathematical Monthly, 103(2):160–161, February 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ferrer:1996:RTF

[579] Jesús Ferrer. Rolle’s theorem fails in l_{2}. American Mathematical Monthly, 103 (2):161–165, February 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Paun:1996:SRS

[580] Gheorghe Păun and Arto Salomaa. Self-reading sequences. American Mathematical Monthly, 103(2):166– 168, February 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beardon:1996:SPI

[581] A. F. Beardon. Sums of powers of integers. American Mathematical Monthly, 103(3):201–213, March 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Remmert:1996:WTA

[582] Reinhold Remmert. Wielandt’s theorem about the Γ-function. American Mathematical Monthly, 103(3): 214–220, March 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zulli:1996:CSE

[583] Louis Zulli. Charting the 3-sphere — an exposition for undergraduates. American Mathematical Monthly, 103 (3):221–229, March 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zeilberger:1996:MUG

[584] Doron Zeilberger. The method of undetermined generalization and specialization. illustrated with Fred Galvin’s amazing proof of the Dinitz conjecture. American Mathematical Monthly, 103(3):233–239, March 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hull:1996:NIP

[585] Thomas Hull. A note on “impossible” paper folding. American Mathematical Monthly, 103(3):240–241, March 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lawlor:1996:NMP

[586] Gary Lawlor. A new minimization proof for the brachistochrone. American Mathematical Monthly, 103(3): 242–249, March 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mansfield:1996:SSS

[587] Richard Mansfield. Strategies for the Shannon switching game. American Mathematical Monthly, 103(3): 250–252, March 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Porter:1996:VGP

[588] Gerald J. Porter. k-volume in R^{n} and the generalized Pythagorean theorem. American Mathematical Monthly, 103(3):252–256, March 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shenitzer:1996:FEM

[589] Abe Shenitzer. A few expository mini-essays. American Mathematical Monthly, 103(3):257–263, March 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Askey:1996:SBF

[590] Richard Askey and Deborah Tepper Haimo. Similarities between Fourier and power series. American Mathematical Monthly, 103(4):297–304, April 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Warren:1996:DC

[591] Richard H. Warren. Disks on a chessboard. American Mathematical Monthly, 103(4):305–307, April 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eisenberg:1996:RTD

[592] Bennett Eisenberg and Rosemary Sullivan. Random triangles in n dimensions. American Mathematical Monthly, 103 (4):308–318, April 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Oldknow:1996:EGS

[593] Adrian Oldknow. The Euler-Gergonne-Soddy triangle of a triangle. American Mathematical Monthly, 103(4): 319–329, April 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hartwig:1996:RRR

[594] Robert E. Hartwig. Roth’s removal rule and the rational canonical form. American Mathematical Monthly, 103 (4):332–335, April 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Krause:1996:SPG

[595] Manfred Krause. A simple proof of the Gale-Ryser theorem. American Mathematical Monthly, 103(4):335–337, April 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Martinez-Maure:1996:NTB

[596] Yves Martinez-Maure. A note on the tennis ball theorem. American Mathematical Monthly, 103(4):338–340, April 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morgan:1996:WS

[597] Frank Morgan. What is a surface. American Mathematical Monthly, 103 (5):369–376, May 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Perline:1996:NEF

[598] Ron Perline. Non-Euclidean flashlights. American Mathematical Monthly, 103 (5):377–385, May 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Elaydi:1996:CST

[599] Saber Elaydi. On a converse of Sharkovsky’s theorem. American Mathematical Monthly, 103(5):386– 392, May 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eriksson:1996:SPT

[600] Kimmo Eriksson. Splitting a polygon into two congruent pieces. American Mathematical Monthly, 103(5): 393–400, May 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hannah:1996:GAD

[601] John Hannah. A geometric approach to determinants. American Mathematical Monthly, 103(5):401–409, May 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Desbrow:1996:ZD

[602] Darrell Desbrow. On zero derivatives. American Mathematical Monthly, 103 (5):410–411, May 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Duren:1996:APH

[603] Peter Duren, Walter Hengartner, and Richard S. Laugesen. The argument principle for harmonic functions. American Mathematical Monthly, 103 (5):411–415, May 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Darst:1996:DPO

[604] Richard B. Darst and Gerald D. Taylor. Differentiating powers of an old friend. American Mathematical Monthly, 103 (5):415–416, May 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kleiner:1996:GAR

[605] Israel Kleiner. The genesis of the abstract ring concept. American Mathematical Monthly, 103(5):417–424, May 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gottlieb:1996:AWG

[606] Daniel Henry Gottlieb. All the way with Gauss–Bonnet and the sociology of mathematics. American Mathematical Monthly, 103(6):457–469, June/ July 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Horn:1996:WDW

[607] Roger A. Horn and Ingram Olkin. When does A^{∗}A = B^{∗}B and why does one want to know. American Mathematical Monthly, 103(6): 470–482, June/July 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bassein:1996:SR

[608] Susan Bassein. A sampler of randomness. American Mathematical Monthly, 103(6):483–490, June/July 1996. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fallis:1996:MPR

[609] Don Fallis. Mathematical proof and the reliability of DNA evidence. American Mathematical Monthly, 103(6): 491–497, June/July 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jepsen:1996:ET

[610] Charles H. Jepsen. Equidissections of trapezoids. American Mathematical Monthly, 103(6):498–500, June/July 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Monsky:1996:CTS

[611] Paul Monsky. Calculating a trapezoidal spectrum. American Mathematical Monthly, 103(6):500–501, June/July 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zeilberger:1996:RCA

[612] Doron Zeilberger. Reverend Charles to the aid of Major Percy and Fields medalist Enrico. American Mathematical Monthly, 103(6):501–502, June/ July 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bukor:1996:APR

[613] József Bukor and János T. Tóth. On accumulation points of ratio sets of positive integers. American Mathematical Monthly, 103(6):502–504, June/July 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ribenboim:1996:CC

[614] Paulo Ribenboim. Catalan’s conjecture. American Mathematical Monthly, 103(7):529–538, August/September 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pinkham:1996:MMT

[615] Roger S. Pinkham. Mathematics and modern technology. American Mathematical Monthly, 103(7):539–545, August/September 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Groetsch:1996:TIP

[616] C. W. Groetsch. Tartaglia’s inverse problem in a resistive medium.

American Mathematical Monthly, 103(7):546–551, August/September 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2974665.

Minassian:1996:CSA

[617] Donald P. Minassian. The current state of actuarial science. American Mathematical Monthly, 103(7):552–561, August/September 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lam:1996:CP

[618] T. Y. Lam and K. H. Leung. On the cyclotomic polynomial Φ_{pq}(X). American Mathematical Monthly, 103(7): 562–564, August/September 1996. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Burger:1996:DRC

[619] Edward B. Burger and Thomas Struppeck. Does ∑_{n=0}^{∞} 1/n! really converge? Infinite series and p-adic analysis or “You can sum some of the series some of the time and some of the series none of the time… but can you sum some of the series all of the time?”. American Mathematical Monthly, 103 (7):565–577, August/September 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Johnson:1996:RB

[620] Charles R. Johnson and Erik A. Schreiner. The relationship between AB and BA. American Mathematical Monthly, 103(7):578–582, August/September 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fejzic:1996:TSC

[621] Hajrudin Fejzić. On thin sets of circles. American Mathematical Monthly, 103 (7):582–585, August/September 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Alzer:1996:PAM

[622] Horst Alzer. A proof of the arithmetic mean-geometric mean inequality. American Mathematical Monthly, 103(7):585, August/September 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Doob:1996:DRM

[623] Joseph L. Doob. The development of rigor in mathematical probability (1900–1950) [in Development of mathematics 1900–1950 (Luxembourg, 1992), 157–170, Birkhäuser, Basel, 1994, MR 95i:60001]. American Mathematical Monthly, 103(7):586–595, August/September 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bartle:1996:RRI

[624] Robert G. Bartle. Return to the Riemann integral. American Mathematical Monthly, 103(8):625–632, October 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Neidinger:1996:RCF

[625] Richard D. Neidinger and R. John Annen, III. The road to chaos is filled with polynomial curves. American Mathematical Monthly, 103(8):

640–653, October 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bedding:1996:IQF

[626] Stephen Bedding and Keith Briggs. Iteration of quaternion functions. American Mathematical Monthly, 103(8): 654–664, October 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ahern:1996:GPG

[627] Patrick Ahern and Walter Rudin. Geometric properties of the gamma function. American Mathematical Monthly, 103(8):678–681, October 1996. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Saeki:1996:PEI

[628] Sadahiro Saeki. A proof of the existence of infinite product probability measures. American Mathematical Monthly, 103(8):682–683, October 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gurin:1996:P

[629] Leo S. Gurin. A problem. American Mathematical Monthly, 103(8): 683–686, October 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bateman:1996:HYP

[630] Paul T. Bateman and Harold G. Diamond. A hundred years of prime numbers. American Mathematical Monthly, 103(9):729–741, November 1996. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Burger:1996:PPP

[631] Thomas Burger, Peter Gritzmann, and Victor Klee. Polytope projection and projection polytopes. American Mathematical Monthly, 103(9): 742–755, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jurisic:1996:MKP

[632] Aleksandar Jurišić. The Mercedes knot problem. American Mathematical Monthly, 103(9):756–770, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

White:1996:FSF

[633] Arthur T. White. Fabian Stedman: the first group theorist. American Mathematical Monthly, 103(9): 771–778, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brown:1996:FWS

[634] Herbert I. Brown, James W. Burgmeier, and David S. Dummit. Functions whose successive tangent lines enclose proportional areas. American Mathematical Monthly, 103(9):779– 787, November 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eggar:1996:DCG

[635] M. H. Eggar. On the discriminant criterion and a generalization. American Mathematical Monthly, 103(9): 788–792, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Snapper:1996:EPQ

[636] Ernst Snapper. Expressing primes as quadratic forms of integers. American Mathematical Monthly, 103(9): 793–794, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bagby:1996:VC

[637] Richard J. Bagby. Volumes of cones. American Mathematical Monthly, 103 (9):794–796, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bertram:1996:RBF

[638] Edward Bertram and Marcel Herzog. On regular bases of finite groups. American Mathematical Monthly, 103 (9):796–799, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dyson:1996:SR

[639] Freeman Dyson. The scientist as rebel. American Mathematical Monthly, 103 (9):800–805, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McMurran:1996:MCM

[640] Shawnee L. McMurran and James J. Tattersall. The mathematical collaboration of M. L. Cartwright and J. E. Littlewood. American Mathematical Monthly, 103(10):833–845, December 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kossak:1996:WIW

[641] Roman Kossak. What are infinitesimals and why they cannot be seen.

American Mathematical Monthly, 103 (10):846–853, December 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brillhart:1996:CSM

[642] John Brillhart and Patrick Morton. A case study in mathematical research: the Golay–Rudin–Shapiro sequence. American Mathematical Monthly, 103 (10):854–869, December 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Aboufadel:1996:MCB

[643] Edward Aboufadel. A mathematician catches a baseball. American Mathematical Monthly, 103(10):870– 878, December 1996. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Powers:1996:HPC

[644] Victoria Powers. Hilbert’s 17th problem and the champagne problem. American Mathematical Monthly, 103 (10):879–887, December 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McCune:1996:CHR

[645] W. McCune and A. D. Sands. Computer and human reasoning: single implicative axioms for groups and for abelian groups. American Mathematical Monthly, 103(10):888–892, December 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Webster:1996:EPH

[646] Roger Webster. An elementary proof of Horn’s theorem. American Mathematical Monthly, 103(10):892–894, December 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Angelos:1996:PFS

[647] James Angelos, George Grossman, Yury Ionin, Edwin Kaufman, Terry Lenker, and Leela Rakesh. Packability of five spheres on a sphere implies packability of six. American Mathematical Monthly, 103(10):894–896, December 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pourciau:1997:RMN

[648] Bruce Pourciau. Reading the master: Newton and the birth of celestial mechanics. American Mathematical Monthly, 104(1):1–19, January 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_jan97_toc.html.

Rogers:1997:ALA

[649] Jack W. Rogers, Jr. Applications of linear algebra in calculus. American Mathematical Monthly, 104(1):20–26, January 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_jan97_toc.html.

Vince:1997:PQB

[650] A. Vince. Periodicity, quasiperiodicity, and Bieberbach’s theorem on crystallographic groups. American Mathematical Monthly, 104(1):27–35, January 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_jan97_toc.html.

Baker:1997:ISD

[651] John A. Baker. Integration over spheres and the divergence theorem for balls. American Mathematical Monthly, 104(1):36–47, January 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_jan97_toc.html.

Burckel:1997:TSA

[652] R. B. Burckel. Three secrets about harmonic functions. American Mathematical Monthly, 104(1):52–56, January 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_jan97_toc.html.

Komornik:1997:SPE

[653] Vilmos Komornik. A short proof of the Erdős–Mordell theorem. American Mathematical Monthly, 104(1):57– 60, January 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_jan97_toc.html.

Buckholtz:1997:IDH

[654] Don Buckholtz. Inverting the difference of Hilbert space projections. American Mathematical Monthly, 104(1):60– 61, January 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_jan97_toc.html.

Bashmakova:1997:GAG

[655] I. G. Bashmakova and E. I. Slavutin. Glimpses of algebraic geometry. Introduction to: A History of Diophantine analysis (Russian) [“Nauka”, Moscow, 1984; MR 85m:01004]. Translated by Abe Shenitzer. American Mathematical Monthly, 104(1):62–67, January 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_jan97_toc.html. In-

troduction to: History of Diophantine analysis (Russian) [“Nauka”, Moscow, 1984; MR 85m:01004], Translated by Abe Shenitzer.

Lacampagne:1997:YGG

[656] Carole B. Lacampagne. Yueh-Gin Gung and Dr. Charles Y. Hu award for distinguished service to Deborah Tepper Haimo. American Mathematical Monthly, 104(2):97–98, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_feb97_toc.html.

Shader:1997:SCF

[657] Bryan L. Shader and Chanyoung Lee Shader. Scheduling conflict-free parties for a dating service. American Mathematical Monthly, 104(2):99–106, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_feb97_toc.html.

Buchholz:1997:ISH

[658] Ralph H. Buchholz and Randall L. Rathbun. An infinite set of Heron triangles with two rational medians. American Mathematical Monthly, 104 (2):107–115, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb97_
toc.html.

Lam:1997:CS

[659] T. K. Lam. Connected Sprouts. American Mathematical Monthly, 104(2): 116–119, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_feb97_
                                        toc.html.

Prather:1997:REP

[660] Ronald E. Prather. Regular expressions for program computations. American Mathematical Monthly, 104(2): 120–130, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_feb97_
                                        toc.html.

Richter:1997:RBC

[661] R. Bruce Richter and Carsten Thomassen. Relations between crossing numbers of complete and complete bipartite graphs. American Mathematical Monthly, 104(2):131–137, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_feb97_toc.html.

Holte:1997:CCA

[662] John M. Holte. Carries, combinatorics, and an amazing matrix. American Mathematical Monthly, 104(2): 138–149, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_feb97_
                                        toc.html.

Nef:1997:NLE

[663] Walter Nef. A new look at Euler’s theorem for polyhedra: a comment.

American Mathematical Monthly, 104 (2):150–151, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb97_
toc.html.

Tokieda:1997:HH

[664] Tadashi F. Tokieda. The hopping hoop. American Mathematical Monthly, 104(2):152–154, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_feb97_toc.html.

Greene:1997:PID

[665] John Greene. Principal ideal domains are almost Euclidean. American Mathematical Monthly, 104(2): 154–156, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb97_
toc.html.

Onn:1997:CDI

[666] Shmuel Onn. A colorful determinantal identity, a conjecture of Rota, and Latin squares. American Mathematical Monthly, 104(2):156–159, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_feb97_toc.html.

Nicholson:1997:VSM

[667] W. K. Nicholson. Very semisimple modules. American Mathematical Monthly, 104(2):159–162, February 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_feb97_toc.html.

Pitman:1997:SPA

[668] Jim Pitman. Some probabilistic aspects of set partitions. American Mathematical Monthly, 104(3):201– 209, March 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_mar97_toc.html.

Parshall:1997:BIR

[669] Karen Hunger Parshall and Eugene Seneta. Building an international reputation: the case of J. J. Sylvester (1814–1897). American Mathematical Monthly, 104(3):210–222, March 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_mar97_toc.html.

Goodson:1997:ISP

[670] Geoffrey R. Goodson. The inversesimilarity problem for real orthogonal matrices. American Mathematical Monthly, 104(3):223–230, March 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_mar97_toc.html.

Tucker:1997:RRC

[671] Thomas W. Tucker. Rethinking rigor in calculus: the role of the mean value theorem. American Mathematical Monthly, 104(3):231–240, March 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_mar97_toc.html.

Swann:1997:CRR

[672] Howard Swann. Commentary on rethinking rigor in calculus: the role of the mean value theorem. American Mathematical Monthly, 104(3): 241–245, March 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_mar97_toc.
html.

Burger:1997:FLT

[673] Edward B. Burger and Frank Morgan. Fermat’s Last Theorem, the Four Color Conjecture, and Bill Clinton for April Fools’ Day. American Mathematical Monthly, 104(3):246–255, March 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_mar97_toc.html.

Newman:1997:EFA

[674] D. J. Newman. Euler’s ϕ function on arithmetic progressions. American Mathematical Monthly, 104(3): 256–257, March 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_mar97_toc.
html.

Fisk:1997:NWI

[675] Steve Fisk. A note on Weyl’s inequality. American Mathematical Monthly, 104(3):257–258, March 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_mar97_toc.html.

Kang:1997:MPC

[676] Ming chang Kang. Minimal polynomials over cyclotomic fields. American Mathematical Monthly, 104(3): 258–260, March 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        maa.org/pubs/monthly_mar97_toc.
                                        html.

Magnus:1997:SMM

[677] Wilhelm Magnus. The significance of mathematics: the mathematicians’ share in the general human condition. American Mathematical Monthly, 104 (3):261–269, March 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_mar97_
                                        toc.html.

Mason:1996:P

[678] Eric H. Mason, Stephen M. Gagola, Jr., John Todd, Theodore J. Rivlin, and David Callan. Polynomialrecurrencology: 10300. American Mathematical Monthly, 103(3):272–273, March 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL https://www.jstor.
                                        org/stable/2975388.

Berrick:1997:RIM

[679] A. J. Berrick and M. E. Keating. Rectangular invertible matrices. American Mathematical Monthly, 104(4): 297–302, April 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        maa.org/pubs/monthly_apr97_toc.
                                        html.

Boerlijst:1997:EPA

[680] Maarten C. Boerlijst, Martin A. Nowak, and Karl Sigmund. Equal pay for all prisoners. American Mathematical Monthly, 104(4):303–305, April 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr97_toc.html.

Sauerberg:1997:LSC

[681] Jim Sauerberg and Linghsueh Shu. The long and the short on counting sequences. American Mathematical Monthly, 104(4):306–317, April 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr97_toc.html.

Velleman:1997:CC

[682] Daniel J. Velleman. Characterizing continuity. American Mathematical Monthly, 104(4):318–322, April 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr97_toc.html.

Bruckstein:1997:PPG

[683] A. M. Bruckstein, C. L. Mallows, and I. A. Wagner. Probabilistic pursuits on the grid. American Mathematical Monthly, 104(4):323–343, April 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr97_toc.html.

Stanley:1997:HPS

[684] Richard P. Stanley. Hipparchus, Plutarch, Schröder, and Hough. American Mathematical Monthly, 104(4): 344–350, April 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_apr97_toc.
html.

Almkvist:1997:MCD

[685] Gert Almkvist. Many correct digits of π, revisited. American Mathematical Monthly, 104(4):351–353, April 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_apr97_toc.html.

Jones:1997:NCC

[686] Martin L. Jones. A note on a cake cutting algorithm of Banach and Knaster. American Mathematical Monthly, 104(4):353–355, April 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_apr97_toc.html.

Sedaghat:1997:IUG

[687] Hassan Sedaghat. The impossibility of unstable, globally attracting fixed points for continuous mappings of the line. American Mathematical Monthly, 104(4):356–358, April 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_apr97_toc.html.

Grant:1997:RVM

[688] Hardy Grant. Reviews: Vita Mathematica: Historical Research and Integration with Teaching, edited by Ronald Calinger. American Mathematical Monthly, 104(5):??, May 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

                                        http://www.maa.org/pubs/monthly_
                                        may97_toc.html.

Toom:1997:RMC

[689] Andre Toom. Reviews: Mathematical Circles (Russian Experience), by Dmitri Fomin, Sergey Genkin, and Ilia Itenberg. American Mathematical Monthly, 104(5):??, May 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.maa.org/pubs/monthly_
may97_toc.html.

Grabiner:1997:WNC

[690] Judith V. Grabiner. Was Newton’s calculus a dead end? the continental influence of Maclaurin’s treatise of fluxions. American Mathematical Monthly, 104(5):393–410, May 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_may97_toc.html.

Touhey:1997:YAD

[691] Pat Touhey. Yet another definition of chaos. American Mathematical Monthly, 104(5):411–414, May 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_may97_toc.html.

McCullough:1997:HCO

[692] Scott A. McCullough and Leiba Rodman. Hereditary classes of operators and matrices. American Mathematical Monthly, 104(5):415–430, May 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_may97_toc.html.

Kimberling:1997:MCT

[693] Clark Kimberling. Major centers of triangles. American Mathematical Monthly, 104(5):431–438, May 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_may97_toc.html.

Laczkovich:1997:LPI

[694] M. Laczkovich. On Lambert’s proof of the irrationality of π. American Mathematical Monthly, 104(5):439–443, May 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_may97_toc.html.

Kohnen:1997:SCM

[695] Winfried Kohnen. A simple congruence modulo p. American Mathematical Monthly, 104(5):444–445, May 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_may97_toc.html.

Polezzi:1997:GMF

[696] Marcelo Polezzi. A geometrical method for finding an explicit formula for the greatest common divisor. American Mathematical Monthly, 104(5): 445–446, May 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        maa.org/pubs/monthly_may97_toc.
                                        html.

Laugwitz:1997:ELHa

[697] Detlef Laugwitz. The evolution of …: On the historical development of infinitesimal mathematics. American Mathematical Monthly, 104(5):

447–455, May 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_may97_toc.
html.

Axler:1997:RLS

[698] Sheldon Axler. Reviews: The Life of Stefan Banach, by Roman Kaluza. American Mathematical Monthly, 104 (6):??, June/July 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_june97_toc.
html.

Colligan:1997:RCM

[699] J. Kevin Colligan. Reviews: 101 Careers in Mathematics, edited by Andrew Sterrett. American Mathematical Monthly, 104(6):??, June/July 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_june97_toc.html.

Denes:1997:UPW

[700] J. Dénes. Unsolved problems: When is there a Latin power set? American Mathematical Monthly, 104(6): ??, June/July 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_june97_toc.
html.

King:1997:LB

[701] Melanie A. King. Lecturing at the ‘bored’. American Mathematical Monthly, 104(6):??, June/July 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_june97_toc.html.

Bhatia:1997:AIE

[702] Rajendra Bhatia and Peter Šemrl. Approximate isometries on Euclidean spaces. American Mathematical Monthly, 104(6):497–504, June/July 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_june97_toc.html.

Nemes:1997:HDM

[703] István Nemes, Marko Petkovšek, Herbert S. Wilf, and Doron Zeilberger. How to do Monthly problems with your computer. American Mathematical Monthly, 104(6):505–519, June/ July 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_june97_toc.html.

Baggett:1997:SSF

[704] Lawrence W. Baggett, Herbert A. Medina, and Kathy D. Merrill. Simultaneously symmetric functions. American Mathematical Monthly, 104(6): 520–528, June/July 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.maa.org/pubs/monthly_june97_
                                        toc.html.

Mollin:1997:PPQ

[705] R. A. Mollin. Prime-producing quadratics. American Mathematical Monthly, 104(6):529–544, June/July 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_june97_toc.html.

Kulpa:1997:PMT

[706] Wladyslaw Kulpa. The Poincaré-Miranda theorem. American Mathematical Monthly, 104(6):545–550, June/July 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_june97_toc.
html.

Bayat:1997:GWT

[707] M. Bayat. A generalization of Wolstenholme’s theorem. American Mathematical Monthly, 104(6):557–560, June/ July 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_june97_toc.html.

Zhang:1997:NMV

[708] Bao lin Zhang. A note on the mean value theorem for integrals. American Mathematical Monthly, 104(6): 561–562, June/July 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.maa.org/pubs/monthly_june97_
toc.html.

Danby:1997:RSJ

[709] J. N. Anthony Danby. Reviews: The Sheer Joy of Celestial Mechanics, by Nathaniel Grossman. American Mathematical Monthly, 104(7):??, August/ September 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_aug97_toc.html.

Galovich:1997:RSN

[710] Jennifer R. Galovich. Reviews: Strength in Numbers: Discovering the Joy and Power of Mathematics in Everyday Life, by Sherman K. Stein. American Mathematical Monthly, 104 (7):??, August/September 1997. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.maa.org/pubs/monthly_
aug97_toc.html.

Merryfield:1997:NWP

[711] Kent G. Merryfield, Ngo Viet, and Saleem Wats. Notes: The wallet paradox. American Mathematical Monthly, 104(7):??, August/September 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_aug97_toc.html.

Coutinho:1997:MAS

[712] S. C. Coutinho. The many avatars of a simple algebra. American Mathematical Monthly, 104(7):593–604, August/ September 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_aug97_toc.html.

Trif:1997:MIS

[713] Tiberiu Trif. Multiple integrals of symmetric functions. American Mathematical Monthly, 104(7):605–608, August/ September 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_aug97_toc.html.

Chollet:1997:SIP

[714] John Chollet. Some inequalities for principal submatrices. American Mathematical Monthly, 104(7):609–617, August/September 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_aug97_toc.
html.

Gutkin:1997:TAC

[715] Eugene Gutkin. Two applications of calculus to triangular billiards. American Mathematical Monthly, 104(7): 618–622, August/September 1997. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.maa.org/pubs/monthly_
aug97_toc.html.

Weintraub:1997:ET

[716] Steven H. Weintraub. Early transcendentals. American Mathematical Monthly, 104(7):623–631, August/ September 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_aug97_toc.html.

Devlin:1997:LSC

[717] Keith Devlin. The logical structure of computer-aided mathematical reasoning. American Mathematical Monthly, 104(7):632–646, August/ September 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_aug97_toc.html.

Gzyl:1997:NWA

[718] Henryk Gzyl and José Luis Palacios. Notes: The Weierstrass approximation theorem and large deviations. American Mathematical Monthly, 104(7): 651–653, August/September 1997. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.maa.org/pubs/monthly_
aug97_toc.html.

Laugwitz:1997:ELHb

[719] Detlef Laugwitz. The evolution of …: On the historical development of infinitesimal mathematics. II. The conceptual thinking of Cauchy. Translated from the German by Abe Shenitzer with the editorial assistance of Hardy Grant. American Mathematical Monthly, 104(7):654–663, August/ September 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_aug97_toc.html.

Translated from the German by Abe Shenitzer with the editorial assistance of Hardy Grant.

Peyerimhoff:1997:AIC

[720] Norbert Peyerimhoff. Areas and intersections in convex domains. American Mathematical Monthly, 104(8): 697–704, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_oct97_
toc.html.

Zagier:1997:NSP

[721] D. Zagier. Newman’s short proof of the Prime Number Theorem. American Mathematical Monthly, 104(8): 705–708, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_oct97_
toc.html.

Callan:1997:EAK

[722] David Callan. An exploratory approach to Kaplansky’s lemma leads to a generalized resultant. American Mathematical Monthly, 104(8):709–712, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_oct97_toc.html.

Richmond:1997:MSW

[723] Bettina Richmond and Thomas Richmond. Metric spaces in which all triangles are degenerate. American Mathematical Monthly, 104(8):713–719, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_oct97_toc.html.

Fathi:1997:PUC

[724] Albert Fathi. Partitions of unity for countable covers. American Mathematical Monthly, 104(8):720–723, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_oct97_toc.html.

Knisley:1997:CMP

[725] Jeff Knisley. Calculus: a modern perspective. American Mathematical Monthly, 104(8):724–??, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_oct97_toc.html.

Zorn:1997:PC

[726] Paul Zorn and Arnold Ostebee. Pro choice. American Mathematical Monthly, 104(8):728–??, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.maa.org/pubs/monthly_
oct97_toc.html.

Kaput:1997:RCL

[727] James J. Kaput. Rethinking calculus: Learning and thinking. American Mathematical Monthly, 104(8):731–??, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_oct97_toc.html.

Askey:1997:WDW

[728] Richard Askey. What do we do about calculus? first, do no harm. American Mathematical Monthly, 104(8):738–??, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_oct97_toc.html.

Alexanderson:1997:FSW

[729] Gerald L. Alexanderson, Leonard F. Klosinski, and Loren C. Larson. The Fifty-Seventh William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 104(8):744–??, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_oct97_toc.html.

Kupka:1997:NQT

[730] Joseph Kupka. Notes: a quadratic trio. American Mathematical Monthly, 104 (8):755–757, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_oct97_
                                        toc.html.

Tyszka:1997:NDF

[731] Apoloniusz Tyszka. Notes: a discrete form of the Beckman-Quarles theorem. American Mathematical Monthly, 104 (8):757–761, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_oct97_
toc.html.

Christensen:1997:UPD

[732] J. Daniel Christensen and Mark Tilford. Unsolved problems: David Gale’s subset take-away game. American Mathematical Monthly, 104(8):762–??, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_oct97_toc.html.

Mickens:1997:RID

[733] Ronald E. Mickens. Reviews: An Introduction to Difference Equations, by Saber Elyadi. American Mathematical Monthly, 104(8):777–??, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_oct97_toc.html.

Roberts:1997:RCL

[734] Wayne Roberts. Reviews: Calculus Lite, by Frank Morgan. American Mathematical Monthly, 104(8):780–??, October 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_oct97_toc.html.

Cobb:1997:MST

[735] George W. Cobb and David S. Moore. Mathematics, statistics, and teaching. American Mathematical Monthly, 104 (9):801–??, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_nov97_
toc.html.

Abate:1997:WLO

[736] Marco Abate. When is a linear operator diagonalizable? American Mathematical Monthly, 104(9): 824–830, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_nov97_
                                        toc.html.

Bach:1997:EAT

[737] Eric Bach. Energy arguments in the theory of algorithms. American Mathematical Monthly, 104(9): 831–837, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_nov97_
                                        toc.html.

Forster:1997:QY

[738] Thomas Forster. Quine’s NF—60 years on. American Mathematical Monthly, 104(9):838–845, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_nov97_toc.html.

Misiurewicz:1997:RST

[739] Michał Misiurewicz. Remarks on Sharkovsky’s theorem. American Mathematical Monthly, 104(9):846– 847, November 1997. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        maa.org/pubs/monthly_nov97_toc.
                                        html.

Granville:1997:CZB

[740] Andrew Granville. Correction to: “Zaphod Beeblebrox’s brain and the fifty-ninth row of Pascal’s triangle”

[Amer. Math. Monthly 99 (1992), no. 4, 318–331; MR 93a:05008]. American Mathematical Monthly, 104(9): 848–851, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_nov97_

toc.html. See [231].

Adamchik:1997:SF

[741] Victor Adamchik and Stan Wagon. A simple formula for π. American Mathematical Monthly, 104(9): 852–855, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_nov97_

toc.html. The authors employ Mathematica to extend earlier work of Bailey, Borwein, and Plouffe, [1122], done in 1995, but only just published, that discovered an amazing formula for π as is a power series in 16^{−k}, enabling any base-16 digit of π to be computed without knowledge of any prior digits. In this paper, Mathematica is used to find several simpler formulas having powers of 4^{−k}. They also note that it has been proven that their methods cannot be used to exhibit similar formulas in powers of 10^{−k}.

Su:1997:BUI

[742] Francis Edward Su. Borsuk–Ulam implies Brouwer: a direct construction. American Mathematical Monthly, 104 (9):855–859, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_nov97_
toc.html.

Pogoda:1997:EOD

[743] Zdzisław Pogoda and Leszek M.

Sokołowski. The evolution of…: Does mathematics distinguish certain dimensions of spaces? Translated by Abe Shenitzer. American Mathematical Monthly, 104(9): 860–869, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_nov97_

toc.html. Translated by Abe Shenitzer.

Dunbar:1997:RKS

[744] William D. Dunbar. Reviews: Knots and Surfaces: A Guide to Discovering Mathematics, by David W. Farmer and Theodore Stanford & Knots and Surfaces, by N. D. Gilbert and T. Porter. American Mathematical Monthly, 104(9):882–??, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_nov97_toc.html.

Bremner:1997:RBN

[745] Andrew Bremner. Reviews: The Book of Numbers, by John Horton Conway and Richard K. Guy. American Mathematical Monthly, 104(9):884–??, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_nov97_toc.html.

Berndt:1997:RAR

[746] Bruce C. Berndt, Heng Huat Chan, and Liang-Cheng Zhang. Ramanujan’s association with radicals in India. American Mathematical Monthly, 104(10): 905–911, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_dec97_
toc.html.

Ono:1997:RTB

[747] Ken Ono. Ramanujan, taxicabs, birthdates, ZIP codes, and twists. American Mathematical Monthly, 104(10): 912–917, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_dec97_
toc.html.

Andrews:1997:SSR

[748] George E. Andrews. Simplicity and surprise in Ramanujan’s “lost” notebook. American Mathematical Monthly, 104 (10):918–925, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_dec97_
toc.html.

Shapiro:1997:CNL

[749] Lou Shapiro, D. G. Rogers, and Wen-Jin Woan. The Catalan numbers, the Lebesgue integral, and 4^{n} − 2. American Mathematical Monthly, 104(10): 926–931, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_dec97_
toc.html.

Richter:1997:GMM

[750] R. Bruce Richter and William P. Wardlaw. Good matrices: Matrices that preserve ideals. American Mathematical Monthly, 104(10):932–938, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_dec97_toc.html.

Braza:1997:CMV

[751] Peter A. Braza and Jingcheng Tong. A converse of the Mean Value Theorem. American Mathematical Monthly, 104 (10):939–942, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_dec97_
                                        toc.html.

Granville:1997:PSL

[752] Andrew Granville, Takashi Agoh, and Paul Erdős. Primes at a (somewhat lengthy) glance. American Mathematical Monthly, 104(10):943–945, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_dec97_toc.html.

Wu:1997:MER

[753] H. Wu. The mathematics education reform: Why you should be concerned and what you can do. American Mathematical Monthly, 104(10):946–??, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_dec97_toc.html.

Kilpatrick:1997:CR

[754] Jeremy Kilpatrick. Confronting reform. American Mathematical Monthly, 104 (10):955–??, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_dec97_
                                        toc.html.

Drost:1997:SPR

[755] John L. Drost. A shorter proof of the Ramanujan congruence modulo 5. American Mathematical Monthly, 104 (10):963–964, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_dec97_
toc.html.

Ahlgren:1997:AR

[756] Scott Ahlgren, Lars English, and Ron Winters. An amusing representation of x/(sin x). American Mathematical Monthly, 104(10):964–966, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_dec97_toc.html.

Nowakowski:1997:MUP

[757] Richard J. Nowakowski and Richard K. Guy. Monthly unsolved problems, 1969–1997. American Mathematical Monthly, 104(10):967–??, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_dec97_toc.html.

Lane:1997:CMF

[758] Saunders Mac Lane. Conceptual Mathematics: A First Introduction to Categories, by F. William Lawvere and Steven Schanuel. American Mathematical Monthly, 104(10):985–??, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_dec97_toc.html.

Brown:1997:EAH

[759] Ezra Brown. Early Astronomy, by Hugh Thurston. American Mathematical Monthly, 104(10):988–??, December 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_dec97_toc.html.

Kazdan:1998:SEE

[760] Jerry L. Kazdan. Solving equations, an elegant legacy. American Mathematical Monthly, 105(1):1–21, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_jan98_toc.html.

Robinson:1998:SSV

[761] Donald W. Robinson. Separation of subspaces by volume. American Mathematical Monthly, 105(1):22–27, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_jan98_toc.html.

Chung:1998:PD

[762] Kai Lai Chung. Probability and Doob. American Mathematical Monthly, 105 (1):28–35, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_jan98_
                                        toc.html.

Andrews:1998:GSC

[763] George E. Andrews. The geometric series in calculus. American Mathematical Monthly, 105(1):36–40, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_jan98_toc.html.

Chow:1998:SEU

[764] Timothy Y. Chow. The surprise examination or unexpected hanging paradox. American Mathematical Monthly, 105(1):41–51, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_jan98_
toc.html.

Hirschhorn:1998:NTT

[765] Michael David Hirschhorn. Notes: Two or three identities of Ramanujan. American Mathematical Monthly, 105(1):52–55, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_jan98_
toc.html.

Hoffman:1998:BSA

[766] Michael E. Hoffman. The bull and the silo: An application of curvature. American Mathematical Monthly, 105(1):55–58, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_jan98_toc.html.

Luzin:1998:ELFa

[767] N. Luzin. The evolution of … function: Part I. American Mathematical Monthly, 105(1):59–67, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.maa.org/pubs/monthly_

jan98_toc.html. Translated from the Polish by Abe Shenitzer.

Troutman:1998:RAH

[768] John Troutman. Reviews: Analysis by its History, by E. Hairer and G. Wanner. American Mathematical Monthly, 105(1):79–??, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_jan98_toc.html.

Ralston:1998:RPP

[769] Anthony Ralston. Reviews: Principles and Practices of Mathematics, by COMAP. American Mathematical Monthly, 105(1):81–??, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_jan98_toc.html.

Toom:1998:RMM

[770] Andre Toom. Reviews: A Mathematical Mosaic: Patterns & Problem Solving, by Ravi Vakil. American Mathematical Monthly, 105(1):86–??, January 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_jan98_toc.html.

Sons:1998:YGG

[771] Linda R. Sons. Yueh-Gin Gung and Dr. Charles Y. Hu Award for Distinguished Service to Alice Turner Schafer. American Mathematical Monthly, 105(2): 105–??, February 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        maa.org/pubs/monthly_feb98_toc.
                                        html.

Gordon:1998:UTP

[772] Russell A. Gordon. The use of tagged partitions in elementary real analysis. American Mathematical Monthly, 105 (2):107–117, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_feb98_
                                        toc.html.

Bassein:1998:DFO

[773] Susan Bassein. The dynamics of a family of one-dimensional maps. American Mathematical Monthly, 105(2): 118–130, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb98_
toc.html.

Snevily:1998:BPS

[774] Hunter S. Snevily and Douglas B. West. The bricklayer problem and the strong cycle lemma. American Mathematical Monthly, 105(2):131–143, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_feb98_toc.html.

Johnson:1998:CSF

[775] Craig M. Johnson. A computer search for free actions on surfaces. American Mathematical Monthly, 105(2): 144–153, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb98_
toc.html.

McConnell:1998:DAB

[776] John C. McConnell. Division algebras — beyond the quaternions. American Mathematical Monthly, 105(2): 154–162, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb98_
toc.html.

Gerard:1998:GDN

[777] Brian Gerard and Lawrence Roberts. Graphical discovery of a new identity for Jacobi polynomials. American Mathematical Monthly, 105(2): 163–166, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb98_
toc.html.

Revoy:1998:GLN

[778] Philippe Revoy. The generalized level of a non prime finite field is two. American Mathematical Monthly, 105 (2):167–168, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb98_
toc.html.

Kuhn:1998:CHD

[779] Wolfgang Kühn and Zuzana Kühn. Cutting high-dimensional cakes. American Mathematical Monthly, 105(2): 168–169, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb98_
toc.html.

Hutchinson:1998:KR

[780] Joan Hutchinson and Stan Wagon. Kempe revisited. American Mathematical Monthly, 105(2):170–174, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_feb98_toc.html.

Pedersen:1998:RIP

[781] Jean Pedersen and Gerald L. Alexanderson. Reviews: “Invertible” Polyhedron Models, distributed by Snyder Engineering. American Mathematical Monthly, 105(2):186–??, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_feb98_toc.html.

Goldman:1998:RTG

[782] William Goldman. Reviews: Topology and Geometry, by Glen E. Bredon. American Mathematical Monthly, 105 (2):192–??, February 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_feb98_
toc.html.

Bollobas:1998:PCP

[783] Béla Bollobás. To prove and conjecture: Paul Erdős and his mathematics. American Mathematical Monthly, 105(3):209–237, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_mar98_
toc.html.

Libeskind-Hadas:1998:SP

[784] Ran Libeskind-Hadas. Sorting in parallel. American Mathematical Monthly, 105(3):238–245, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_mar98_
toc.html.

Schechter:1998:IPA

[785] Murray Schechter. Integration over a polyhedron: An application of the Fourier-Motzkin elimination method. American Mathematical Monthly, 105 (3):246–251, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_mar98_
toc.html.

Ross:1998:DPP

[786] Kenneth A. Ross. Doing and proving: The place of algorithms and proofs in school mathematics. American Mathematical Monthly, 105(3): 252–255, March 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        maa.org/pubs/monthly_mar98_toc.
                                        html.

Beckwith:1998:LPP

[787] David Beckwith. Legendre polynomials and polygon dissections? American Mathematical Monthly, 105(3): 256–257, March 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        maa.org/pubs/monthly_mar98_toc.
                                        html.

Furuta:1998:GMC

[788] Takayuki Furuta and Masahiro Yanagida. Generalized means and convexity of inversion for positive operators. American Mathematical Monthly, 105(3): 258–259, March 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        maa.org/pubs/monthly_mar98_toc.
                                        html.

Saleh:1998:NRH

[789] Mohammad Saleh and Hasan Yousef. The number of ring homomorphisms from Z_{m_{1}} ×⋯×Z_{m_{r}}intoZ_{k_{1}} ×⋯×Z_{k_{s}}. American Mathematical Monthly, 105 (3):259–260, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_mar98_
                                        toc.html.

Mirzakhani:1998:SPT

[790] M. Mirzakhani. A simple proof of a theorem of Schur. American Mathematical Monthly, 105(3):260–262, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_mar98_toc.html.

Luzin:1998:ELFb

[791] N. Luzin. The evolution of … function: Part II. American Mathematical Monthly, 105(3):263–270, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.maa.org/pubs/monthly_

mar98_toc.html. Translated by Abe Schenitzer.

Anonymous:1998:PSa

[792] Anonymous. Problems and solutions. American Mathematical Monthly, 105 (3):271–283, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cooke:1998:EG

[793] Roger Cooke. Évariste Galois 1811– 1832, by Laura Toti Rigatelli. translated from the Italian by John Denton. birkhäuser, 1996, 168pp., $29.50. American Mathematical Monthly, 105 (3):284–288, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_mar98_
toc.html.

Wimp:1998:FFT

[794] Jet Wimp. Reviews: The Fast Fourier Transform Workshop, by Richard and Sandra McPeak. American Mathematical Monthly, 105(3):289–292, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_mar98_toc.html.

Ullman:1998:RWW

[795] Daniel H. Ullman. Reviews: Which Way Did the Bicycle Go?, by Joseph D. E. Konhauser, Dan Velleman, and Stan Wagon and Problems and Solutions from the Mathematical Visitor, edited by Stanley Rabinowitz. American Mathematical Monthly, 105 (3):292–296, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_mar98_
toc.html.

Ostebee:1998:TRa

[796] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 105 (3):297–301, March 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Vardi:1998:ACP

[797] Ilan Vardi. Archimedes’ cattle problem. American Mathematical Monthly, 105(4):305–319, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_apr98_
toc.html.

Groetsch:1998:LGD

[798] C. W. Groetsch. Lanczos’ generalized derivative. American Mathematical Monthly, 105(4):320–326, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr98_toc.html.

Gethner:1998:STG

[799] Ellen Gethner, Stan Wagon, and Brian Wick. A stroll through the Gaussian primes. American Mathematical Monthly, 105(4):327–337, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr98_toc.html.

Erlichson:1998:GWS

[800] Herman Erlichson. Galileo’s work on swiftest descent from a circle and how he almost proved the circle itself was the minimum time path. American Mathematical Monthly, 105(4): 338–347, April 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_apr98_toc.
html.

Sheard:1998:IHW

[801] Michael Sheard. Induction the hard way. American Mathematical Monthly, 105(4):348–353, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_apr98_
toc.html.

Howard:1998:GPD

[802] C. Douglas Howard. Good paths don’t double back. American Mathematical Monthly, 105(4):354–357, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr98_toc.html.

Wang:1998:NCS

[803] Xianfu Wang. A note on the Clarke subdifferential. American Mathematical Monthly, 105(4):357–359, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_apr98_toc.html.

Reid:1998:NCC

[804] Michael Reid. The number of conjugacy classes. American Mathematical Monthly, 105(4):359–361, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_apr98_toc.html.

Grunbaum:1998:WCS

[805] Branko Grünbaum. Which coronas are simple? American Mathematical Monthly, 105(4):362–365, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_apr98_toc.html.

Anonymous:1998:PSb

[806] Anonymous. Problems and solutions. American Mathematical Monthly, 105 (4):366–376, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_apr98_
                                        toc.html.

Sanchez:1998:ODE

[807] David A. Sánchez. Ordinary differential equations texts. American Mathematical Monthly, 105(4):377–383, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_apr98_toc.html.

Thomas:1998:SWN

[808] Owen Thomas. The Schools We Need, by E. D. Hirsch, Jr. American Mathematical Monthly, 105(4):384–386, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr98_toc.html.

Katz:1998:FGN

[809] Victor J. Katz. Force and Geometry in Newton’s Principia, by François De Gandt, translated by Curtis Wilson. American Mathematical Monthly, 105(4):386–392, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_apr98_toc.html.

Ostebee:1998:TRb

[810] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 105 (4):393–397, April 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_apr98_
toc.html.

Case:1998:MAF

[811] James Case. The modeling and analysis of financial time series. American Mathematical Monthly, 105(5): 401–411, May 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_may98_toc.
html.

Lowenthal:1998:MPD

[812] Franklin Lowenthal, Arnold Langsen, and Clark T. Benson. Merton’s partial differential equation and fixed point theory. American Mathematical Monthly, 105(5):412–420, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_may98_toc.html.

Knorr:1998:RDD

[813] Wilbur R. Knorr. “rational diameters” and the discovery of incommensurability. American Mathematical Monthly, 105(5):421–429, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_may98_toc.html.

Balinski:1998:GM

[814] Michel Balinski and Guillaume Ratier. Graphs and marriages. American Mathematical Monthly, 105(5):430– 445, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

org/pubs/monthly_may98_toc.html.

Habsieger:1998:SNP

[815] Laurent Habsieger, Maxim Kazarian, and Sergei Lando. On the second number of Plutarch. American Mathematical Monthly, 105(5):446, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_may98_toc.html.

Anderson:1998:DRS

[816] Bruce Anderson, Jeffrey Jackson, and Meera Sitharam. Descartes’ rule of signs revisited. American Mathematical Monthly, 105(5):447–451, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/
pubs/monthly_may98_toc.html.

Faudree:1998:CE

[817] Ralph Faudree. A conjecture of Erdős. American Mathematical Monthly, 105 (5):451–453, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_may98_
toc.html.

Driver:1998:TLI

[818] R. D. Driver. Toricelli’s law — an ideal example of an elementary ODE. American Mathematical Monthly, 105 (5):453–455, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_may98_
toc.html.

Pogoda:1998:DMD

[819] Zdzisław Pogoda and Leszek M. Sokołowski. Does mathematics distinguish certain dimensions of spaces? part II. American Mathematical Monthly, 105(5):456–463, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.maa.org/pubs/monthly_

may98_toc.html. Translated from the Polish by Abe Shenitzer.

Anonymous:1998:PSc

[820] Anonymous. Problems and solutions. American Mathematical Monthly, 105 (5):464–476, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/www.maa.org/pubs/monthly_may98_
toc.html.

Schwalbe:1998:CEF

[821] Daniel Schwalbe and Hal R. Varian. Computational economics and finance: Modeling and analysis with Mathematica. American Mathematical Monthly, 105(5):477–481, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_may98_toc.html.

Hilton:1998:PC

[822] Peter Hilton and T. W. Korner. The pleasures of counting. American Mathematical Monthly, 105(5):481–485, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.maa.org/
                                        pubs/monthly_may98_toc.html.

Gillman:1998:WM

[823] Leonard Gillman, Richard Courant, Herbert Robbins, and Ian Stewart. What is mathematics? American Mathematical Monthly, 105(5):485– 488, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.maa.

                                        org/pubs/monthly_may98_toc.html.

Ostebee:1998:TRc

[824] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 105 (5):489–493, May 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /www.maa.org/pubs/monthly_may98_
                                        toc.html.

Horn:1998:EE

[825] Roger A. Horn. Editor’s endnotes. American Mathematical Monthly, 105 (5):494, May 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

maa.org/pubs/monthly_may98_toc.
html.

Harel:1998:TDA

[826] Guershon Harel. Two dual assertions: The first on learning and the second on teaching (or vice versa). American Mathematical Monthly, 105 (6):497–??, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gluchoff:1998:UPN

[827] Alan Gluchoff and Frederick Hartmann. Univalent polynomials and nonnegative trigonometric sums. American Mathematical Monthly, 105(6): 508–522, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Neumann:1998:RRS

[828] Peter M. Neumann. Reflections on reflection in a spherical mirror. American Mathematical Monthly, 105(6): 523–528, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Blecksmith:1998:SRI

[829] Richard Blecksmith, Michael McCallum, and J. L. Selfridge. 3-smooth representations of integers. American Mathematical Monthly, 105(6): 529–543, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Groetsch:1998:AF

[830] C. W. Groetsch. Aristotle’s fall. American Mathematical Monthly, 105(6): 544–547, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Banaschewski:1998:PEC

[831] B. Banaschewski. On proving the existence of complete ordered fields. American Mathematical Monthly, 105 (6):548–551, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kicey:1998:S

[832] Charles Kicey and Sudhir Goel. A series for ln k. American Mathematical Monthly, 105(6):552–554, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Herzog:1998:SCC

[833] Gerd Herzog. C^{1}-solutions of x = f(x′) are convex or concave. American Mathematical Monthly, 105(6): 554–555, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lampert:1998:PKC

[834] Douglas E. Lampert and Peter J. Slater. Parallel knockouts in the complete graph. American Mathematical Monthly, 105(6):556–558, June/July 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Farris:1998:VCA

[835] Frank A. Farris. Visual Complex Analysis, by Tristan Needham. American Mathematical Monthly, 105(6):

570–??, June/July 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Messer:1998:LAP

[836] Robert Messer. Linear Algebra Problem Book, by Paul R. Halmos. American Mathematical Monthly, 105(6): 577–??, June/July 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Singmaster:1998:AMP

[837] David Singmaster. After Math. Puzzles and Brainteasers, by Ed Barbeau; The Chicken from Minsk, by Yuri B. Chernyak and Robert M. Rose; New Mathematical Diversions, revised edition, by Martin Gardner. American Mathematical Monthly, 105(6): 579–??, June/July 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bosch:1998:SHA

[838] Robert A. Bosch and Jason A. Smith. Separating hyperplanes and the authorship of the disputed Federalist papers. American Mathematical Monthly, 105(7):601–608, August/ September 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Frantz:1998:TFW

[839] Marc Frantz. Two functions whose powers make fractals. American Mathematical Monthly, 105(7):609–617, August/September 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Eggar:1998:PCP

[840] M. H. Eggar. Pinhole cameras, perspective, and projective geometry. American Mathematical Monthly, 105(7): 618–630, August/September 1998. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dosen:1998:FR

[841] Kosta Došen. Functions redefined. American Mathematical Monthly, 105 (7):631–635, August/September 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Komornik:1998:UDN

[842] Vilmos Komornik and Paola Loreti. Unique developments in non-integer bases. American Mathematical Monthly, 105(7):636–639, August/ September 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hamming:1998:MDP

[843] R. W. Hamming. Mathematics on a distant planet. American Mathematical Monthly, 105(7):640–650, August/September 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lam:1998:TBM

[844] T. Y. Lam. A theorem of Burnside on matrix rings. American Mathematical Monthly, 105(7):651–653, August/September 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dierolf:1998:PCV

[845] Peter Dierolf and Volker Schmidt. A proof of the change of variable formula for d-dimensional integrals. American Mathematical Monthly, 105(7): 654–656, August/September 1998. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Yu:1998:DE

[846] Hong Bing Yu. On the Diophantine equation (x + 1)^{y} − x^{z} = 1. American Mathematical Monthly, 105 (7):656–657, August/September 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gessel:1998:WR

[847] Ira M. Gessel. Wolstenholme revisited. American Mathematical Monthly, 105 (7):657–658, August/September 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Medvedev:1998:NAH

[848] F. A. Medvedev. Nonstandard analysis and the history of classical analysis. American Mathematical Monthly, 105 (7):659–664, August/September 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). Translated by Abe Shenitzer.

Krantz:1998:ICC

[849] Steven G. Krantz. Introduction to Calculus and Classical Analysis, by Omar Hijab; Introduction to Mathematical Structures and Proofs, by Larry J. Gerstein. American Mathematical Monthly, 105(7):677–??, August/September 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anthony:1998:MRR

[850] Joby Milo Anthony. Mathematical Reflections: In a Room with Many Mirrors, by Peter Hilton, Derek Holton, and Jean Pedersen. American Mathematical Monthly, 105(7):682–??, August/September 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Devaney:1998:IDE

[851] Robert L. Devaney. Interactive Differential Equations, by Beverly West, Steven Strogatz, Jean Marie McDill, and John Cantwell; software designer Hubert Hohn. American Mathematical Monthly, 105(7):687–??, August/September 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tokieda:1998:MIG

[852] Tadashi F. Tokieda. Mechanical ideas in geometry. American Mathematical Monthly, 105(8):697–703, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cheng:1998:WD

[853] R. Cheng, A. Dasgupta, B. R. Ebanks, L. F. Kinch, L. M. Larson, and R. B. McFadden. When does f − 1 = 1/f ? American Mathematical Monthly, 105 (8):704–717, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kraft:1998:GCS

[854] Roger L. Kraft. A golden Cantor set. American Mathematical Monthly, 105 (8):718–725, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Desbrow:1998:ECI

[855] Darrell Desbrow. On evaluating ∫_{−∞}^{+∞} e^{ax(x−2b)}dx by contour integration round a parallelogram. American Mathematical Monthly, 105(8): 726–731, October 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gross:1998:GCG

[856] C. Groß and T.-K. Strempel. On generalizations of conics and on a generalization of the Fermat–Torricelli problem. American Mathematical Monthly, 105 (8):732–743, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Klosinski:1998:FEW

[857] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Fifty-Eighth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 105 (8):744–??, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Korner:1998:BHS

[858] T. W. Körner. Butterfly hunting and the symmetry of mixed partial derivatives. American Mathematical Monthly, 105(8):756–758, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gagola:1998:NNE

[859] Stephen M. Gagola, Jr. A note on the non-existence of regular bases for finite groups. American Mathematical Monthly, 105(8):759–761, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nathan:1998:INR

[860] Joseph Amal Nathan. The irrationality of e^{x} for nonzero rational x. American Mathematical Monthly, 105(8): 762–763, October 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Warrack:1998:NVR

[861] Giles Warrack. A note on variable replacement rates in urn models. American Mathematical Monthly, 105(8): 764–765, October 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Guy:1998:FF

[862] Richard K. Guy and John L. Selfridge. Factoring factorial n. American Mathematical Monthly, 105(8): 766–??, October 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Folland:1998:HWM

[863] Gerald B. Folland. Handbook of Writing for the Mathematical Sciences, by Nicholas J. Higham; A Primer of Mathematical Writing, by Steven G. Krantz. American Mathematical Monthly, 105 (8):779–??, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hogg:1998:LPS

[864] Robert V. Hogg. Leading Personalities in Statistical Sciences, edited by Norman L. Johnson and Samuel Kotz. American Mathematical Monthly, 105 (8):782–??, October 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Manaster:1998:SCE

[865] Alfred B. Manaster. Some characteristics of eighth grade mathematics classes in the TIMSS videotape study. American Mathematical Monthly, 105 (9):793–??, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Yau:1998:NHM

[866] Lily Yau and Adi Ben-Israel. The Newton and Halley methods for complex roots. American Mathematical Monthly, 105(9):806–818, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kedlaya:1998:RHC

[867] Kiran S. Kedlaya and Lenhard L. Ng. The rook on the half-chessboard, or how not to diagonalize a matrix. American Mathematical Monthly, 105(9): 819–824, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cassidy:1998:LRB

[868] Jack Cassidy. The last round of betting in poker. American Mathematical Monthly, 105(9):825–831, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Benyamini:1998:AUS

[869] Yoav Benyamini. Applications of the universal surjectivity of the Cantor set. American Mathematical Monthly, 105 (9):832–839, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Thoma:1998:GRP

[870] Apostolos Thoma. The group Z_{2} × Z_{n} and regular polygonal paths. American Mathematical Monthly, 105(9): 840–841, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chisala:1998:QCH

[871] Busiso P. Chisala. A quick Cayley– Hamilton. American Mathematical Monthly, 105(9):842–844, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ash:1998:PTG

[872] J. Marshall Ash. The probability of a tie in an n-game match. American Mathematical Monthly, 105(9): 844–846, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:1998:CMV

[873] J. M. Borwein and Xianfu Wang. The converse of the mean value theorem may fail generically. American Mathematical Monthly, 105(9): 847–848, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ewell:1998:TTN

[874] John A. Ewell. A trio of triangular number theorems. American Mathematical Monthly, 105(9): 848–849, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stillwell:1998:EO

[875] John Stillwell. Exceptional objects. American Mathematical Monthly, 105 (9):850–858, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Schnabel:1998:GDK

[876] Dan Schnabel. Goodbye, Descartes, by Keith Devlin. American Mathematical Monthly, 105(9):870–??, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nunemacher:1998:FAB

[877] Jeffrey Nunemacher. Fourier Analysis and Boundary Value Problems, by Enrique A. González-Velasco. American Mathematical Monthly, 105(9): 874–??, November 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lucas:1998:FDC

[878] William F. Lucas. Fair Division: From Cake-Cutting to Dispute Resolution, by Steven J. Brams and Alan D. Taylor. American Mathematical Monthly, 105 (9):877–??, November 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ipsen:1998:IBK

[879] Ilse C. F. Ipsen and Carl D. Meyer. The idea behind Krylov methods. American Mathematical Monthly, 105(10): 889–899, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kedlaya:1998:PFS

[880] Kiran S. Kedlaya. Product-free subsets of groups. American Mathematical Monthly, 105(10):900–906, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sturmfels:1998:PEC

[881] Bernd Sturmfels. Polynomial equations and convex polytopes. American Mathematical Monthly, 105(10): 907–922, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fraenkel:1998:MII

[882] Aviezri S. Fraenkel. Multivision: An intractable impartial game with a linear winning strategy. American Mathematical Monthly, 105(10):923–928, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McCartin:1998:SDS

[883] Brian J. McCartin. Seven deadly sins of numerical computation. American Mathematical Monthly, 105(10): 929–941, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Woon:1998:GIE

[884] S. C. Woon. Generating integrals and an elementary proof of the finite harmonic series theorem by fractional calculus. American Mathematical Monthly, 105(10):942–945, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Redheffer:1998:MM

[885] Raymond M. Redheffer. Monotonicity of the maxima. American Mathematical Monthly, 105(10):945–948, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Komornik:1998:SPF

[886] Vilmos Komornik. A simple proof of Farkas’ lemma. American Mathematical Monthly, 105(10):949–950, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Guy:1998:NNN

[887] Richard K. Guy. Nothing’s new in number theory? American Mathematical Monthly, 105(10):951–954, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Blum:1998:JLM

[888] Lenore Blum. Julia, A Life in Mathematics, by Constance Reid. American Mathematical Monthly, 105(10): 964–??, December 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maki:1998:WWJ

[889] Daniel P. Maki. Would-be Worlds, by John L. Casti. American Mathematical Monthly, 105(10):972–??, December 1998. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hersh:1998:NSH

[890] Reuben Hersh. The Number Sense: How the Mind Creates Mathematics, by Stanislas Dehaene. American Mathematical Monthly, 105(10): 975–??, December 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McKenna:1999:LTO

[891] P. J. McKenna. Large torsional oscillations in suspension bridges revisited: fixing an old approximation. American Mathematical Monthly, 106(1):1– 18, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goodson:1999:ICR

[892] Geoffrey R. Goodson. Inverse conjugacies and reversing symmetry groups. American Mathematical Monthly, 106 (1):19–26, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fuchs:1999:MP

[893] Dmitry Fuchs and Serge Tabachnikov. More on paperfolding. American Mathematical Monthly, 106(1):27–35, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bragg:1999:TIH

[894] L. R. Bragg. Trigonometric integrals and Hadamard products. American Mathematical Monthly, 106(1):36–42, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Blecksmith:1999:CP

[895] Richard Blecksmith, Paul Erdős, and J. L. Selfridge. Cluster primes. American Mathematical Monthly, 106(1):43– 48, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zuehlke:1999:FLT

[896] John A. Zuehlke. Fermat’s last theorem for Gaussian integer exponents. American Mathematical Monthly, 106(1):49, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jamieson:1999:RFA

[897] M. J. Jamieson. On rational function approximations to square roots. American Mathematical Monthly, 106(1):50– 52, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

vanderPoorten:1999:NJS

[898] A. J. van der Poorten and P. G. Walsh. A note on Jacobi symbols and continued fractions. American Mathematical Monthly, 106(1):52–56, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bashmakova:1999:BLA

[899] I. G. Bashmakova and G. S. Smirnova. The birth of literal algebra. Translated from the Russian and edited by Abe Shenitzer. American Mathematical Monthly, 106(1):57–66, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ross:1999:YGG

[900] Kenneth A. Ross. Yueh-Gin Gung and Dr. Charles Y. Hu Award for Distinguished Service to Leonard Gillman. American Mathematical Monthly, 106(2):97–98, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Feferman:1999:DMN

[901] Solomon Feferman. Does mathematics need new axioms. American Mathematical Monthly, 106(2):99–111, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goodman:1999:SIN

[902] Gerald S. Goodman. Statistical independence and normal numbers: an aftermath to Mark Kac’s Carus Monograph. American Mathematical Monthly, 106(2):112–126, 1999. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Long:1999:VDA

[903] Lyle N. Long and Howard Weiss. The velocity dependence of aerodynamic drag: a primer for mathematicians. American Mathematical Monthly, 106 (2):127–135, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cannell:1999:GGE

[904] D. M. Cannell. George Green: an enigmatic mathematician. American Mathematical Monthly, 106(2): 136–151, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Benjamin:1999:MSI

[905] Arthur T. Benjamin and Kan Yasuda. Magic “squares” indeed. American Mathematical Monthly, 106(2): 152–156, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sasvari:1999:EPB

[906] Zoltán Sasvári. An elementary proof of Binet’s formula for the gamma function. American Mathematical Monthly, 106(2):156–158, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Swan:1999:SPR

[907] Richard G. Swan. A simple proof of Rankin’s campanological theorem. American Mathematical Monthly, 106 (2):159–161, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sekino:1999:EMD

[908] Junpei Sekino. n-ellipses and the minimum distance sum problem. American Mathematical Monthly, 106(3): 193–202, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grinshpan:1999:BCM

[909] Arcadii Z. Grinshpan. The Bieberbach conjecture and Milin’s functionals. American Mathematical Monthly, 106(3):203–214, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sardelis:1999:DMG

[910] Dimitris A. Sardelis and Theodoros M. Valahas. Decision making: a golden rule. American Mathematical Monthly, 106(3):215–226, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Totik:1999:TTI

[911] Vilmos Totik. A tale of two integrals. American Mathematical Monthly, 106 (3):227–240, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Paradis:1999:MET

[912] Jaume Paradís, Pelegrí Viader, and Lluís Bibiloni. A mathematical excursion: from the three-door problem to a Cantor-type set. American Mathematical Monthly, 106(3):241–251, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Devaney:1999:MSF

[913] Robert L. Devaney. The Mandelbrot set, the Farey tree, and the Fibonacci sequence. American Mathematical Monthly, 106(4):289–302, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Richman:1999:EP

[914] Fred Richman. Existence proofs. American Mathematical Monthly, 106 (4):303–308, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Corazza:1999:IMP

[915] Paul Corazza. Introduction to metricpreserving functions. American Mathematical Monthly, 106(4):309–323, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Flury:1999:MD

[916] Bernard D. Flury, Robert Irving, and M. N. Goria. Magic dice. American Mathematical Monthly, 106(4):

324–337, 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Granville:1999:SDG

[917] Andrew Granville and Friedrich Roesler. The set of differences of a given set. American Mathematical Monthly, 106(4):338–344, 1999. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anoulova:1999:SWL

[918] S. Anoulova, J. Bennies, J. Lenhard, D. Metzler, Y. Sung, and A. Weber. Six ways of looking at Burtin’s lemma. American Mathematical Monthly, 106 (4):345–351, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maehara:1999:LTI

[919] Hiroshi Maehara. Lexell’s theorem via an inscribed angle theorem. American Mathematical Monthly, 106(4): 352–353, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589558.

Boyadzhiev:1999:CPD

[920] Khristo Boyadzhiev. A characteristic property of differentiation. American Mathematical Monthly, 106(4): 353–355, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589559.

Kedlaya:1999:WMM

[921] Kiran S. Kedlaya. A weighted mixedmean inequality. American Mathematical Monthly, 106(4):355–358, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589560.

Caines:1999:PTS

[922] Ian Caines, Carrie Gates, Richard K. Guy, and Richard J. Nowakowski. Periods in taking and splitting games. American Mathematical Monthly, 106 (4):359–361, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589561.

Mihai:1999:P

[923] Vasile Mihai. Problem 10725. American Mathematical Monthly, 106(4):362, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589562.

Knuth:1999:P

[924] Donald E. Knuth. Problem 10726. American Mathematical Monthly, 106 (4):362, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589563.

Anglesio:1999:P

[925] Jean Anglesio. Problem 10727. American Mathematical Monthly, 106(4):362, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589564.

Andreescu:1999:P

[926] Titu Andreescu. Problem 10728. American Mathematical Monthly, 106 (4):362, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589565.

Bellamy:1999:P

[927] David P. Bellamy and Felix Lazebnik. Problem 10729. American Mathematical Monthly, 106(4):362, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589566.

Janous:1999:P

[928] Walther Janous. Problem 10730. American Mathematical Monthly, 106 (4):363, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589567.

Pelling:1999:Pa

[929] M. J. Pelling. Problem 10731. American Mathematical Monthly, 106(4):363, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589568.

Goffinet:1999:CSP

[930] Daniel Goffinet, Kiran S. Kedlaya, Kenneth Schilling, Arlo W. Schurle, Fredric D. Ancel, Phil Bowers, and John Bryant. Connected sets of periodic functions: 10434. American Mathematical Monthly, 106(4): 363–364, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589569.

Cavachi:1999:PCD

[931] Marius Cavachi, Sam Northshield, Simeon T. Stefanov, and Victor Klee. The plane covered by disks: 10440. American Mathematical Monthly, 106 (4):364–365, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589570.

Rey:1999:RPM

[932] Joaquin Gomez Rey and Jose Heber Nieto. Random perfect matchings: 10587. American Mathematical Monthly, 106 (4):365, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589571.

Mazur:1999:CMT

[933] Marcin Mazur. Characterizations of the medial triangle: 10588. American Mathematical Monthly, 106(4): 365–366, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589572.

Bateman:1999:BEK

[934] Paul Bateman, David Bradley, David Callan, and Richard Stong. Binary expansions and k th powers: 10596. American Mathematical Monthly, 106 (4):366–367, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589573.

Rosenblatt:1999:SSD

[935] Joseph Rosenblatt, Thomas Hermann, and John H. Lindsey, II. A supremum of sine differences: 10604. American Mathematical Monthly, 106(4):368, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589574.

Vukmirovic:1999:CS

[936] Jovan Vukmirovic and Kenneth Schilling. A convergent series: 10657. American Mathematical Monthly, 106 (4):368, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589575.

Rothman:1999:BRF

[937] Tony Rothman. Book review: The French Mathematician, by Tom Petsinis. American Mathematical Monthly, 106(4):369–373, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589576.

Gold:1999:BRS

[938] Bonnie Gold. Book review: Social Constructivism as a Philosophy of Mathematics, by Paul Ernest; What is Mathematics, Really? by Reuben Hersh. American Mathematical Monthly, 106 (4):373–380, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589577.

Ostebee:1999:TRa

[939] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 106 (4):381–384, April 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589578.

Anonymous:1999:BMa

[940] Anonymous. Back matter. American Mathematical Monthly, 106(4):??, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589579.

Anonymous:1999:FMa

[941] Anonymous. Front matter. American Mathematical Monthly, 106(4):??, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589551.

Francis:1999:CZP

[942] George K. Francis and Jeffrey R. Weeks. Conway’s ZIP proof. American Mathematical Monthly, 106(5): 393–399, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589143.

Kraft:1999:CCS

[943] Roger L. Kraft. Chaos, Cantor sets, and hyperbolicity for the logistic maps. American Mathematical Monthly, 106 (5):400–408, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589144.

Apostol:1999:EVE

[944] Tom M. Apostol. An elementary view of Euler’s summation formula. American Mathematical Monthly, 106 (5):409–418, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589145.

Leep:1999:MMS

[945] David B. Leep and Gerry Myerson. Marriage, magic, and solitaire. American Mathematical Monthly, 106(5): 419–429, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589146.

Howards:1999:IPS

[946] Hugh Howards, Michael Hutchings, and Frank Morgan. The isoperimetric problem on surfaces. American Mathematical Monthly, 106(5): 430–439, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589147.

Chow:1999:WCF

[947] Timothy . Chow. What is a closedform number? American Mathematical Monthly, 106(5):440–448, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589148.

Martin:1999:SSP

[948] Greg Martin. The smallest solution of ϕ(30n + 1) is …. American Mathematical Monthly, 106(5): 449–451, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589149.

Kenter:1999:MRE

[949] Frank K. Kenter. A matrix representation for Euler’s constant, γ. American Mathematical Monthly, 106(5):

452–454, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589150.

Fejzic:1999:MMV

[950] H. Fejzic and D. Rinne. More on a Mean Value Theorem converse. American Mathematical Monthly, 106(5): 454–455, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589151.

Lange:1999:ECF

[951] L. J. Lange. An elegant continued fraction for π. American Mathematical Monthly, 106(5):456–458, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589152.

Beck:1999:RLD

[952] Matthias Beck. The reciprocity law for Dedekind sums via the constant Ehrhart coefficient. American Mathematical Monthly, 106(5):459–462, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589153.

Laugwitz:1999:RDE

[953] Detlef Laugwitz. Riemann’s dissertation and its effect on the evolution of mathematics. American Mathematical Monthly, 106(5):463–469, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589154.

Deshpande:1999:P

[954] M. N. Deshpande. Problems: 10732. American Mathematical Monthly, 106 (5):470, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589155.

Kim:1999:Pa

[955] Sung Soo Kim. Problems: 10733. American Mathematical Monthly, 106 (5):470, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589156.

vanLamoen:1999:P

[956] Floor van Lamoen. Problems: 10734. American Mathematical Monthly, 106 (5):470, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589157.

Simmons:1999:P

[957] Gustavus J. Simmons. Problems: 10735. American Mathematical Monthly, 106(5):470, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589158.

Khan:1999:P

[958] Mizan R. Khan. Problems: 10736. American Mathematical Monthly, 106 (5):470, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589159.

Ali:1999:P

[959] Hassan Ali Shah Ali. Problems: 10737. American Mathematical Monthly, 106 (5):471, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589160.

Theodorescu:1999:P

[960] Radu Theodorescu. Problems: 10738. American Mathematical Monthly, 106 (5):471, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589161.

Chang:1999:MRC

[961] Fu-Chuen Chang, Paul Deiermann, Walter Van Assche, Franz Peherstorfer, and Johannes Kepler. Moments of roots of Chebyshev polynomials: 10448. American Mathematical Monthly, 106(5):471–472, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589162.

Keller:1999:IN

[962] Tim Keller and Gchq Problem Solving Group. Indecomposable numbers: 10589. American Mathematical Monthly, 106(5):473, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589163.

Knuth:1999:NCV

[963] Donald E. Knuth and Gchq Problem Solving Group. Negatively correlated vectors of signs: 10593. American Mathematical Monthly, 106(5):

473–474, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589164.

Cox:1999:TWE

[964] David Cox and Gchq Problem Solving Group. n-tuples whose elements divide their sum: 10597. American Mathematical Monthly, 106(5): 474–475, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589165.

Krafft:1999:BR

[965] Olaf Krafft, Martin Schaefer, Nora Thornber, The National Security Agency Problems Group, and Ulrich Abel. Binomial ratios: 10625. American Mathematical Monthly, 106(5): 475–476, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589166.

Bencze:1999:TI

[966] Mihaly Bencze and Gchq Problems Group. A triangle inequality: 10644. American Mathematical Monthly, 106 (5):476, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589167.

Bhatia:1999:LR

[967] N. P. Bhatia. Limit of a recurrence: 10648. American Mathematical Monthly, 106(5):476–477, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589168.

Mallows:1999:RPR

[968] Colin L. Mallows and Kenneth Schilling. Random polynomials with real roots: 10660. American Mathematical Monthly, 106(5):477, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589169.

Schnabel:1999:BRL

[969] Dan Schnabel. Book review: Life’s Other Secret, by Ian Stewart; The Magical Maze, by Ian Stewart. American Mathematical Monthly, 106(5): 478–481, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589170.

Woodburn:1999:BRI

[970] Cynthia Woodburn. Book review: An Introductory Course in Commutative Algebra. by A. W. Chatters, C. R. Hajarnavis; Introduction to Algebra, by Peter J. Cameron. American Mathematical Monthly, 106(5): 481–483, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589171.

Ostebee:1999:TRb

[971] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 106 (5):484–487, May 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589172.

Anonymous:1999:BMb

[972] Anonymous. Back matter. American Mathematical Monthly, 106(5):??, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589173.

Anonymous:1999:FMb

[973] Anonymous. Front matter. American Mathematical Monthly, 106(5):??, May 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589142.

Lax:1999:CVM

[974] Peter D. Lax. Change of variables in multiple integrals. American Mathematical Monthly, 106(6): 497–501, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589462.

Billera:1999:RWP

[975] Louis J. Billera, Kenneth S. Brown, and Persi Diaconis. Random walks and plane arrangements in three dimensions. American Mathematical Monthly, 106(6):502–524, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589463.

Khan:1999:CFP

[976] Mizan R. Khan. A counting formula for primitive tetrahedra in Z^{3}. American Mathematical Monthly, 106 (6):525–533, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589464.

Rice:1999:WMG

[977] Adrian Rice. What makes a great mathematics teacher? The case of Augustus De Morgan. American Mathematical Monthly, 106(6): 534–552, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589465.

Gerth:1999:TAI

[978] Frank Gerth III. The TEAM approach to investing. American Mathematical Monthly, 106(6):553–558, June/ July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589466.

Schwind:1999:PMF

[979] William J. Schwind, Jun Ji, and Daniel E. Koditschek. A physically motivated further note on the mean value theorem for integrals. American Mathematical Monthly, 106(6): 559–564, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589467.

Butler:1999:HHD

[980] James P. Butler. Hopping hoops don’t hop. American Mathematical Monthly, 106(6):565–568, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589468.

Mathe:1999:AHC

[981] Peter Mathe. Approximation of Holder continuous functions by Bernstein polynomials. American Mathematical Monthly, 106(6):568–574, June/ July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589469.

deGuzman:1999:EWS

[982] Miguel de Guzman. An extension of the Wallace–Simson Theorem: Projecting in arbitrary directions. American Mathematical Monthly, 106(6): 574–580, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589470.

Hirschhorn:1999:ASP

[983] Michael D. Hirschhorn. Another short proof of Ramanujan’s mod5 partition congruence, and more. American Mathematical Monthly, 106(6): 580–583, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589471.

Snevily:1999:CAT

[984] Hunter S. Snevily. The Cayley addition table of Z_{n}. American Mathematical Monthly, 106(6):584–585, June/ July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589472.

Ciaurri:1999:P

[985] Oscar Ciaurri. Problems: 10739. American Mathematical Monthly, 106 (6):586, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589473.

VandenEynden:1999:P

[986] Charles Vanden Eynden. Problems: 10740. American Mathematical Monthly, 106(6):586, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589474.

Keller:1999:P

[987] Tim Keller. Problems: 10741. American Mathematical Monthly, 106(6): 586, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589475.

Alkan:1999:P

[988] Emre Alkan. Problems: 10742. American Mathematical Monthly, 106(6): 586, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589476.

Popescu:1999:Pa

[989] Călin Popescu. Problems: 10743. American Mathematical Monthly, 106 (6):586, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589477.

Lindqvist:1999:P

[990] Peter Lindqvist. Problems: 10744. American Mathematical Monthly, 106 (6):587, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589478.

Pelling:1999:Pb

[991] M. J. Pelling. Problems: 10745. American Mathematical Monthly, 106 (6):587, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589479.

Tabov:1999:UWF

[992] Jordan Tabov and Murray S. Klamkin. Using the walls to find the center: 10386. American Mathematical Monthly, 106(6):587, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589480.

Bang:1999:RSI

[993] Seung-Jin Bang, Jeremy E. Dawson, A. N. ’t Woord, O. P. Lossers, and Victor Hernandez. A reciprocal summation identity: 10490. American Mathematical Monthly, 106(6): 588–590, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589481.

Ionin:1999:LP

[994] Yury J. Ionin and Robin R. Lewis. A limit of periods: 10603. American Mathematical Monthly, 106(6): 590, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589482.

Zaslavsky:1999:AI

[995] Thomas Zaslavsky and Stephen M. Gagola, Jr. Avoiding the identity: 10606. American Mathematical Monthly, 106(6):590–591, June/ July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589483.

Conolly:1999:SSR

[996] Brian Conolly, Allen Stenger, and Thomas Hermann. Some sums require care: 10638. American Mathematical Monthly, 106(6):591–592, June/ July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589484.

Diaz:1999:BRI

[997] Ricardo Diaz. Book review: An Imaginary Tale: The Story of \sqrt{−1}, by Paul J. Nahin. American Mathematical Monthly, 106(6):593–594, June/ July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589485.

Beebee:1999:BRL

[998] John Beebee and Karen Willmore. Book review: Leaning towards Infinity, by Sue Woolfe. American Mathematical Monthly, 106(6):594–596, June/ July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589486.

Ostebee:1999:TRc

[999] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 106 (6):597–600, June/July 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589487.

Anonymous:1999:BMc

[1000] Anonymous. Back matter. American Mathematical Monthly, 106(6): ??, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589488.

Anonymous:1999:FMc

[1001] Anonymous. Front matter. American Mathematical Monthly, 106(6): ??, June/July 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589461.

Pritchett:1999:HHR

[1002] Timothy Pritchett. The hopping hoop revisited. American Mathematical Monthly, 106(7):609–617, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589490.

Fenster:1999:WDL

[1003] Della Dumbaugh Fenster. Why Dickson left quadratic reciprocity out of his history of the theory of numbers. American Mathematical Monthly, 106(7):618–627, August/September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589491.

Bridger:1999:UCL

[1004] Mark Bridger and Gabriel Stolzenberg. Uniform calculus and the law of bounded change. American Mathematical Monthly, 106(7):628–635, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589492.

Gordon:1999:AWQ

[1005] Gary Gordon. The answer is 2^{n}.n! What’s the question? American Mathematical Monthly, 106(7):636–645, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589493.

Trench:1999:CCI

[1006] William F. Trench. Conditional convergence of infinite products. American Mathematical Monthly, 106(7):646– 651, August/September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589494.

Clarke:1999:HJS

[1007] F. W. Clarke, W. N. Everitt, L. L. Littlejohn, and S. J. R. Vorster. H. J. S. Smith and the Fermat two squares theorem. American Mathematical Monthly, 106(7):652–665, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589495.

Nathanson:1999:NTS

[1008] Melvyn B. Nathanson. Number theory and semigroups of intermediate growth. American Mathematical Monthly, 106(7):666–669, August/ September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.
org/stable/2589496.

McCutcheon:1999:GHT

[1009] Randall McCutcheon. The Gottschalk– Hedlund theorem. American Mathematical Monthly, 106(7):670–672, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589497.

Tokieda:1999:MVT

[1010] Tadashi F. Tokieda. A mean value theorem. American Mathematical Monthly, 106(7):673–674, August/ September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.
org/stable/2589498.

Sury:1999:EJ

[1011] B. Sury. On an example of Jacobson. American Mathematical Monthly, 106(7):675–676, August/September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589499.

Kleiner:1999:FTEa

[1012] Israel Kleiner. Field theory: From equations to axiomatization. American Mathematical Monthly, 106(7):677– 684, August/September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589500.

Popescu:1999:Pb

[1013] Călin Popescu. Problems: 10743.

American Mathematical Monthly, 106 (7):685, August/September 1999. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589501.

Tersian:1999:P

[1014] Stepan Tersian. Problems: 10746. American Mathematical Monthly, 106 (7):685, August/September 1999. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

                                        http://www.jstor.org/stable/2589502.

Kalakos:1999:P

[1015] Athanasios Kalakos. Problems: 10747. American Mathematical Monthly, 106 (7):685, August/September 1999. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

                                        http://www.jstor.org/stable/2589503.

Borosh:1999:P

[1016] Itshak Borosh, Douglas A. Hensley, and Joel Zinn. Problems: 10748. American Mathematical Monthly, 106(7):685, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589504.

Grigis:1999:P

[1017] Alain Grigis. Problems: 10749. American Mathematical Monthly, 106(7):685, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589505.

Smiley:1999:P

[1018] Leonard Smiley. Problems: 10750. American Mathematical Monthly, 106 (7):686, August/September 1999. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

                                        http://www.jstor.org/stable/2589506.

Deutsch:1999:P

[1019] Emeric Deutsch. Problems: 10751.

American Mathematical Monthly, 106 (7):686, August/September 1999. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589507.

Costovici:1999:P

[1020] Gh. Costovici. Problems: 10752.

American Mathematical Monthly, 106 (7):686, August/September 1999. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2589508.

Silverman:1999:ZFR

[1021] Joseph H. Silverman, David Bradley, and Donald A. Darling. A zeta function over a recurrent sequence: 10486. American Mathematical Monthly, 106(7):686–688, August/ September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.
org/stable/2589509.

Galvin:1999:MI

[1022] Fred Galvin, Frank Jelen, and Eberhard Triesch. A matrix of inequalities: 10599. American Mathematical Monthly, 106(7):688, August/ September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.
org/stable/2589510.

Ma:1999:CD

[1023] Wen-Xiu Ma, Robin J. Chapman, and Joseph J. Rushanan. A complex determinant: 10601. American Mathematical Monthly, 106(7):688–689, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589511.

Knuth:1999:CEN

[1024] Donald E. Knuth and David Callan. On a convolution of Eulerian numbers: 10609. American Mathematical Monthly, 106(7):690–691, August/ September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.
org/stable/2589512.

Hall:1999:VAB

[1025] Richard Hall, National Security Agency Problems Group, and Gchq Problems Group. A variation on additive bases: 10610. American Mathematical Monthly, 106(7):691–692, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589513.

Rey:1999:IIR

[1026] Joaquin Gomez Rey, Anchorage Math Solutions Group, and Gchq Problems Group. An identity involving rooted trees: 10615. American Mathematical Monthly, 106(7):692–693, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589514.

Merickel:1999:DSD

[1027] James G. Merickel and John P. Robertson. Divisors of sums of divisors: 10617. American Mathematical Monthly, 106(7):693, August/ September 1999. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.
org/stable/2589515.

Mullin:1999:BRM

[1028] Albert A. Mullin. Book review: My Brain is Open: The Mathematical Journeys of Paul Erdős, by Bruce Schechter; The Man Who Loved Only Numbers: The Story of Paul Erdős and the Search for Mathematical Truth, by Paul Hoffman. American Mathematical Monthly, 106(7):694–696, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589516.

Ostebee:1999:TRd

[1029] Arnold Ostebee. Telegraphic reviews. American Mathematical Monthly, 106(7):697–700, August/ September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.
org/stable/2589517.

Anonymous:1999:LRF

[1030] Anonymous. Lester R. Ford awards for 1998. American Mathematical Monthly, 106(7):701–702, August/ September 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.
org/stable/2589518.

Anonymous:1999:BMd

[1031] Anonymous. Back matter. American Mathematical Monthly, 106(7):??, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589519.

Anonymous:1999:FMd

[1032] Anonymous. Front matter. American Mathematical Monthly, 106(7):??, August/September 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589489.

Terrell:1999:SFCa

[1033] William J. Terrell. Some fundamental control theory I: Controllability, observability, and duality. American Mathematical Monthly, 106(8): 705–719, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589018.

Pourciau:1999:EPM

[1034] Bruce Pourciau. The education of a pure mathematician. American Mathematical Monthly, 106(8):720–732, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589019.

Velleman:1999:MCP

[1035] Daniel J. Velleman. Multivariable calculus and the plus topology. American Mathematical Monthly, 106(8): 733–740, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589020.

Hubbard:1999:FDP

[1036] John H. Hubbard. The forced damped pendulum: Chaos, complication and control. American Mathematical Monthly, 106(8):741–758, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589021.

Ungar:1999:HPT

[1037] Abraham A. Ungar. The hyperbolic Pythagorean theorem in the Poincaré disc model of hyperbolic geometry. American Mathematical Monthly, 106 (8):759–763, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589022.

Lu:1999:CFI

[1038] Jitan Lu. Is the composite function integrable? American Mathematical Monthly, 106(8):763–766, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589023.

Shen:1999:GLG

[1039] Jianhong Shen. On the generalized ”Lanczos’ generalized derivative”. American Mathematical Monthly, 106 (8):766–768, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589024.

Rudin:1999:ST

[1040] Walter Rudin. A stability theorem.

American Mathematical Monthly, 106 (8):768–770, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589025.

Alperin:1999:RMG

[1041] Roger C. Alperin. Rationals and the modular group. American Mathematical Monthly, 106(8):771–773, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589026.

Osler:1999:UVW

[1042] Thomas J. Osler. The union of Vieta’s and Wallis’s products for pi. American Mathematical Monthly, 106 (8):774–776, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589027.

Shapiro:1999:P

[1043] Louis Shapiro. Problems: 10753. American Mathematical Monthly, 106 (8):777, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589028.

Bracken:1999:P

[1044] Paul Bracken. Problems: 10754. American Mathematical Monthly, 106(8): 777, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589029.

Fukuta:1999:P

[1045] Jiro Fukuta. Problems: 10755. American Mathematical Monthly, 106(8): 777, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589030.

Iannucci:1999:P

[1046] Douglas Iannucci. Problems: 10756. American Mathematical Monthly, 106 (8):778, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589031.

Kidwell:1999:P

[1047] Mark Kidwell. Problems: 10757.

American Mathematical Monthly, 106 (8):778, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589032.

Sapir:1999:P

[1048] Mark Sapir. Problems: 10758. American Mathematical Monthly, 106(8): 778, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589033.

Popescu:1999:Pc

[1049] Călin Popescu. Problems: 10759.

American Mathematical Monthly, 106 (8):778, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589034.

Kedlaya:1999:CEC

[1050] Kiran S. Kedlaya and Alain Tissier.

Common eigenvector of commuting matrices: 10633. American Mathematical Monthly, 106(8):778–779, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589035.

Parry:1999:RCL

[1051] C. F. Parry and Robert L. Young. Reflected concurrent lines: 10637. American Mathematical Monthly, 106(8): 779–780, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589036.

Ali:1999:CM

[1052] Hassan Ali Shah Ali and Patrick A. Staley. A constrained maximization: 10646. American Mathematical Monthly, 106(8):780–781, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589037.

Sasvari:1999:PSE

[1053] Zoltan Sasvari and William F. Trench. A Pólya–Szegő exercise revisited: 10650. American Mathematical Monthly, 106(8):781–782, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589038.

Hayman:1999:HPH

[1054] W. K. Hayman, Tewodros Amdeberhan, and Irl C. Bivens. Harmonic products of harmonic functions: 10651. American Mathematical Monthly, 106 (8):782, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589039.

Bellamy:1999:LVT

[1055] David P. Bellamy, Felix Lazebnik, Jeffrey Lagarias, and Stephen M. Gagola, Jr. Large values of tangent: 10656. American Mathematical Monthly, 106 (8):782–784, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589040.

Mihai:1999:EPC

[1056] Vasile A. Mihai and J. Schaer. The ellipse in a paper cup: 10664. American Mathematical Monthly, 106(8): 784, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589041.

Koch:1999:BRF

[1057] John A. Koch. Book review: The Four-Color Theorem. by Rudolf Fritsch, Gerda Fritsch, J. Peschke. American Mathematical Monthly, 106(8): 785–787, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589042.

Ostebee:1999:TRe

[1058] Arnold Ostebee. Telegraphic reviews.

American Mathematical Monthly, 106 (8):788–791, October 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589043.

Anonymous:1999:BMe

[1059] Anonymous. Back matter. American Mathematical Monthly, 106(8): ??, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589044.

Anonymous:1999:FMe

[1060] Anonymous. Front matter. American Mathematical Monthly, 106(8): ??, October 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589017.

Archer:1999:MTP

[1061] Aaron F. Archer. A modern treatment of the 15 puzzle. American Mathematical Monthly, 106(9):793–799, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589612.

Hurley:1999:ECR

[1062] James F. Hurley, Uwe Koehn, and Susan L. Ganter. Effects of calculus reform: Local and national. American Mathematical Monthly, 106(9): 800–811, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589613.

Terrell:1999:SFCb

[1063] William J. Terrell. Some fundamental control theory II: Feedback linearization of single input nonlinear systems. American Mathematical Monthly, 106 (9):812–828, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589614.

Baker:1999:DPE

[1064] John A. Baker. The Dirichlet problem for ellipsoids. American Mathematical Monthly, 106(9):829–834, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589615.

Singer:1999:CWC

[1065] David A. Singer. Curves whose curvature depends on distance from the origin. American Mathematical Monthly, 106(9):835–841, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589616.

Klosinski:1999:FNW

[1066] Leonard F. Klosinski, Gerald L.

Alexanderson, and Loren C. Larson. The Fifty-Ninth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 106 (9):842–849, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589617.

Thomassen:1999:NUD

[1067] Carsten Thomassen. On the Nelson unit distance coloring problem. American Mathematical Monthly, 106 (9):850–853, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589618.

Grabiner:1999:DRS

[1068] David J. Grabiner. Descartes’ rule of signs: Another construction. American Mathematical Monthly, 106(9): 854–856, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589619.

McCammond:1999:IPS

[1069] Jonathan P. McCammond. Integrating polynomials in secant and tangent. American Mathematical Monthly, 106 (9):856–858, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589620.

Kleiner:1999:FTEb

[1070] Israel Kleiner. Field theory: From equations to axiomatization. American Mathematical Monthly, 106(9): 859–863, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589621.

Reznick:1999:P

[1071] Bruce Reznick. Problems: 10760. American Mathematical Monthly, 106 (9):864, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589622.

Galvin:1999:P

[1072] Fred Galvin. Problems: 10761. American Mathematical Monthly, 106(9): 864, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589623.

Quet:1999:P

[1073] Leroy Quet. Problems: 10762. American Mathematical Monthly, 106(9): 864, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589624.

Anglesio:1999:Pa

[1074] Jean Anglesio. Problems: 10763. American Mathematical Monthly, 106 (9):864, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589625.

Redheffer:1999:P

[1075] Ray Redheffer. Problems: 10764.

American Mathematical Monthly, 106 (9):864, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589626.

Ferraro:1999:P

[1076] Peter J. Ferraro. Problems: 10765.

American Mathematical Monthly, 106 (9):864, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589627.

Andras:1999:P

[1077] Szilard Andras. Problems: 10766.

American Mathematical Monthly, 106 (9):865, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589628.

Lakshminarayanan:1999:CRN

[1078] S. Lakshminarayanan, S. L. Shah, K. Nandakumar, and Gchq Problems Group. Cramer’s rule for non-square matrices: 10618. American Mathematical Monthly, 106(9):865, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589629.

Propp:1999:ISC

[1079] James Propp. An identity for strongly connected digraphs: 10620. American Mathematical Monthly, 106(9): 865–867, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589630.

Deshpande:1999:SSA

[1080] M. N. Deshpande, Hansruedi Widmer, and Zachary Franco. Simultaneous squares from arithmetic progressions: 10622. American Mathematical Monthly, 106(9):867–868, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589631.

Barbara:1999:DPI

[1081] Roy Barbara, Nasha Komanda, and Robin J. Chapman. The divisibility poset inside itself: 10623. American Mathematical Monthly, 106(9): 868–869, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589632.

Luca:1999:EIT

[1082] Florian Luca and Nasha Komanda. An equation involving the totient: 10626. American Mathematical Monthly, 106 (9):869–870, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589633.

Fukuta:1999:TCC

[1083] Jiro Fukuta, Gchq Problems Group, and Călin Popescu. Three congruent circles between two triangles: 10659. American Mathematical Monthly, 106 (9):870, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589634.

Diamond:1999:LCS

[1084] Harold G. Diamond and Nathaniel Grossman. Logarithmic convexity of Stirling’s ratio: 10680. American Mathematical Monthly, 106(9): 871, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589636.

Konhauser:1999:QCG

[1085] Joseph D. E. Konhauser, Stan Wagon, and Con Amore Problems Group. Quadrilateral center of gravity: 10662. American Mathematical Monthly, 106 (9):871, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589635.

Rauff:1999:BRA

[1086] James V. Rauff. Book review: African Fractals: Modern Computing and Indigenous Design. by Ron Eglash. American Mathematical Monthly, 106 (9):872–875, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589637.

Ostebee:1999:TRf

[1087] Arnold Ostebee. Telegraphic reviews.

American Mathematical Monthly, 106 (9):876–879, November 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589638.

Anonymous:1999:BMf

[1088] Anonymous. Back matter. American Mathematical Monthly, 106(9): ??, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589639.

Anonymous:1999:FMf

[1089] Anonymous. Front matter. American Mathematical Monthly, 106(9): ??, November 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589611.

Borwein:1999:ETE

[1090] Jonathan M. Borwein and Robert M. Corless. Emerging tools for experimental mathematics. American Mathematical Monthly, 106(10):889–909, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589743.

Tucker:1999:RTS

[1091] Thomas W. Tucker. Reform, tradition, and synthesis. American Mathematical Monthly, 106(10):910–914, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589744.

Krantz:1999:DN

[1092] Steven G. Krantz. You don’t need a weatherman to know which way the wind blows. American Mathematical Monthly, 106(10):915–918, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589745.

Shapiro:1999:WC

[1093] Helene Shapiro. The Weyr characteristic. American Mathematical Monthly, 106(10):919–929, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589746.

Su:1999:RHS

[1094] Francis Edward Su. Rental harmony:

Sperner’s lemma in fair division. American Mathematical Monthly, 106(10): 930–942, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589747.

Hanin:1999:WTE

[1095] Leonid G. Hanin. Which tanks empty faster? American Mathematical Monthly, 106(10):943–947, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589748.

Callan:1999:TUD

[1096] David Callan. Two uniformly distributed parameters defining Catalan numbers. American Mathematical Monthly, 106(10):948–949, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589749.

Kovacs:1999:DCB

[1097] Istvan Kovacs, Daniel S. Silver, and Susan G. Williams. Determinants of commuting-block matrices. American Mathematical Monthly, 106(10): 950–952, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589750.

iu:1999:MI

[1098] Paul iu. Mixtilinear incircles. American Mathematical Monthly, 106(10): 952–955, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589751.

Etter:1999:AMP

[1099] David N. Etter. The area of the medial parallelogram of a tetrahedron. American Mathematical Monthly, 106 (10):956–958, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589752.

Nowakowski:1999:UP

[1100] Richard Nowakowski. Unsolved problems, 1969–1999. American Mathematical Monthly, 106(10):959–962, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589753.

Dearden:1999:P

[1101] Bruce Dearden and Jerry Metzger. Problems: 10767. American Mathematical Monthly, 106(10):963, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2589754.

Kim:1999:Pb

[1102] Sung Soo Kim. Problems: 10768. American Mathematical Monthly, 106 (10):963, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589755.

Blatter:1999:P

[1103] Christian Blatter. Problems: 10769.

American Mathematical Monthly, 106 (10):963, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589756.

Popescu:1999:Pd

[1104] Călin Popescu. Problems: 10770.

American Mathematical Monthly, 106 (10):963, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589757.

Hajja:1999:P

[1105] Mowaffaq Hajja and Peter Walker.

Problems: 10771. American Mathematical Monthly, 106(10):963, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589758.

Waterhouse:1999:P

[1106] William C. Waterhouse. Problems: 10772. American Mathematical Monthly, 106(10):963, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589759.

Anglesio:1999:Pb

[1107] Jean Anglesio. Problems: 10773.

American Mathematical Monthly, 106 (10):964, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589760.

Huber:1999:TI

[1108] Greg Huber and Gchq Problems Group. Tracking the incenters: 10631. American Mathematical Monthly, 106 (10):964–965, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589761.

Trench:1999:ABF

[1109] William F. Trench and Ronald A. Kopas. An appearance of the beta function: 10632. American Mathematical Monthly, 106(10):965, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589762.

Farnum:1999:AC

[1110] Nicholas R. Farnum and Alain Tissier. Apery’s constant: 10635. American Mathematical Monthly, 106(10): 965–966, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2589763.

Kagan:1999:EDF

[1111] Abram Kagan, Larry Shepp, and Kenneth Schilling. Essentially discontinuous functions: 10668. American Mathematical Monthly, 106(10):966– 967, December 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589764.

Benchimol:1999:TRR

[1112] Salomon Benchimol, Elliott Cohen, and Con Amore Problems Group. Two recurrence relations, one easy, one hard: 10670. American Mathematical Monthly, 106(10):967–968, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589765.

Rothe:1999:NZM

[1113] F. Rothe and Sung Soo Kim. The number of zeros of a Maclaurin polynomial: 10671. American Mathematical Monthly, 106(10):968, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2589766.

Kumar:1999:GV

[1114] V. Anil Kumar and John H. Lindsey, II.

An AM–GM variation: 10672. American Mathematical Monthly, 106(10): 969, December 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589767.

Tamvakis:1999:FPA

[1115] Harry Tamvakis and Gchq Problems Group. Functions with a polynomial addition formula: 10675. American Mathematical Monthly, 106(10): 969, December 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589768.

Alzer:1999:UI

[1116] Horst Alzer and M. J. Pelling. An unsettled inequality: 10337. American Mathematical Monthly, 106(10): 970, December 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589769.

Aboufadel:1999:BRW

[1117] Edward Aboufadel, Matthew Boelkins, and Steven Schlicker. Book review: Wavelets: A Primer. by Christian Blatter, A. K. Peters; Wavelets in a Box, by Charles K. Chui, Andrew K. Chan, C. Steve Liu; A Primer on Wavelets for Scientists and Engineers, by James S. Walker; Wavelet Analysis: The Scalable Structure of Information, by Howard L. Resnikoff, Raymond O. Wells, Jr. American Mathematical Monthly, 106(10):971– 977, December 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589770.

Gottlieb:1999:BRP

[1118] Daniel Henry Gottlieb. Book review: Poincaré and the Three Body Problem. by June Barrow-Green. American Mathematical Monthly, 106(10): 977–980, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589771.

Anonymous:1999:VI

[1119] Anonymous. Volume information. American Mathematical Monthly, 106 (10):981–992, December 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

                                        www.jstor.org/stable/2589741.

Anonymous:1999:BMg

[1120] Anonymous. Back matter. American Mathematical Monthly, 106(10): ??, December 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

                                        jstor.org/stable/2589772.

Anonymous:1999:FMg

[1121] Anonymous. Front matter. American Mathematical Monthly, 106(10): ??, December 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2589742.

Bailey:1997:RCV

[1122] David Bailey, Peter Borwein, and Simon Plouffe. On the rapid computation of various polylogarithmic constants. Mathematics of Computation, 66(218):903–913, April 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/
journals/mcom/1997-66-218/S0025-
5718-97-00856-9/S0025-5718-97-00856-

9.pdf. See [741].